### v1

# A Unified Implementation of Quasi-Monte Carlo Generators, Randomization Routines, and Fast Kernel Methods

Abstract.
Quasi-random sequences, also called low-discrepancy sequences, have been extensively used as efficient experimental designs across many scientific disciplines. This article provides a unified description and software API for methods pertaining to low-discrepancy point sets. These methods include low discrepancy point set generators, randomization techniques, and fast kernel methods. Specifically, we provide generators for lattices, digital nets, and Halton point sets. Supported randomization techniques include random permutations / shifts, linear matrix scrambling, and nested uniform scrambling. Routines for working with higher-order digital nets and scramblings are also detailed. For kernel methods, we provide implementations of special shift-invariant and digitally-shift invariant kernels along with fast Gram matrix operations facilitated by the bit-reversed FFT, the bit-reversed IFFT, and the FWHT. A new digitally-shift-invariant kernel of higher-order smoothness is also derived. We also describe methods to quickly update the matrix-vector product or linear system solution after doubling the number of points in a lattice or digital net in natural order.
Low-discrepancy sequences, low-discrepancy point set randomizations, higher-order digital net scrambling, randomized Quasi-Monte Carlo, fast kernel methods, shift-invariant kernels, digitally-shift-invariant kernels
copyright: acmlicensedjournalyear: 2025doi: 0.0journal: TOMSjournalvolume: 0journalnumber: 0article: 0publicationmonth: 0submissionid: 0-0-0ccs: Mathematics of computing Statistical softwareccs: Mathematics of computing Mathematical software performanceccs: Mathematics of computing Solversccs: Mathematics of computing Probabilistic algorithms
1. Introduction
This article describes and implements a number of algorithms related to low-discrepancy point sets. These include implementations of
  • low-discrepancy point set generators and randomizations,
  • higher-order digital net generators and randomizations,
  • shift-invariant and digitally-shift-invariant kernels of higher order smoothness,
  • and fast transforms which enable fast Gram matrix operations when special kernels are paired with certain low-discrepancy point sets.
While these methods have been developed and often extensively studied throughout the literature, we found implementations to be scattered or even missing in some popular programming languages. Our main contribution is to provide a unified and accessible software for computations using low-discrepancy sequences in the QMCPy [^1] https://qmcpy.readthedocs.io Python package (Choi et al., 2022b, a). We also integrate with the new LDData [^2] https://github.com/QMCSoftware/LDData repository which provides generating vectors for lattices and generating matrices for digital nets in standardized formats. This collection is, so far, mainly an aggregation of results from the Magic Point Shop [^3] https://people.cs.kuleuven.be/~dirk.nuyens/qmc-generators/ (Kuo and Nuyens, 2016a) and the websites of Frances Kuo on lattices [^4] https://web.maths.unsw.edu.au/~fkuo/lattice/index.html (Cools et al., 2006; Nuyens and Cools, 2006) and Sobol’ points [^5] https://web.maths.unsw.edu.au/~fkuo/sobol/index.html (Joe and Kuo, 2003, 2008).
Below we summarize the novel components of our QMCPy software which we detail in this paper. Comparisons to existing implementations are also discussed. A more comprehensive review of Quasi-Monte Carlo software is given in (Choi et al., 2022b), along with more details on the QMCPy framework.
Lattice point sets:
We support generating points in linear, natural, or Gray code order and randomizing points with shifts modulo one. Other implementations of shifted lattices include those in MATLAB’s GAIL [^6] http://gailgithub.github.io/GAIL_Dev/ (Guaranteed Automatic Integration Library) (Tong et al., 2022; Hickernell et al., 2018), the C++ LatNet Builder [^7] https://github.com/umontreal-simul/latnetbuilder software (L’Ecuyer et al., 2021), and the multi-language MPS [^8] https://people.cs.kuleuven.be/~dirk.nuyens/qmc-generators/ (Magic Point Shop) (Kuo and Nuyens, 2016b). LatNet Builder and MPS also provide search routines for finding good generating vectors which define lattice point sets.
Digital nets in base $2$:
We support generating digital nets in natural or Gray code order and randomizing with linear matrix scrambling (LMS) (Owen, 2003), digital shifts, nested uniform scrambling (NUS) (Owen, 1995), and permutation scrambling. To the best of our knowledge, we are also the only Python implementation to support digital interlacing for the creating of higher order digital nets and higher order scrambling with either LMS or NUS.
Early implementations of unrandomized digital sequences, including the Faure, Sobol’ and Niederreiter-Xing constructions, can be found in (Fox, 1986; Bratley et al., 1992; Bratley and Fox, 2003; Pirsic, 2002). Considerations for implementing scrambles for digital sequences was discussed in (Hong and Hickernell, 2003). Recent digital net implementations supporting LMS and digital shifts include those in MPS, PyTorch (Paszke et al., 2019), SciPy (Virtanen et al., 2020), and MATLAB (Inc., 2022). LatNet Builder provides robust support for all randomization routines and digital interlacing for higher order nets and scramblings. LatNet Builder and MPS also provide search routines for finding good generating matrices which define digital nets.
Halton point sets:
As with digital nets, we support randomizing Halton point sets with LMS, digital shifts, NUS, and permutation scrambles. The implementation of Halton point sets and randomizations have been treated in (Owen, 2017; Wang and Hickernell, 2000). The QRNG [^9] https://cran.r-project.org/web/packages/qrng/qrng.pdf (Quasi-Random Number Generators) R package (Hofert and Lemieux, 2023) implements generalized digital nets which use optimized digital permutation scrambles; these are also supported in QMCPy. LatNet Builder also provides support for Halton points and their randomizations.
Kernel Methods and Fast Transforms:
We provide routines to evaluate special shift-invariant and digitally-shift-invariant kernels. This includes a new digitally-shift-invariant kernel of higher-order smoothness derived in this paper. These are not readily available elsewhere.
We also provide routines for fast matrix-vector multiplication and inversion of the nicely-structured Gram matrices produced when pairing these special kernels with certain low-discrepancy sequences. We also describe novel routines to efficiently update the matrix-vector products and linear system solves in the Gram matrix after doubling the number of lattice or digital net points. These fast operations are backed by implementations of the Fast Walsh Hadamard Transform (FWHT), the bit-reversed fast Fourier Transform (FFT), and the bit-reversed inverse FFT (IFFT). We are among the first to provide efficient and accessible Python interfaces to such fast kernel methods and the underlying fast transforms. GAIL has implemented fast Bayesian cubature routines utilizing this machinery.
These above routines are most commonly used in Quasi-Monte Carlo methods and for fast kernel methods, which we describe in the following subsections.
1.1. Quasi-Monte Carlo methods
Monte Carlo (MC) methods approximate a high dimensional integral over the unit cube by the sample average of function evaluations at certain sampling locations:
(1) $\mu:=\int_{[0,1]^{d}}f({\boldsymbol{x}})\mathrm{d}{\boldsymbol{x}}\approx\frac{1}{n}\sum_{i=0}^{n-1}f({\boldsymbol{x}}_{i})=:{\widehat{\mu}}.$
Here $f:[0,1]^{d}\to{\mathbb{R}}$ is a given integrand and $\{{\boldsymbol{x}}_{i}\}_{i=0}^{n-1}\in[0,1]^{n\times d}$ is a point set. The integral on the left hand side may be viewed as taking the expectation ${\mathbb{E}}[f({\boldsymbol{X}})]$ where ${\boldsymbol{X}}$ is a standard uniform ${\boldsymbol{X}}\sim{\mathcal{U}}[0,1]^{d}$. Classic MC methods choose the sampling locations to be independent and identically distributed (IID) $d$-dimensional standard uniforms ${\boldsymbol{x}}_{0},\dots,{\boldsymbol{x}}_{n-1}\overset{\mathrm{IID}}{\sim}{\mathcal{U}}[0,1]^{d}$. IID-MC methods for (1) have a root-mean-squared-error (RMSE) of ${\mathcal{O}}(n^{-1/2})$.
Figure 1. An independent identically distributed (IID) point set and low-discrepancy (quasi-random) point sets of size $n=2^{13}=8192$. Notice the quasi-random points more evenly fill the space than IID points, which leads to faster convergence of Quasi-Monte Carlo methods compared to IID Monte Carlo methods. Randomized lattices, digital nets in base $b=2$ (DN2), and Halton points are shown. Randomizations include shifts, linear matrix scramblings (LMS), digital shifts (DS), and nested uniform scramblings (NUS). Digital interlacing of order $\alpha$ is used to create higher-order randomized digital nets (HOα).
[…]
Quasi-Monte Carlo (QMC) methods (Niederreiter, 1992; Dick and Pillichshammer, 2010; Kroese et al., 2013; Dick et al., 2022; Lemieux, 2009; Sloan and Joe, 1994) replace IID point sets with low-discrepancy (LD) point sets which more evenly cover the unit cube $[0,1]^{d}$. For integrands with bounded variation, plugging LD point sets into ${\widehat{\mu}}$ (1) yields a worst-case error rate of ${\mathcal{O}}(n^{-1+\delta})$ with $\delta>0$ arbitrarily small. Some popular LD point sets are plotted in fig. 1 including lattices, digital nets (including higher-order versions), and Halton point sets.
Randomized Quasi-Monte Carlo (RQMC) uses randomized LD point sets to give improved convergence rates and enable practical error estimation. Specifically, if we again assume the integrand has bounded variation, then certain RQMC methods can achieve a RMSE of ${\mathcal{O}}(n^{-3/2+\delta})$. Moreover, let $\{{\boldsymbol{x}}_{i}^{1}\}_{i=0}^{n-1},\dots,\{{\boldsymbol{x}}_{i}^{R}\}_{i=0}^{n-1}$ denote $R$ IID randomizations of a low-discrepancy point set where typically $R$ is small e.g. $R=15$. Then the RQMC estimate
(2) ${\widehat{\mu}}_{R}=\frac{1}{R}\sum_{r=1}^{R}\hat{\mu}^{r}\qquad\text{where}\quad\hat{\mu}^{r}=\frac{1}{n}\sum_{i=0}^{n-1}f({\boldsymbol{x}}_{i}^{r})$
may be used to construct a practical (approximate) $100(1-\tau)\%$ confidence interval ${\widehat{\mu}}_{R}\pm t_{R-1,\tau/2}{\widehat{\sigma}}_{R}/\sqrt{R}$ for $\mu$ where ${\widehat{\sigma}}_{R}^{2}=R^{-1}(R-1)^{-1}\sum_{r=1}^{R}({\widehat{\mu}}^{r}-{\widehat{\mu}}_{R})^{2}$ and $t_{R-1,\tau/2}$ is the $\tau/2$ quantile of a Student distribution with $R-1$ degrees of freedom. RQMC estimates are detailed in (Owen, 2018, Chapter 17), and a recent comparison of RQMC confidence interval methods can be found in (L’Ecuyer et al., 2023). QMC error estimation is treated more broadly in (Owen, 2024), while (Clancy et al., 2014; Hickernell et al., 2017; Clancy et al., 2014) detail additional considerations for adaptive QMC algorithms. For lattices, randomization is typically done using random shifts modulo $1$. For digital nets, randomization is typically done using nested uniform scrambling (NUS) or the cheaper combination of linear matrix scrambling (LMS) with digital shifts / permutations (Owen, 2003, 2017; Owen and Pan, 2024).
Higher-order LD point sets were designed to yield faster convergence for integrands with additional smoothness. For integrands with square integrable mixed partial derivatives up to order $\alpha>1$, plugging higher-order digital nets into ${\widehat{\mu}}$ (1) yields a worst-case error rate of ${\mathcal{O}}(n^{-\alpha+\delta})$ (Dick, 2008, 2009b, 2009a). RQMC using scrambled digital nets (via either NUS or LMS with digital permutations) combined with digital interlacing has been shown to achieve an RMSE of order ${\mathcal{O}}(n^{-\alpha-1/2+\delta})$ (Dick, 2011).
1.2. Kernel computations
Low-discrepancy (LD) point sets can also been used to accelerated many kernel computations. Let $\{{\boldsymbol{x}}_{i}\}_{i=0}^{n-1}$ be a point set, $K:[0,1]^{d}\times[0,1]^{d}\to{\mathbb{R}}$ be a symmetric positive definite (SPD) kernel, and ${\mathsf{K}}=\left(K({\boldsymbol{x}}_{i},{\boldsymbol{x}}_{k})\right)_{i,k=0}^{n-1}$ be the $n\times n$ SPD Gram matrix. Two motivating kernel computations are described below
Discrepancy Computation:
The error of the Monte Carlo approximation (1) can be bounded by the product two terms: a measure of variation of the function $f$ and a measure of the discrepancy of the point set $\{{\boldsymbol{x}}_{i}\}_{i=1}^{n-1}$. The most well known bound of this type is the Koksma-Hlawka inequality (Hickernell, 1998; Dick et al., 2013; Hickernell, 1999; Niederreiter, 1992). The discrepancy is frequently computed when designing and evaluating low-discrepancy point sets, see (Rusch et al., 2024) for a newer idea in this area where good point sets are generated using a neural networks trained with a discrepancy-based loss function.
Let us consider the more general setting of approximating the mean $\mu$ in (1) by a weighted cubature rule $\sum_{i=0}^{n-1}\omega_{i}f({\boldsymbol{x}}_{i})$. If $f$ is assumed to lie in the RKHS $H$, the discrepancy in the error bound of such a cubature rule takes the form
$\int\int K({\boldsymbol{u}},{\boldsymbol{v}})\mathrm{d}{\boldsymbol{u}}\mathrm{d}{\boldsymbol{v}}-2\sum_{i=0}^{n-1}\omega_{i}\int K({\boldsymbol{u}},{\boldsymbol{x}}_{i})\mathrm{d}{\boldsymbol{u}}+\sum_{i,k=0}^{n-1}\omega_{i}\omega_{k}K({\boldsymbol{x}}_{i},{\boldsymbol{x}}_{k})$
following (Hickernell, 1998). Here the integrals are understood to be over the unit cube $[0,1]^{d}$. Evaluating this discrepancy above requires computing ${\mathsf{K}}{\boldsymbol{\omega}}$ where ${\boldsymbol{\omega}}=\{\omega_{i}\}_{i=0}^{n-1}$. Moreover, the discrepancy is minimized for a given point set by setting the weights to ${\boldsymbol{\omega}}^{\star}={\mathsf{K}}^{-1}{\boldsymbol{\kappa}}$ where $\kappa_{i}=\int K({\boldsymbol{u}},{\boldsymbol{x}}_{i})\mathrm{d}{\boldsymbol{u}}$ for $0\leq i<n$.
Kernel Interpolation:
Suppose we would like to approximate $f:[0,1)^{d}\to{\mathbb{R}}$ given observations ${\boldsymbol{y}}=\{y_{i}\}_{i=0}^{n-1}$ of $f$ at $\{{\boldsymbol{x}}_{i}\}_{i=0}^{n-1}$ satisfying $y_{i}=f(x_{i})$. Then a kernel interpolant approximates $f$ by ${\widehat{f}}\in H$ where
${\widehat{f}}({\boldsymbol{x}})=\sum_{i=0}^{n-1}\omega_{i}K({\boldsymbol{x}},{\boldsymbol{x}}_{i})$
and ${\boldsymbol{\omega}}={\mathsf{K}}^{-1}{\boldsymbol{y}}$. The above kernel interpolant may be reinterpreted as the posterior mean in Gaussian process regression, see (Williams and Rasmussen, 2006). Fitting a Gaussian process regression model often includes optimizing a kernel’s hyperparameters, which also may be done by computing ${\mathsf{K}}{\boldsymbol{\omega}}$ and ${\mathsf{K}}^{-1}{\boldsymbol{y}}$.
Underlying these problems, and many others, is the requirement to compute the matrix-vector product ${\mathsf{K}}{\boldsymbol{y}}$ or solve the linear system ${\mathsf{K}}\boldsymbol{\omega}={\boldsymbol{y}}$ for some length $n$ vector ${\boldsymbol{y}}$. The standard cost of these computations are ${\mathcal{O}}(n^{2})$ and ${\mathcal{O}}(n^{3})$ respectively. One method to reduce these high costs is to induce structure into ${\mathsf{K}}$. Structure-inducing methods pairing lattices and digital nets with special kernels were proposed in (Zeng et al., 2009) and shown to reduce the cost of both computations to ${\mathcal{O}}(n\log n)$. Specifically, when a lattice point set $\{{\boldsymbol{x}}_{i}\}_{i=0}^{n-1}$ in linear order is paired with a shift-invariant kernel, ${\mathsf{K}}$ becomes circulant and computations can be performed performed using the fast Fourier transforms (FFT) and inverse FFT (IFFT). Similarly, when a base $2$ digital net $\{{\boldsymbol{x}}_{i}\}_{i=0}^{n-1}$ is paired with a digitally-shift-invariant kernel, ${\mathsf{K}}$ becomes nested-block-Toeplitz and these computations can be performed using the Fast Walsh-Hadamard Transforms (FWHT) (Fino and Algazi, 1976). Such methods have recently been used to accelerate Bayesian cubature (Jagadeeswaran and Hickernell, 2019, 2022; Rathinavel, 2019) and solving PDEs with random coefficients (Kaarnioja et al., 2022, 2023).
We provide novel methods which exploit past solutions to ${\mathsf{K}}{\boldsymbol{y}}$ or ${\mathsf{K}}^{-1}{\boldsymbol{y}}$ when doubling the sample size from the LD sequence. This requires lattices are generated in natural order so we do not need to reorder outputs after doubling the sample size. We show that the same speedups can be achieved for lattices in natural order using bit-reversed FFT and bit-reversed IFFT algorithms. These bit-reversed algorithms simply skip the initial and final bit-reversal steps in the decimation-in-time FFT and IFFT respectively. We provide implementations of the bit-reversed FFT, bit-reversed IFFT, and FWHT.
Shift-invariant kernels of arbitrary smoothness are well known and can be computed based on the Bernoulli polynomials (Kaarnioja et al., 2022, 2023; Cools et al., 2021, 2020; Sloan and Woźniakowski, 2001; Kuo et al., 2004). Until now, digitally-shift-invariant kernels were only known for order $1$ smoothness (Dick and Pillichshammer, 2005). We present higher-order digitally-shift-invariant kernels whose corresponding RKHSs contain functions of arbitrary smoothness. We derive closed forms of low order kernels (up to order $4$ smoothness) in base $b=2$, with the order $4$ kernel not appearing elsewhere in the literature. These kernels are also of interest for computing the worst-case error of higher-order polynomial-lattice rules (Baldeaux et al., 2012), although we do not explore this connection here.
1.3. Outline
The remainder of this article is organized as follows. First, we describe our implementation in section 2. Section 3 discusses notation. Section 4 details rank $1$ lattices, digital sequences, and Halton sequences along with their randomization routines. Section 5 describes accelerated kernel computation with LD sequences, matching kernels, and fast transform algorithms. Section 6 provides numerical experiments. Finally, section 7 gives a brief conclusion and outlines directions for future work.
2. Implementation
We have implemented this work into QMCPy [^10] https://qmcpy.readthedocs.io version 1.6, installable with pip install -U qmcpy. The architecture and philosophy underlying QMCPy are described in (Choi et al., 2022b, a). To follow along with code snippets presented in this article, please import QMCPy and NumPy (Oliphant, 2006) via
import qmcpy as qp
import numpy as np
These code snippets and code producing the plots in this article are available in notebook form [^11] https://github.com/QMCSoftware/QMCSoftware/blob/master/demos/ld_randomizations_and_higher_order_nets.ipynb.
Most of the QMCPy objects and methods presented here are wrappers around the lower level QMCToolsCL [^12] https://qmcsoftware.github.io/QMCToolsCL/ package. The Python functions in QMCToolsCL often call lower level C implementations for improved performance. QMCToolsCL interfaces for other languages which support C extensions is a direction of future work.
3. Notation
Bold symbols denote vectors which are assumed to be column vectors e.g. ${\boldsymbol{x}}\in[0,1]^{d}$. Capital letters in serif font denote matrices e.g. ${\mathsf{K}}\in{\mathbb{R}}^{n\times n}$. Lower case letters in serif font denote digits in a base $b$ expansion e.g. $i=\sum_{t=0}^{m-1}{\mathsf{i}}_{t}b^{t}$ for $0\leq i<2^{m}$. This may be combined with bold notation when denoting the vector of base $b$ digits e.g. ${\boldsymbol{\mathsf{i}}}=({\mathsf{i}}_{0},{\mathsf{i}}_{1},\dots,{\mathsf{i}}_{m-1})^{\intercal}$. Modulo is always taken component-wise e.g. ${\boldsymbol{x}}\mod 1=(x_{1}\mod 1,\dots,x_{d}\mod 1)^{\intercal}$. Permutations may be specified by vectors e.g. $\pi:\{0,1,2\}\to\{0,1,2\}$ with $\pi(0)=2$, $\pi(1)=0$, and $\pi(2)=1$ may be denoted by $\pi=(2,0,1)$.
4. Low-discrepancy sequences and randomization routines
For a fixed base $b\in{\mathbb{N}}$, write $i\in{\mathbb{N}}_{0}$ as $i=\sum_{t=0}^{\infty}{\mathsf{i}}_{t}b^{t}$ so ${\mathsf{i}}_{t}$ is the $t^{\mathrm{th}}$ digit in the base $b$ expansion of $i$. We denote the vector of the first $m$ base $b$ digits of $i$ by
${\boldsymbol{D}}_{m}(i)=({\mathsf{i}}_{0},{\mathsf{i}}_{1},\dots,{\mathsf{i}}_{m-1})^{\intercal}.$
For ${\boldsymbol{\mathsf{i}}}={\boldsymbol{D}}_{m}(i)$ we use the notation
$F_{m}({\boldsymbol{\mathsf{i}}})=\sum_{t=1}^{m}{\mathsf{i}}_{t-1}b^{-t}$
to flip the digits of $i$ about the decimal and convert back to a floating point number in $[0,1)$. Finally, let
(3) $v(i)=\sum_{t=1}^{\infty}{\mathsf{i}}_{t-1}b^{-t}$
so that $v(i)=F_{m}({\boldsymbol{D}}_{m}(i))$ when $i<b^{m}$, which is always the case in this article. The Van der Corput sequence in base $b$ is $(v(i))_{i\geq 0}$.
[…]
$x$$z$shifted digital net
$x$$z$permuted digital net
$x$$z$NUS digital net
Figure 2. Low-discrepancy randomization routines in dimension $d=1$. Each vertical line is a unit interval with 00 at the bottom and $1$ at the top. A given interval is partitioned at the horizontal ticks extending to the right, and then rearranged following the arrows to create the partition of the right interval as shown by the ticks extending to the left. For the shifted lattice and digitally shifted digital net, when the arrow hits a horizontal gray bar it is wrapped around to the next gray bar below. See for example the dotted line in the shifted digital net. The lattice shift is $\Delta=0.227$. All digital nets use base $b=3$ and $t_{\mathrm{max}}=2$ digits of precision. The digital shift is ${\boldsymbol{\Delta}}=(2,1)^{\intercal}$. Dropping the $j$ subscript for dimension, the digital permutations in the third panel are $\pi_{1}=(2,0,1)$ and $\pi_{2}=(2,1,0)$. Notice $\pi_{1}$ is equivalent to a digital shift by $2$, but $\pi_{2}$ cannot be written as a digital shift. The nested uniform scramble (NUS) has digital permutations $\pi=(2,0,1)$, $\pi_{0}=(2,1,0)$, $\pi_{1}=(0,1,2)$, and $\pi_{2}=(1,2,0)$. Notice the permuted digital net in third panel has permutations depending only the current digit in the base $b$ expansion. In contrast, the full NUS scrambling in the fourth panel has permutations which depend on all previous digits in the base $b$ expansion.
4.1. Rank 1 lattices
Consider a fixed generating vector ${\boldsymbol{g}}\in{\mathbb{N}}^{d}$ and fixed prime base $b$. Then we define the lattice sequence
${\boldsymbol{z}}_{i}=v(i){\boldsymbol{g}}\mod 1,\qquad i\geq 0.$
If $n=b^{m}$ for some $m\in{\mathbb{N}}_{0}$, then the lattice point set $\{{\boldsymbol{z}}_{i}\}_{i=0}^{n-1}\in[0,1)^{n\times d}$ is equivalent to $\{i/n{\boldsymbol{g}}\mod 1\}_{i=0}^{n-1}$ where we say the former is in natural order while the latter is in linear order.
Shifted lattice:
For a shift ${\boldsymbol{\Delta}}\in[0,1)^{d}$, we define the shifted point set
${\boldsymbol{x}}_{i}=({\boldsymbol{z}}_{i}+{\boldsymbol{\Delta}})\mod 1$
where ${\boldsymbol{z}}_{i}$ are lattice point set. The shift operation mod 1 is visualized in fig. 2. Randomized lattices use ${\boldsymbol{\Delta}}\sim{\mathcal{U}}[0,1]^{d}$. In the following code snippet we generate $R$ independently shifted lattices with shifts ${\boldsymbol{\Delta}}_{1},\dots,{\boldsymbol{\Delta}}_{R}\overset{\mathrm{IID}}{\sim}{\mathcal{U}}[0,1]^{d}$.
lattice = qp.Lattice(
dimension = 52,
randomize = "shift", # for unrandomized lattice set randomize = False
replications = 16, # R
order = "natural", # also supports "linear" and "Gray" for Gray code
seed = None, # pass integer seed for reproducibility
generating_vector = "kuo.lattice-33002-1024-1048576.9125.txt")
x = lattice(2**16) # a numpy.ndarray with shape 16 x 65536 x 52
Here we have used a generating vector from (Cools et al., 2006) which is stored in a standardized format in the LDData repository. Other generating vectors from LDData may be used by passing in a file name from https://github.com/QMCSoftware/LDData/tree/main/lattice/ or by passing an explicit array.
4.2. Digital nets
Consider a fixed prime base $b$ and generating matrices ${\mathsf{C}}_{1},\dots\mathsf{C}_{d}\in\{0,...,b-1\}^{t_{\mathrm{max}}\times m}$ where $m\in{\mathbb{N}}$ is fixed and $b$ is a given prime base. One may relax the assumption that $b$ is prime, but we do not consider that here or in the implementation. The first $n=b^{m}$ points of a digital sequence form a digital net $\{{\boldsymbol{z}}_{i}\}_{i=1}^{b^{m}-1}\in[0,1)^{b^{m}\times d}$ in natural order. For $1\leq j\leq d$ and $0\leq i<b^{m}$ with digits ${\boldsymbol{\mathsf{i}}}={\boldsymbol{D}}_{m}(i)$,
${\boldsymbol{\mathsf{z}}}_{ij}={\mathsf{C}}_{j}{\boldsymbol{\mathsf{i}}}\mod b\qquad\text{and}\qquad z_{ij}=F_{t_{\mathrm{max}}}({\boldsymbol{\mathsf{z}}}_{ij})$
where ${\boldsymbol{\mathsf{z}}}_{ij}$ is the base $b$ expansion vector of $z_{ij}$.
Digitally shifted digital net:
Similar to lattices, one may apply a shift ${\mathsf{\Delta}}\in[0,1)^{t_{\mathrm{max}}\times d}$ to the digital net to get a digitally-shifted digital net $\{{\boldsymbol{x}}_{i}\}_{i=0}^{b^{m}-1}$ where
$x_{ij}=F_{t_{\mathrm{max}}}(({\boldsymbol{\mathsf{z}}}_{ij}+{\boldsymbol{\Delta}}_{j})\mod b)$
and ${\boldsymbol{\Delta}}_{j}$ is the $j^{\mathrm{th}}$ column of ${\mathsf{\Delta}}$. Randomly shifted digital nets use ${\mathsf{\Delta}}\overset{\mathrm{IID}}{\sim}{\mathcal{U}}\{0,\dots,b-1\}^{t_{\mathrm{max}}\times d}$ i.e. is each digit is chosen uniformly from $\{0,\dots,b-1\}$.
Digitally permuted digital net:
In what follows will denote permutations of $\{0,\dots,b-1\}$ by $\pi$. Suppose we are given a set of permutations
${\mathsf{\Pi}}=\{\pi_{j,t}:\quad 1\leq j\leq d,\quad 0\leq t<t_{\mathrm{max}}\}.$
Then we may construct the digitally permuted digital net $\{{\boldsymbol{x}}_{i}\}_{i=1}^{b^{m}-1}$ where
$x_{ij}=F_{t_{\mathrm{max}}}((\pi_{j,0}({\mathsf{z}}_{ij0}),\pi_{j,1}({\mathsf{z}}_{ij1}),\dots,\pi_{j,t_{\mathrm{max}}-1}({\mathsf{z}}_{ij(t_{\mathrm{max}}-1)}))^{\intercal}).$
Randomly permuted digital nets use independent permutations chosen uniformly over all permutations of $\{0,\dots,b-1\}$.
Nested Uniform Scrambling (NUS):
NUS is often called Owen scrambling for its conception in (Owen, 1995). As before, $\pi$ values denote permutations of $\{0,\dots,b-1\}$. Now suppose we are given a set of permutations
${\mathcal{P}}=\{\pi_{j,{\mathsf{v}}_{1}\cdots{\mathsf{v}}_{t}}:\quad 1\leq j\leq d,\quad 0\leq t<t_{\mathrm{max}},\quad{\mathsf{v}}_{k}\in\{0,\dots,b-1\}\text{ for }0\leq k\leq t\}.$
Then a nested uniform scrambling of a digital net is $\{{\boldsymbol{x}}_{i}\}_{i=1}^{b^{m}-1}$ where
$x_{ij}=F_{t_{\mathrm{max}}}((\pi_{j,}({\mathsf{z}}_{ij0}),\pi_{j,{\mathsf{z}}_{ij0}}({\mathsf{z}}_{ij1}),\pi_{j,{\mathsf{z}}_{ij0}{\mathsf{z}}_{ij1}}({\mathsf{z}}_{ij2}),\dots,\pi_{j,{\mathsf{z}}_{ij0}{\mathsf{z}}_{ij1}\cdots{\mathsf{z}}_{ij(t_{\mathrm{max}}-2)}}({\mathsf{z}}_{ij(t_{\mathrm{max}}-1)}))^{\intercal}).$
As the number of elements in ${\mathcal{P}}$ is
$\lvert{\mathcal{P}}\rvert=d(1+b+b^{2}+\dots+b^{t_{\mathrm{max}}-1})=d\frac{b^{t_{\mathrm{max}}}-1}{b-1},$
our implementation cannot afford to generated all permutations apriori. Instead, permutations are generated and stored only as needed. As with digitally-permuted digital nets, NUS uses independent uniform random permutations.
The randomization routines described above are visualized in fig. 2.
Linear matrix scrambling (LMS) is a computationally cheaper but less complete version of NUS which has proven sufficient for many practical problems. LMS uses scrambling matrices ${\mathsf{S}}_{1},\dots,{\mathsf{S}}_{d}\in\{0,\dots,b\}^{t_{\mathrm{max}}\times t_{\mathrm{max}}}$ and sets the LMS generating matrices ${\widetilde{\mathsf{C}}}_{1},\dots,{\widetilde{\mathsf{C}}}_{d}\in\{0,\dots,b\}^{t_{\mathrm{max}}\times m}$ so that
${\widetilde{\mathsf{C}}}_{j}={\mathsf{S}}_{j}{\mathsf{C}}_{j}\mod b$
for $j=1,\dots,d$. Following (Owen, 2003), let us denote elements in $\{1,\dots,b\}$ by $h$ and elements in $\{0,\dots,b\}$ by $g$. Then common structures for ${\mathsf{S}}_{j}$ include
$\begin{pmatrix}h_{11}&0&0&0&\dots\\ g_{21}&h_{22}&0&0&\dots\\ g_{31}&g_{32}&h_{33}&0&\dots\\ g_{41}&g_{42}&g_{43}&h_{44}&\dots\\ \vdots&\vdots&\vdots&\vdots&\ddots\end{pmatrix},\begin{pmatrix}h_{1}&0&0&0&\dots\\ g_{2}&h_{1}&0&0&\dots\\ g_{3}&g_{2}&h_{1}&0&\dots\\ g_{4}&g_{3}&g_{2}&h_{1}&\dots\\ \vdots&\vdots&\vdots&\vdots&\ddots\end{pmatrix},\quad\text{and}\quad\begin{pmatrix}h_{1}&0&0&0&\dots\\ h_{1}&h_{2}&0&0&\dots\\ h_{1}&h_{2}&h_{3}&0&\dots\\ h_{1}&h_{2}&h_{3}&h_{4}&\dots\\ \vdots&\vdots&\vdots&\vdots&\ddots\end{pmatrix}$
which corresponds to Matousek’s linear scrambling (Matoušek, 1998), Tezuka’s $i$-binomial scrambling (Tezuka, 2002), and Owen’s striped LMS (Owen, 2003) (not to be confused with NUS which is often called Owen scrambling). Random LMS chooses $g$ and $h$ values all independently and uniformly from $\{1,\dots,b\}$ and $\{0,\dots,b-1\}$ respectively. These scrambling are described in more detail and connected to NUS in (Owen, 2003).
Digital interlacing enables the construction of higher-order digital nets. Denote the digitally interlaced version of ${\mathsf{C}}_{1},{\mathsf{C}}_{2},\dots\in\{0,\dots,b\}^{t_{\mathrm{max}}\times m}$ with interlacing factor $\alpha\in{\mathbb{N}}$ by ${\widetilde{\mathsf{C}}}_{1},{\widetilde{\mathsf{C}}}_{2},\dots\in\{0,\dots,b\}^{\alpha t_{\mathrm{max}}\times m}$. Then we have ${\widetilde{\mathsf{C}}}_{jtk}=C_{{\widetilde{j}},\tt,k}$ where ${\widetilde{j}}=\alpha(j-1)+(t\mod\alpha)+1$ and $\tt=\lfloor t/\alpha\rfloor$ for $j\geq 1$ and $0\leq t<\alpha t_{\mathrm{max}}$ and $1\leq k\leq m$. For example, with $m=2$, $t_{\mathrm{max}}=2$, and $\alpha=2$ we have the following
${\mathsf{C}}_{1}=\begin{pmatrix}c_{101}&c_{102}\\ c_{111}&c_{112}\end{pmatrix}\qquad{\mathsf{C}}_{2}=\begin{pmatrix}c_{201}&c_{202}\\ c_{211}&c_{212}\end{pmatrix}\qquad{\mathsf{C}}_{3}=\begin{pmatrix}c_{301}&c_{302}\\ c_{311}&c_{312}\end{pmatrix}\qquad{\mathsf{C}}_{4}=\begin{pmatrix}c_{401}&c_{402}\\ c_{411}&c_{412}\end{pmatrix}\qquad\dots$
${\widetilde{\mathsf{C}}}_{1}=\begin{pmatrix}c_{101}&c_{102}\\ c_{201}&c_{202}\\ c_{111}&c_{112}\\ c_{211}&c_{212}\end{pmatrix}\qquad{\widetilde{\mathsf{C}}}_{2}=\begin{pmatrix}c_{301}&c_{302}\\ c_{401}&c_{402}\\ c_{311}&c_{312}\\ c_{411}&c_{412}\end{pmatrix}\qquad\dots$
Higher-order NUS requires NUS be applied to the digital net point sets generated by ${\mathsf{C}}_{1},\dots,{\mathsf{C}}_{\alpha d}$ and then interlacing be performed to the digits $\{{\boldsymbol{\mathsf{z}}}_{i1}\}_{i=0}^{b^{m}-1},\dots,\{{\boldsymbol{\mathsf{z}}}_{i,\alpha d}\}_{i=0}^{b^{m}-1}$ (Dick, 2011). For higher-order LMS, we apply LMS to the generating matrices ${\mathsf{C}}_{1},\dots,{\mathsf{C}}_{\alpha d}$, then interlace them, then generate the digital net. As we shown in the numerical experiments in section 6, LMS is significantly faster than NUS (especially for higher-order nets) while still achieving higher-order rates of RMSE convergence.
A subtle difference between the above presentation and practical implementation is that $t_{\mathrm{max}}$ may change with randomization in practice. For example, suppose we are given generating matrices ${\mathsf{C}}_{1},\dots,{\mathsf{C}}_{d}\in\{0,\dots,b-1\}^{32\times 32}$ but would like the shifted digital net to have $64$ digits of precision. Then we should generate ${\boldsymbol{\Delta}}\in\{0,\dots,b-1\}^{t_{\mathrm{max}}\times d}$ with $t_{\mathrm{max}}=64$ and treat ${\mathsf{C}}_{j}$ as $t_{\mathrm{max}}\times t_{\mathrm{max}}$ matrices with appropriate rows and columns zeroed out.
Gray code ordering of digital nets enables computing the next point ${\boldsymbol{x}}_{i+1}$ from ${\boldsymbol{x}}_{i}$ by only adding a single column from each generating matrix rather. Specifically, the $q^{\mathrm{th}}$ column of each generating matrix gets digitally added to the previous point where $q-1$ is the index of the only digit to be changed in Gray code ordering. Gray code orderings for $b=2$ and $b=3$ are shown in table 1.
$i$ $i_{2}$ Gray code $i_{2}$ $i_{3}$ Gray code $i_{3}$
00 $0000_{2}$ $0000_{2}$ $00_{3}$ $00_{3}$
$1$ $0001_{2}$ $0001_{2}$ $01_{3}$ $01_{3}$
$2$ $0010_{2}$ $0011_{2}$ $02_{3}$ $02_{3}$
$3$ $0011_{2}$ $0010_{2}$ $10_{3}$ $12_{3}$
$4$ $0100_{2}$ $0110_{2}$ $11_{3}$ $11_{3}$
$5$ $0101_{2}$ $0111_{2}$ $12_{3}$ $10_{3}$
$6$ $0110_{2}$ $0101_{2}$ $20_{3}$ $20_{3}$
$7$ $0111_{2}$ $0100_{2}$ $21_{3}$ $21_{3}$
$8$ $1000_{2}$ $1000_{2}$ $22_{3}$ $22_{3}$
Table 1. Gray code order for bases $b=2$ and $b=3$. In Gray code order only one digit is incremented or decremented by $1$ (modulo $b$) when $i$ is incremented by $1$.
The following code generates a base $2$ digital net with $R$ independent LMS, $R$ independent digital shifts, and $\alpha=2$ interlacing.
dnb2 = qp.DigitalNetB2(
dimension = 52,
randomize = "LMS_DS", # Matousek’s LMS with a digital shift
# options are [LMS_DS, NUS, DS, LMS, False]
t_lms = 64, # number of LMS bits i.e. number of rows in S_j
alpha = 2, # interlacing factor for higher order digital nets
replications = 15, # R
graycode = False, # use natural order instead of Gray code order
seed = None, # pass integer seed for reproducibility
generating_matrices = "LDData:mps.sobol_Cs.txt")
x = dnb2(2**16) # a numpy.ndarray with shape 15 x 65536 x 52
Here we have used a set of Sobol’ generating matrices from Joe and Kuo131313the “new-joe-kuo-6.21201” direction numbers from https://web.maths.unsw.edu.au/~fkuo/sobol/index.html (Joe and Kuo, 2008) which are stored in a standardized format in the LDData repository. Other generating matrices from LDData may be used by passing in a file name from https://github.com/QMCSoftware/LDData/blob/main/dnet/ or by passing an explicit array.
4.3. Halton sequences
The digital sequences described in the previous section used a fix prime base $b$. One may allow each dimension $j\in\{1,\dots,d\}$ to have its own prime base $b_{j}$. The most popular of such constructions is the Halton sequence which sets $b_{j}$ to the $j^{\mathrm{th}}$ prime, sets ${\mathsf{C}}_{j}$ to the identity matrix, and sets $t_{\mathrm{max}}=m$. This enables the simplified construction of Halton points $\{{\boldsymbol{x}}_{i}\}_{i=0}^{n-1}$ via
${\boldsymbol{x}}_{i}=(v_{b_{1}}(i),\dots,v_{b_{d}}(i))^{\intercal}$
where we have added a subscript to $v$ in (3) to denote the base dependence.
Almost all the methods described for digital sequences are immediately applicable to Halton sequences after accounting for the differing bases across dimensions. However, digital interlacing is not generally applicable when the bases differ. Halton with random starting points has also been explored in (Wang and Hickernell, 2000), although we do not consider this here. The following code generates a Halton point set with $R$ independent LMS and $R$ independent digital permutations.
halton = qp.Halton(
dimension = 52,
randomize = "LMS_PERM", # Matousek’s LMS with a digital permutation
# options are [LMS_PERM, LMS_DS, LMS, PERM, DS, NUS, QRNG, FALSE]
t_lms = 64, # number of LMS digits i.e. number of rows in S_j
replications = 15, # R
seed = None) # pass integer seed for reproducibility
x = halton(2**10) # a numpy.ndarray with shape 15 x 1024 x 52
The QRNG randomization follows the QRNG software package (Hofert and Lemieux, 2023) in generating a generalized Halton point set (Faure and Lemieux, 2009) using a deterministic set of permutation scrambles and random digital shifts.
5. Fast transforms and kernel computations
Recall from the introduction that we would like to compute the matrix-vector product ${\mathsf{K}}{\boldsymbol{y}}$ and solve the linear system ${\mathsf{K}}^{-1}{\boldsymbol{y}}$ where ${\boldsymbol{y}}$ is some known length $n$ vector and ${\mathsf{K}}=\{K({\boldsymbol{x}}_{i},{\boldsymbol{x}}_{k})\}_{i,k=0}^{n-1}$ is a $n\times n$ symmetric positive definite (SPD) Gram matrix based on an SPD kernel $K$ and point set $\{{\boldsymbol{x}}_{i}\}_{i=0}^{n-1}$. We may reduce the standard ${\mathcal{O}}(n^{2})$ and ${\mathcal{O}}(n^{3})$ costs of solving these respective problems to both be ${\mathcal{O}}(n\log n)$ by inducing structure into ${\mathsf{K}}$ using special point set-kernel pairings. Specifically, pairing lattice points with a shift invariance kernel yields a circulant Gram matrix ${\mathsf{K}}$ for which computations can be done quickly using the bit-reversed fast Fourier transform (FFT) and bit-reversed inverse FFT (IFFT). Similarly, pairing a digital net with digitally-shift-invariant kernel yields a nested block Toeplitz matrix for which computations can be done quickly using the fast Walsh Hadamard Transform (FWHT). Shift-invariant and digitally-shift-invariant kernels are shown in fig. 3. The reduced costs achieved by inducing these structures are summarized in table 2.
Figure 3. Shift-invariant kernels (top row) and digitally-shift-invariant kernels (bottom row). Here $\alpha$ denotes the smoothness of the kernel. The $\alpha=3$ and $\alpha=4$ digitally-shift-invariant kernels are nearly indistinguishable. Shift-invariant kernels are described in section 5.1 while digitally-shift-invariant kernels are described in section 5.2.
[…]
$\{{\boldsymbol{x}}_{i}\}_{i=0}^{n-1}$ $K$ structure ${\mathsf{K}}$ storage ${\mathsf{K}}$ decomposition ${\mathsf{K}}{\boldsymbol{y}}$ cost ${\mathsf{K}}^{-1}{\boldsymbol{y}}$ cost methods
general SPD ${\mathcal{O}}(n^{2})$ ${\mathcal{O}}(n^{3})$ ${\mathcal{O}}(n^{2})$ ${\mathcal{O}}(n^{2})$ standard
lattice SPD SI ${\mathcal{O}}(n)$ ${\mathcal{O}}(n\log n)$ ${\mathcal{O}}(n\log n)$ ${\mathcal{O}}(n\log n)$ FFT-based
digital net SPD DSI ${\mathcal{O}}(n\log n)$ ${\mathcal{O}}(n)$ ${\mathcal{O}}(n\log n)$ ${\mathcal{O}}(n\log n)$ FWHT-based
Table 2. Comparison of storage and cost requirements for kernel methods. Decomposition of the symmetric positive definite (SPD) Gram matrix ${\mathsf{K}}$ is the cost of computing the eigendecomposition or Cholesky factorization. The costs of matrix-vector multiplication and solving a linear system are the costs after performing the decomposition. Both storage and kernel computation costs are greatly reduced by pairing certain low-discrepancy point sets with special shift-invariant (SI) or digitally-shift-invariant (DSI) kernels.
The following code can be used to construct a Gram matrix with a lattice point set and shift-invariant kernel. The Gram matrix is then used perform matrix-vector multiplication and solve a linear system. Analogous code using qp.DigitalNetB2, qp.KernelDigShiftInvar, and qp.FastGramMatrixDigitalNetB2 is also supported.
d = 5 # dimension
n = 2**20 # 2^20 points
dd_obj = qp.Lattice(dimension=d, seed=7)
# qp.Lattice defaults to natural ordering with a single random shift
kernel_obj = qp.KernelShiftInvar( # matching kernel
dimension = d,
lengthscales = 0.1*np.ones(d,dtype=float), # product weights
alpha = 4*np.ones(d,dtype=int)) # smoothness parameters
K = qp.FastGramMatrixLattice( # n x n Gram matrix
kernel_obj = kernel_obj,
dd_obj = dd_obj,
n1 = n,
n2 = n)
y = np.random.rand(n) # get a random length n vector
a = K@y # compute the matrix-vector product a=Ky at O(n log n) cost
b = K.solve(y) # solve linear system Kb=y for b at O(n log n) cost
Schematics of the bit-reversed FFT and FWHT underlying these fast computations are shown in fig. 4. The one-dimensional bit-reversed FFT, one-dimensional bit-reversed IFFT, and one-dimensional FWHT are implemented into the QMCPy functions qp.fftbr, qp.ifftbr, and qp.fwht respectively. These fast transforms act along the last dimension which is required to to have length $2^{m}$ for some $m\geq 0$. The following subsections detail the matching point set-kernel pairings and provide methods to quickly update the Gram matrix product or linear system solution after doubling the point size.
[…]
Figure 4. Bit-reversed FFT and FWHT schemes. The bit-reversed FFT is performed via a decimation in time algorithm without the initial reordering of inputs. The bit-reversed IFFT may be performed by propagating $\{{\widetilde{y}}_{i}\}_{i=0}^{2^{m}-1}$ to $\{y_{i}\}_{i=0}^{2^{m}-1}$ in the other direction through the bit-reversed FFT algorithm.
[…]
5.1. Shift invariant kernels, lattices, and the bit-reversed FFT/IFFT
A kernel is said to be shift-invariant when $K({\boldsymbol{u}},{\boldsymbol{v}})={\widetilde{K}}(({\boldsymbol{u}}-{\boldsymbol{v}})\mod 1)$ for some ${\widetilde{K}}$ i.e. the kernel is only a function of the component-wise difference between inputs modulo $1$. One set of shift-invariant kernels take the form
(4) $K({\boldsymbol{u}},{\boldsymbol{v}})=\sum_{{\mathfrak{u}}\subseteq{\mathcal{D}}}\gamma_{\mathfrak{u}}\prod_{j\in{\mathfrak{u}}}\eta_{\alpha_{j}}((u_{j}-v_{j})\mod 1)$
where
$\eta_{\alpha}(\delta)=\frac{(2\pi)^{2\alpha}}{(-1)^{\alpha+1}(2\alpha)!}B_{2\alpha}(\delta)$
with ${\boldsymbol{\alpha}}\in{\mathbb{N}}^{d}$, ${\mathcal{D}}=\{1,\dots,d\}$, weights $\{\gamma_{\mathfrak{u}}\}_{{\mathfrak{u}}\subseteq{\mathcal{D}}}$, and $B_{\ell}$ denoting the Bernoulli polynomial of degree $\ell$. The corresponding reproducing kernel Hilbert space (RKHS) $H$ is a weighted periodic unanchored Sobolev space of smoothness ${\boldsymbol{\alpha}}\in{\mathbb{N}}^{d}$ with norm
$\lVert f\rVert_{H}^{2}:=\sum_{{\mathfrak{u}}\subseteq{\mathcal{D}}}\frac{1}{(2\pi)^{2\lvert{\mathfrak{u}}\rvert}\gamma_{\mathfrak{u}}}\int_{[0,1]^{\lvert{\mathfrak{u}}\rvert}}\left\lvert\int_{[0,1]^{s-\lvert{\mathfrak{u}}\rvert}}\left(\prod_{j\in{\mathfrak{u}}}\frac{\partial^{\alpha_{j}}}{\partial y_{j}^{\alpha_{j}}}\right)f({\boldsymbol{y}})\mathrm{d}{\boldsymbol{y}}_{-{\mathfrak{u}}}\right\rvert^{2}\mathrm{d}{\boldsymbol{y}}_{\mathfrak{u}},$
where $f:[0,1]^{d}\to{\mathbb{R}}$, ${\boldsymbol{y}}_{\mathfrak{u}}=({\boldsymbol{y}}_{j})_{j\in{\mathfrak{u}}}$, ${\boldsymbol{y}}_{-{\mathfrak{u}}}:=(y_{j})_{j\in{\mathcal{D}}\setminus{\mathfrak{u}}}$, and $\lvert{\mathfrak{u}}\rvert$ is the cardinality of ${\mathfrak{u}}$. The space $H$ is a special case of the weighted Korobov space which has real smoothness parameter $\alpha$ characterizing the rate of decay of Fourier coefficients, see e.g. (Kaarnioja et al., 2022, 2023; Cools et al., 2021, 2020; Sloan and Woźniakowski, 2001; Kuo et al., 2004).
The kernel (4) is the sum over $2^{d}$ terms and thus becomes impractical to compute for large $d$. A simplified space assumes $\gamma_{\mathfrak{u}}$ are product weights which take the form $\gamma_{\mathfrak{u}}=\prod_{j\in{\mathfrak{u}}}\gamma_{j}$ for some $\{\gamma_{j}\}_{j=1}^{d}$. The kernel with product weights takes the simplified form
(5) $K({\boldsymbol{u}},{\boldsymbol{v}})=\prod_{j=1}^{d}\left(1+\gamma_{j}\eta_{\alpha_{j}}((u_{j}-v_{j})\mod 1)\right).$
Figure 3 shows one dimensional shift-invariant kernels with unit weight and varying smoothness. See (Kaarnioja et al., 2023) for a review of lattice-based kernel interpolation in such weighted spaces along with the development of serendipitous weights, which are a special form of high-performance product weights.
Suppose $n=2^{m}$ for some $m\in{\mathbb{N}}_{0}$ and let $\{{\boldsymbol{x}}_{i}\}_{i=0}^{b^{m}-1}$ be a shifted lattice with shift ${\boldsymbol{\Delta}}\in[0,1)^{d}$ generated in linear order. For $0\leq i,k<n$
$K({\boldsymbol{x}}_{i},{\boldsymbol{x}}_{k})=K\left(\frac{(i-k){\boldsymbol{g}}}{n},{\boldsymbol{0}}\right),$
so the Gram matrix ${\mathsf{K}}$ is circulant. Let $R_{m}(i)$ flip the first $m$ bits of $0\leq i<2^{m}$ in base $b=2$ so that if $i=\sum_{t=0}^{m-1}{\mathsf{i}}_{t}2^{t}$ then $R(i)=\sum_{t=0}^{m-1}{\mathsf{i}}_{m-t-1}2^{t}$. The first step in the FFT with decimation-in-time is to reorder the inputs $\{y_{i}\}_{i=0}^{2^{m}-1}$ into bit reversed order $\{y_{R(i)}\}_{i=0}^{2^{m}-1}$. The last step in computing the IFFT is to reorder the outputs $\{y_{R(i)}\}_{i=0}^{2^{m}-1}$ into bit reversed order $\{y_{i}\}_{i=0}^{2^{m}-1}$. Therefore, one should generate lattice points in natural order and skip the first step of the FFT and last step of the inverse FFT.
The Fourier matrix is $\overline{{\mathsf{F}}^{(m)}}=\{W_{m}^{ij}\}_{i,j=0}^{2^{m}-1}$ where $W_{m}=\exp(-2\pi\sqrt{-1}/2^{m})$. Let $\overline{{\mathsf{T}}^{(m)}}=\{W_{m}^{iR_{m}(j)}\}_{i,j=0}^{2^{m}-1}$ so that
$\overline{{\mathsf{F}}^{(m)}\{y_{i}\}_{j=0}^{2^{m}-1}}=\overline{{\mathsf{T}}^{(m)}\{y_{R(j)}\}_{j=0}^{2^{m}-1}}.$
For $\{{\boldsymbol{x}}_{i}\}_{i=0}^{2^{m}-1}$ a lattice in natural order, we have
${\mathsf{K}}=\frac{1}{n}{\mathsf{T}}^{(m)}{\mathsf{\Lambda}}^{(m)}\overline{{\mathsf{T}}^{(m)}}.$
Notice that $\overline{{\mathsf{T}}^{(m+1)}}=\left\{W_{m+1}^{iR_{m+1}(j)}\right\}_{i,j=0}^{2^{m+1}-1}$. For $0\leq j<2^{m}$ we have $R_{m+1}(j)=2R_{m}(j)$ so $W^{iR_{m+1}(j)}_{m+1}=W^{iR_{m}(j)}_{m}$. For $2^{m}\leq j<2^{m+1}$ we have $R_{m+1}(j)=2R_{m}(j-2^{m})+1$ so $W^{iR_{m+1}(j)}_{m+1}=W_{m}^{iR(j-2^{m})}W_{m+1}^{i}$. Moreover, for $0\leq i<2^{m}$ we have $W_{m+1}^{2^{m}+i}=-W_{m+1}^{i}$. Define ${\widetilde{\boldsymbol{w}}}^{(m)}=\{W_{m+1}^{i}\}_{i=0}^{2^{m}-1}$. Then
$\overline{{\mathsf{T}}^{(m+1)}}=\begin{pmatrix}\overline{{\mathsf{T}}^{(m)}}&\mathrm{diag}({\widetilde{\boldsymbol{w}}}^{(m)})\overline{{\mathsf{T}}^{(m)}}\\ \overline{{\mathsf{T}}^{(m)}}&-\mathrm{diag}({\widetilde{\boldsymbol{w}}}^{(m)})\overline{{\mathsf{T}}^{(m)}}\end{pmatrix}$
and ${\widetilde{\boldsymbol{w}}}^{(m+1)}$ is the $\alpha=2$ interlacing of ${\widetilde{\boldsymbol{w}}}^{(m)}$ and $W_{m+1}{\widetilde{\boldsymbol{w}}}^{(m)}$.
5.2. Digital nets, digitally-shift-invariant kernels, and the FWHT
For $x,y\in[0,1)$ with $x=\sum_{t=1}^{m}{\mathsf{x}}_{t-1}b^{-t}$ and $y=\sum_{t=1}^{m}{\mathsf{y}}_{t-1}b^{-t}$ let
$x\oplus y=\sum_{t=1}^{m}(({\mathsf{x}}_{t-1}+{\mathsf{y}}_{t-1})\mod b)b^{-t}\qquad\text{and}\qquad x\ominus y=\sum_{t=1}^{m}(({\mathsf{x}}_{t-1}-{\mathsf{y}}_{t-1})\mod b)b^{-t}$
denote digital addition and subtraction respectively. Note that in base $b=2$ these operations are both equivalent to the exclusive or (XOR) between bits. For vectors ${\boldsymbol{x}},{\boldsymbol{y}}\in[0,1)^{d}$, digital operations ${\boldsymbol{x}}\oplus{\boldsymbol{y}}$ and ${\boldsymbol{x}}\ominus{\boldsymbol{y}}$ act component-wise. A kernel is said to be digitally-shift-invariant when
$K({\boldsymbol{u}},{\boldsymbol{v}})={\widetilde{K}}({\boldsymbol{u}}\ominus{\boldsymbol{v}})$
for some ${\widetilde{K}}$.
Suppose $n=b^{m}$ for some $m\in{\mathbb{N}}_{0}$ and let $\{{\boldsymbol{x}}_{i}\}_{i=0}^{b^{m}-1}$ be a digitally shifted digital net in natural order (possibly of higher order and/or with LMS). Following (Jagadeeswaran and Hickernell, 2022, Theorem 5.3.1,Theorem 5.3.2), for $b=2$, the Gram matrix ${\mathsf{K}}$ is nested block Toeplitz which implies the eigendecomposition
${\mathsf{K}}=\frac{1}{n}{\mathsf{H}}^{(m)}{\mathsf{\Lambda}}^{(m)}{\mathsf{H}}^{(m)}$
where ${\mathsf{H}}^{(m)}$ is the $2^{m}\times 2^{m}$ Hadamard matrix defined by ${\mathsf{H}}^{(0)}=(1)$ and the relationship
${\mathsf{H}}^{(m+1)}=\begin{pmatrix}{\mathsf{H}}^{(m)}&{\mathsf{H}}^{(m)}\\ {\mathsf{H}}^{(m)}&-{\mathsf{H}}^{(m)}\end{pmatrix}.$
Multiplying by ${\mathsf{H}}^{(m)}$ can be done at at ${\mathcal{O}}(n\log n)$ cost using the FWHT. Note that ${\mathsf{K}}$ is not necessarily nested block Toeplitz under NUS.
We now describe one dimensional digitally-shift-invariant kernels $K(u,v)={\widetilde{K}}(u\ominus v)$ of higher-order smoothness $\alpha$. Weighted sums over products these one dimensional kernels may be used to construct higher-dimensional kernels e.g. those with product weights
$K({\boldsymbol{u}},{\boldsymbol{v}})=\prod_{j=1}^{d}\left(1+\gamma_{j}\left(K(u_{j},v_{j})-1\right)\right).$
To begin, let us write $k\in{\mathbb{N}}_{0}$ as $k=\sum_{\ell=1}^{\#k}{\mathsf{k}}_{a_{\ell}}b^{a_{\ell}}$ where $a_{1}>\cdots>a_{\#k}\geq 0$ are the #$\#k$ indices of non-zero digits ${\mathsf{k}}_{a_{\ell}}\in\{1,\dots,b-1\}$ in the base $b$ expansion of $k$. Then the $k^{\mathrm{th}}$ Walsh coefficient is
$\mathrm{wal}_{k}(x)=e^{2\pi\sqrt{-1}/b\sum_{\ell=1}^{\#k}{\mathsf{k}}_{a_{\ell}}{\mathsf{x}}_{a_{\ell}+1}}.$
For $\alpha\in{\mathbb{N}}$ we also define the weight function
$\mu_{\alpha}(k)=\sum_{\ell=1}^{\min\{\alpha,\#k\}}(a_{\ell}+1).$
For $\alpha>1$, let the digitally-shift-invariant kernel
(6) $K(x,z)=\sum_{k\in{\mathbb{N}}_{0}}\frac{\mathrm{wal}_{k}(x\ominus z)}{b^{\mu_{\alpha}(k)}}={\widetilde{K}}_{\alpha}(x\ominus z)$
have corresponding RKHS ${\widetilde{H}}_{\alpha}$. The following theorem shows ${\widetilde{H}}_{\alpha}$ contains smooth functions. The case when $\alpha=1$ is treated separately in (Dick and Pillichshammer, 2005).
Theorem 5.1.
${\widetilde{H}}_{\alpha}\supset H_{\alpha}$ where $H_{\alpha}$ is an RKHS with inner product
$\langle f,g\rangle_{H_{\alpha}}=\sum_{l=1}^{\alpha-1}\int_{0}^{1}f^{(l)}(x)\mathrm{d}x\int_{0}^{1}g^{(l)}(x)\mathrm{d}x+\int_{0}^{1}f^{(\alpha)}(x)g^{(\alpha)}(x)\mathrm{d}x.$
Proof.
See appendix A
The following theorem gives explicit forms of a few higher-order digitally-shift-invariant kernels in base $b=2$ which are plotted in fig. 3. These kernels are also useful in the computation of the worst-case error of QMC with higher order polynomial-lattice rules as in (Baldeaux et al., 2012). In fact, their paper details how to compute ${\widetilde{K}}_{\alpha}$ at ${\mathcal{O}}(\alpha\#x)$ cost where #$\#x$ is the number of non-zero digits in the base $b$ expansion of $x$
Theorem 5.2.
Fix the base $b=2$. Let $\beta(x)=-\lfloor\log_{2}(x)\rfloor$ and $t_{\nu}(x)=2^{-\nu\beta(x)}$ where for $x=0$ we set $\beta(x)=t_{\nu}(x)=0$. Then
${\widetilde{K}}_{2}(x)$ $=-\beta(x)x+\frac{5}{2}\left[1-t_{1}(x)\right],$
${\widetilde{K}}_{3}(x)$ $=\beta(x)x^{2}-5\left[1-t_{1}(x)\right]x+\frac{43}{18}\left[1-t_{2}(x)\right],$
${\widetilde{K}}_{4}(x)$ $=-\frac{2}{3}\beta(x)x^{3}+5\left[1-t_{1}(x)\right]x^{2}-\frac{43}{9}\left[1-t_{2}(x)\right]x$
$+\frac{701}{294}\left[1-t_{3}(x)\right]+\beta(x)\left[\frac{1}{48}\sum_{a=0}^{\infty}\frac{\mathrm{wal}_{2^{a}}(x)}{2^{3a}}-\frac{1}{42}\right].$
Proof.
The $\alpha\in\{2,3\}$ forms are due to (Baldeaux et al., 2012). The $\alpha=4$ form is derived in appendix A. ∎
5.3. Fast kernel computations and efficient updates
Let us make the general assumptions satisfied by the formulations in the previous two subsections.
  1. (1)
    For $n=2^{m}$, the Gram matrix ${\mathsf{K}}^{(m)}\in{\mathbb{R}}^{n\times n}$ can be written as
    ${\mathsf{K}}^{m}={\mathsf{V}}^{(m)}{\mathsf{\Lambda}}^{(m)}\overline{{\mathsf{V}}^{(m)}}$
    where ${\mathsf{V}}^{(m)}\overline{{\mathsf{V}}^{(m)}}={\mathsf{I}}$ and the first column of ${\mathsf{V}}^{(m)}$, denoted ${\boldsymbol{v}}_{1}^{(m)}$, is the constant $1/\sqrt{n}$
  2. (2)
    $V^{(m)}\in{\mathbb{C}}^{n\times n}$ satisfies ${\mathsf{V}}^{(0)}=(1)$ and
    ${\mathsf{V}}^{(m+1)}=\begin{pmatrix}{\mathsf{V}}^{(m)}&{\mathsf{V}}^{(m)}\\ {\mathsf{V}}^{(m)}\overline{\mathrm{diag}({\widetilde{\boldsymbol{w}}}^{(m)})}&-{\mathsf{V}}^{(m)}\overline{\mathrm{diag}({\widetilde{\boldsymbol{w}}}^{(m)})}\end{pmatrix}/\sqrt{2}.$
  3. (3)
    ${\mathsf{V}}^{(m)}{\boldsymbol{y}}$ and $\overline{{\mathsf{V}}^{(m)}}{\boldsymbol{y}}$ are each computable at cost ${\mathcal{O}}(2^{m}m)$.
For the case of lattice points with shift-invariant kernels and resulting circulant Gram matrices, ${\mathsf{V}}^{(m)}={\mathsf{T}}^{(m)}/\sqrt{2^{m}}$ is the scaled and permuted Fourier matrix and ${\widetilde{\boldsymbol{w}}}^{(m)}=\{\exp(-\pi\sqrt{-1}/2^{m}i)\}_{i=0}^{2^{m}-1}$, so ${\mathsf{V}}^{(m)}{\boldsymbol{y}}$ and $\overline{{\mathsf{V}}^{(m)}}{\boldsymbol{y}}$ can be computed using a bit-reversed FFT and bit-reversed IFFT respectively. For the case of digital nets with digitally-shift-invariant kernels and resulting nested block Toeplitz Gram matrices, ${\mathsf{V}}^{(m)}={\mathsf{H}}^{(m)}/\sqrt{n}$ is the scaled Hadamard matrix and ${\widetilde{\boldsymbol{w}}}^{(m)}=\{1\}_{i=0}^{2^{m}-1}$, so ${\mathsf{V}}^{(m)}{\boldsymbol{y}}$ and $\overline{{\mathsf{V}}^{(m)}}{\boldsymbol{y}}$ can be computed using a FWHT and inverse FWHT respectively.
For ${\mathsf{\Lambda}}=\mathrm{diag}({\boldsymbol{\lambda}})$, we have
${\boldsymbol{\lambda}}=\sqrt{n}{\mathsf{\Lambda}}^{(m)}\overline{{\boldsymbol{v}}_{1}^{(m)}}=\sqrt{n}\overline{{\mathsf{V}}^{(m)}}\left({\mathsf{V}}^{(m)}{\mathsf{\Lambda}}^{(m)}\overline{{\boldsymbol{v}}_{1}^{(m)}}\right)=\sqrt{n}\overline{{\mathsf{V}}^{(m)}}{\boldsymbol{k}}_{1}$
which can be computed at ${\mathcal{O}}(n\log n)$ cost. Moreover,
${\mathsf{K}}{\boldsymbol{y}}={\mathsf{V}}^{(m)}({\widetilde{\boldsymbol{y}}}\odot{\boldsymbol{\lambda}})\qquad\text{and}\qquad{\mathsf{K}}^{-1}{\boldsymbol{y}}={\mathsf{V}}^{(m)}({\widetilde{\boldsymbol{y}}}./{\boldsymbol{\lambda}})$
may each be evaluated at cost ${\mathcal{O}}(n\log n)$ where $\odot$ denotes the Hadamard (component-wise) product and ././. / denotes component-wise division.
To update the solutions of the motivating problems, it is easily shown that
${\mathsf{K}}^{(m+1)}{\boldsymbol{y}}^{(m+1)}$ $=\sqrt{2}\begin{pmatrix}{\mathsf{V}}^{(m)}\left({\widetilde{\boldsymbol{y}}}^{(1)}\odot{\boldsymbol{\lambda}}^{(1)}+{\widetilde{\boldsymbol{w}}}^{(m)}\odot{\widetilde{\boldsymbol{w}}}^{(m)}\odot{\widetilde{\boldsymbol{y}}}^{(2)}\odot{\boldsymbol{\lambda}}^{(2)}\right)\\ {\mathsf{V}}^{(m)}\left({\widetilde{\boldsymbol{y}}}^{(2)}\odot{\boldsymbol{\lambda}}^{(1)}+{\widetilde{\boldsymbol{y}}}^{(1)}\odot{\boldsymbol{\lambda}}^{(2)}\right)\end{pmatrix}$
${\mathsf{K}}^{-(m+1)}{\boldsymbol{y}}^{(m+1)}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}{\mathsf{V}}^{(m)}\left(\left({\widetilde{\boldsymbol{y}}}^{(1)}\odot{\boldsymbol{\lambda}}^{(1)}-{\widetilde{\boldsymbol{w}}}^{(m)}\odot{\widetilde{\boldsymbol{w}}}^{(m)}\odot{\widetilde{\boldsymbol{y}}}^{(2)}\odot{\boldsymbol{\lambda}}^{(2)}\right)./{\widetilde{\boldsymbol{\gamma}}}\right)\\ {\mathsf{V}}^{(m)}\left(\left({\widetilde{\boldsymbol{y}}}^{(2)}\odot{\boldsymbol{\lambda}}^{(1)}-{\widetilde{\boldsymbol{y}}}^{(1)}\odot{\boldsymbol{\lambda}}^{(2)}\right)./{\widetilde{\boldsymbol{\gamma}}}\right)\end{pmatrix}$
where ${\widetilde{\boldsymbol{\gamma}}}={\boldsymbol{\lambda}}^{(1)}\odot{\boldsymbol{\lambda}}^{(1)}-{\widetilde{\boldsymbol{w}}}^{(m)}\odot{\widetilde{\boldsymbol{w}}}^{(m)}\odot{\boldsymbol{\lambda}}^{(2)}\odot{\boldsymbol{\lambda}}^{(2)}$.
6. Numerical experiments
All numerical experiments were carried out on a 2019 MacBook Pro with an Intel Core i7 processor. Figure 5 compares the wall-clock time required to generate point sets and perform fast transforms. Since scaling in the number dimensions and number of randomizations is linear, point sets are only generated in $d=1$ with a single randomization. IID points, lattices with shifts, and digital nets with LMS and digital shifts (including higher-order versions) are the fastest sequences to generate. Digital nets in base $b=2$ exploit Graycode order, integer storage of bit-vectors, and exclusive or (XOR) operations to perform digital addition. Generators of Halton point sets are slower to generate as they cannot exploit these advantages. NUS, especially higher order versions, are significantly slower to generate then LMS randomizations. Even so, higher-order LMS scrambling with digital shifts are sufficient to achieve higher-order RMSE convergence as we shown in the next experiment.
Figure 5. Comparison of time required to generate IID and low-discrepancy point sets and perform fast transforms. All point sets are generated in a single dimension and with a single randomizations. BR indicates the bit-reversed FFT and IFFT.
[…]
Figure 6 shows higher-order LMS with digital shifts for base $b=2$ digital nets achieve higher-order RMSE convergence for RQMC methods. Specifically, we are able to achieve an RMSE of ${\mathcal{O}}(n^{-\min\{\alpha,{\widetilde{\alpha}}\}-1/2-\delta})$ where $\alpha$ is the higher-order digital interlacing of the net, ${\widetilde{\alpha}}$ is the smoothness of the integrand, and $\delta>0$ is arbitrarily small. The first two integrands reproduce the experiments of (Dick, 2011). The next two problems are test functions from Derek Bingham’s Virtual Library of Simulation Experiments (VLSE) [^14] https://www.sfu.ca/~ssurjano/uq.html. The last two Genz functions are used for forward uncertainty quantification in the Dakota software package (Adams et al., 2020) among other places. For each problem, the RMSE of the (Q)MC estimator (1) with ${\boldsymbol{X}}\sim{\mathcal{U}}[0,1]^{d}$ is approximated using $300$ independent randomizations of a LD (or IID) point sets. For lattices, we periodize the integrand using a Baker-transform i.e. we use $\tilde{f}({\boldsymbol{x}})=f(1-2\lvert{\boldsymbol{x}}-1/2\rvert)$ in place of $f({\boldsymbol{x}})$ in (1). The Baker-transform does not change the expectation. A description of the integrands follows
Simple function, $d=1$:
has $f(X)=Xe^{X}-1$. This was used in (Dick, 2011) where higher-order digital net scrambling was first proposed.
Simple function, $d=2$:
has $f({\boldsymbol{X}})=X_{2}e^{X_{1}X_{2}}/(e-2)-1$. This was also used in (Dick, 2011).
Oakley & O’Hagan, $d=2$:
has $f({\boldsymbol{X}})=g(({\boldsymbol{X}}-1/2)/50)$ for $g({\boldsymbol{T}})=5+T_{1}+T_{2}+2\cos(T_{1})+2\cos(T_{2})$, see (Oakley and O’hagan, 2002) or the VLSE14.
G-Function, $d=3$:
has $f({\boldsymbol{X}})=\prod_{j=1}^{d}\frac{\lvert 4X_{j}-2\rvert-a_{j}}{1+a_{j}}$ with $a_{j}=(j-2)/2$ for $1\leq j\leq d$, see (Crestaux et al., 2007; Marrel et al., 2009) or the VLSE14.
Oscillatory Genz, $d=3$:
has $f({\boldsymbol{X}})=\cos\left(-\sum_{j=1}^{d}c_{j}X_{j}\right)$ with coefficients of the third kind $c_{j}=\exp\left(k\log\left(10^{-8}\right)/d\right)$. This is a common test function for uncertainty quantification which is is available in the Dakota software (Adams et al., 2020) among other.
Corner-peak Genz, $d=3$:
has $f({\boldsymbol{x}})=\left(1+\sum_{j=1}^{d}c_{j}X_{j}\right)^{-(d+1)}$ with coefficients of the second kind $c_{j}=1/j^{2}$. This is also used in the Dakota software (Adams et al., 2020) among others.
Figure 6. The RMSE of the RQMC estimate for a few different integrands. Higher-order digital nets with linear matrix scrambling and digital shifts achieve higher-order convergence.
[…]
7. Conclusion and future work
This work has reviewed routines for generating low-discrepancy point sets, randomizing them, and applying them to fast kernel methods. We have reviewed lattices, digital nets, and Halton point sets with randomizations spanning shifts, digital permutations, digital shifts, linear matrix scrambling, and nested uniform scrambling. Higher-order scramblings for digital nets were also considered. For kernel methods, we presented classes of kernels which pair with low-discrepancy sequences to accelerate kernel computations. This includes new higher-order digitally shift invariant kernels. The accelerated methods utilize modified fast Fourier transforms and the fast Walsh-Hadamard transform. The presented material has been implemented into the accessible open-source Python library QMCPy.
Acknowledgements.
Thank you to Fred J Hickernell, Sou-Cheng T Choi, and Aadit Jain for helpful comments and prototyping. The author would like to acknowledge the support of NSF Grant DMS-2316011.
Appendix A Proofs of theorems
For $k\in{\mathbb{N}}_{0}$ write
${\widehat{f}}(k)=\int_{0}^{1}f(x)\overline{\mathrm{wal}_{k}(x)}\mathrm{d}x.$
Proof of theorem 5.1.
Suppose $f\in H_{\alpha}$. Using the fact that (6) is a Mercer kernel and (Dick, 2009a, Theorem 14, Remark 19), we have
$\left\lVert f\right\rVert_{{\widetilde{H}}_{\alpha}}^{2}=\sum_{k\in{\mathbb{N}}_{0}}\left\lvert{\widehat{f}}(k)\right\rvert^{2}b^{\mu_{\alpha}(k)}\leq C_{\alpha}\sum_{k\in{\mathbb{N}}}b^{-\mu_{\alpha}(k)}<\infty$
for some $C_{\alpha}<\infty$, so $f\in{\widetilde{H}}_{K}$. ∎
Lemma A.1 (Walsh coefficients of low order monomials).
Fix $b=2$. Let $f_{p}(x):=x^{p}$. When $k\in{\mathbb{N}}$ write
$k=2^{a_{1}}+\dots+2^{a_{\#k}}$
where $a_{1}>a_{2}>\dots>a_{\#k}\geq 0$. Then we have
${\widehat{f}}_{1}(k)$ =$=\begin{cases}1/2,&k=0\\ -2^{-a_{1}-2},&k=2^{a_{1}}\\ 0,&\mathrm{otherwise}\end{cases},$
${\widehat{f}}_{2}(k)$ =$=\begin{cases}1/3,&k=0\\ -2^{-a_{1}-2},&k=2^{a_{1}}\\ 2^{-a_{1}-a_{2}-3},&k=2^{a_{1}}+2^{a_{2}}\\ 0,&\mathrm{otherwise}\end{cases},$
${\widehat{f}}_{3}(k)$ =$=\begin{cases}1/4,&k=0\\ -2^{-a_{1}-2}+2^{-3a_{1}-5},&k=2^{a_{1}}\\ 3\cdot 2^{-a_{1}-a_{2}-4},&k=2^{a_{1}}+2^{a_{2}}\\ -3\cdot 2^{-a_{1}-a_{2}-a_{3}-5},&k=2^{a_{1}}+2^{a_{2}}+2^{a_{3}}\\ 0,&\mathrm{otherwise}\end{cases}.$
Proof.
The forms for ${\widehat{f}}_{1}$ and ${\widehat{f}}_{2}$ follow from (Dick and Pillichshammer, 2010, Example 14.2, Example 14.3). For $k=0$ and any $x\in[0,1)$ we have $\mathrm{wal}_{0}(x)=1$, so
${\widehat{f}}_{3}(x)=\int_{0}^{1}x^{3}\mathrm{d}x=1/4.$
Assume $k\in{\mathbb{N}}$ going forward. For $k=2^{a_{1}}+k^{\prime}$ with $0\leq k^{\prime}<2^{a_{1}}$, (Fine, 1949, Equation 3.6) implies
$J_{k}(x):=\int_{0}^{x}\mathrm{wal}_{k}(t)\mathrm{d}t=2^{-a_{1}-2}\left[\mathrm{wal}_{k^{\prime}}(x)-\sum_{r=1}^{\infty}2^{-r}\mathrm{wal}_{2^{a_{1}+r}+k}(x)\right].$
Using integration by parts and the fact that $J_{k}(0)=J_{k}(1)=0$
${\widehat{f}}_{3}(k)$ $=\int_{0}^{1}x^{3}\mathrm{wal}_{k}(x)\mathrm{d}x=\left[x^{3}J_{k}(x)\right]_{x=0}^{x=1}-3\int_{0}^{1}x^{2}J_{k}(x)\mathrm{d}x$
$=-3*2^{-a_{1}-2}\left[{\widehat{f}}_{2}(k^{\prime})-\sum_{r=1}^{\infty}2^{-r}{\widehat{f}}_{2}(2^{a_{1}+r}+k)\right].$
  • If #$\#k=1$, i.e. $k=2^{a_{1}}$ then
    ${\widehat{f}}_{3}(k)$ $=-3*2^{-a_{1}-2}\left[{\widehat{f}}_{2}(0)-\sum_{r=1}^{\infty}2^{-r}{\widehat{f}}_{2}(2^{a_{1}+r}+2^{a_{1}})\right]$
    $=-3*2^{-a_{1}-2}\left[\frac{1}{3}-\sum_{r=1}^{\infty}2^{-r}2^{-(a_{1}+r)-a_{1}-3}\right]$
    $=2^{-3a_{1}-5}-2^{-a_{1}-2}.$
  • If #$\#k=2$ then
    ${\widehat{f}}_{3}(k)=-3*2^{-a_{1}-2}{\widehat{f}}_{2}(2^{a_{2}})=3\cdot 2^{-a_{1}-a_{2}-4}.$
  • If #$\#k=3$ then
    ${\widehat{f}}_{3}(k)=-3*2^{-a_{1}-2}{\widehat{f}}_{2}(2^{a_{2}}+2^{a_{3}})=-3\cdot 2^{-a_{1}-a_{2}-a_{3}-5}.$
  • If #$\#k>3$ then ${\widehat{f}}_{3}(k)=0$.
Proof of theorem 5.2.
Write
${\widetilde{K}}_{\alpha}(x)=1+\sum_{1\leq\nu<\alpha}s_{\nu}(x)+{\widetilde{s}}_{\alpha}(x)$
where $s_{\nu}$ sums over all $k$ with #$\#k=\nu$ and ${\widetilde{s}}_{\alpha}$ sums over all $k$ with #$\#k\geq\alpha$. In (Baldeaux et al., 2012, Corollary 1) it was shown that
$s_{1}(x)$ $=-2x+1,$
$s_{2}(x)$ $=2x^{2}-2x+\frac{1}{3},$
${\widetilde{s}}_{2}(x)$ $=\left[2-\beta(x)\right]x+\frac{1}{2}\left[1-5t_{1}(x)\right],$
${\widetilde{s}}_{3}(x)$ $=-\left[2-\beta(x)\right]x^{2}-\left[1-5t_{1}(x)\right]x+\frac{1}{18}\left[1-43t_{2}(x)\right]$
from which ${\widetilde{K}}_{2}$ and ${\widetilde{K}}_{3}$ follow. We now find expressions for $s_{3}$ and ${\widetilde{s}}_{4}$ from which ${\widetilde{K}}_{4}$ follows.
Assume sums over $a_{i}$ are over ${\mathbb{N}}_{0}$ unless otherwise restricted. Recall from Lemma A.1 that
$x$ $=\frac{1}{2}-\sum_{a_{1}}\frac{\mathrm{wal}_{2^{a_{1}}}(x)}{2^{a_{1}+2}},$
$x^{2}$ $=\frac{1}{3}-\sum_{a_{1}}\frac{\mathrm{wal}_{2^{a_{1}}}(x)}{2^{a_{1}+2}}+\sum_{a_{1}>a_{2}}\frac{\mathrm{wal}_{2^{a_{1}}+2^{a_{2}}}(x)}{2^{a_{1}+a_{2}+3}},$
$x^{3}$ $=\frac{1}{4}-\sum_{a_{1}}\frac{\mathrm{wal}_{2^{a_{1}}}(x)}{2^{a_{1}+2}}+\frac{3}{2}\sum_{a_{1}>a_{2}}\frac{\mathrm{wal}_{2^{a_{1}}+2^{a_{2}}}(x)}{2^{a_{1}+a_{2}+3}}-\frac{3}{2}\sum_{a_{1}>a_{2}>a_{3}}\frac{\mathrm{wal}_{2^{a_{1}}+2^{a_{2}}+2^{a_{3}}}(x)}{2^{a_{1}+a_{2}+a_{3}+4}}+\sum_{a_{1}}\frac{\mathrm{wal}_{2^{a_{1}}}(x)}{2^{3a_{1}+5}}$
so
$x^{3}-\frac{3}{2}x^{2}+\frac{1}{2}x=\frac{1}{32}\sum_{a_{1}}\frac{\mathrm{wal}_{2^{a_{1}}}(x)}{2^{3a_{1}}}-\frac{3}{4}\sum_{a_{1}>a_{2}>a_{3}}\frac{\mathrm{wal}_{2^{a_{1}}+2^{a_{2}}+2^{a_{3}}}(x)}{2^{a_{1}+a_{2}+a_{3}+3}}$
and
$s^{[3}(x)=\sum_{a_{1}>a_{2}>a_{3}}\frac{\mathrm{wal}_{2^{a_{1}}+2^{a_{2}}+2^{a_{3}}}(x)}{2^{a_{1}+a_{2}+a_{3}+3}}=-\frac{4}{3}x^{3}+2x^{2}-\frac{2}{3}x+\frac{1}{24}\sum_{a_{1}}\frac{\mathrm{wal}_{2^{a_{1}}}(x)}{2^{3a_{1}}}.$
Now,
${\widetilde{s}}_{4}(x)$ $=\sum_{\begin{subarray}{c}a_{1}>a_{2}>a_{3}>a_{4}\\ 0\leq k<2^{a_{4}}\end{subarray}}\frac{\mathrm{wal}_{2^{a_{1}}+2^{a_{2}}+2^{a_{3}}+2^{a_{4}}+k}(x)}{2^{a_{1}+a_{2}+a_{3}+a_{4}+4}}$
$=\sum_{a_{1}>a_{2}>a_{3}>a_{4}}\frac{\mathrm{wal}_{2^{a_{1}}+2^{a_{2}}+2^{a_{3}}+2^{a_{4}}}(x)}{2^{a_{1}+a_{2}+a_{3}+a_{4}+4}}\sum_{0\leq k<2^{a_{4}}}\mathrm{wal}_{k}(x)$
If $x=0$ then
${\widetilde{s}}_{4}(0)=\sum_{a_{1}>a_{2}>a_{3}>a_{4}}\frac{1}{2^{a_{1}+a_{2}+a_{3}+4}}=\frac{1}{294}.$
Going forward, assume $x\in(0,1)$ so $\beta(x)=-\lfloor\log_{2}(x)\rfloor$ is finite. Recall that
$\sum_{0\leq k<2^{a_{4}}}\mathrm{wal}_{k}(x)=\begin{cases}2^{a_{4}},&a_{4}\leq\beta(x)-1\\ 0,&a_{4}>\beta(x)-1\end{cases}.$
Moreover, since $\beta(x)$ is the index of the first $1$ in the base $2$ expansion of $x$, when $a_{4}<\beta(x)-1$ we have $\mathrm{wal}_{2^{a_{4}}}(x)=(-1)^{{\mathsf{x}}_{a_{4}+1}}=1$ and when $a_{4}=\beta(x)-1$ we have $\mathrm{wal}_{2^{a_{4}}}(x)=-1$. This implies
${\widetilde{s}}_{4}(x)$ $=\sum_{\begin{subarray}{c}a_{1}>a_{2}>a_{3}>a_{4}\\ \beta(x)-1\geq a_{4}\end{subarray}}\frac{\mathrm{wal}_{2^{a_{1}}+2^{a_{2}}+2^{a_{3}}+2^{a_{4}}}(x)}{2^{a_{1}+a_{2}+a_{3}+4}}$
$=\sum_{\begin{subarray}{c}a_{1}>a_{2}>a_{3}>a_{4}\\ \beta(x)-1>a_{4}\end{subarray}}\frac{\mathrm{wal}_{2^{a_{1}}+2^{a_{2}}+2^{a_{3}}}(x)}{2^{a_{1}+a_{2}+a_{3}+4}}-\sum_{a_{1}>a_{2}>a_{3}>\beta(x)-1}\frac{\mathrm{wal}_{2^{a_{1}}+2^{a_{2}}+2^{a_{3}}}(x)}{2^{a_{1}+a_{2}+a_{3}+4}}$
$=:T_{1}-T_{2}.$
The first term is
$T_{1}$ $=\sum_{\beta(x)-1>a_{4}}\bigg{(}\sum_{a_{1}>a_{2}>a_{3}}\frac{\mathrm{wal}_{2^{a_{1}}+2^{a_{2}}+2^{a_{3}}}(x)}{2^{a_{1}+a_{2}+a_{3}+4}}-\sum_{a_{4}\geq a_{1}>a_{2}>a_{3}}\frac{1}{2^{a_{1}+a_{2}+a_{3}+4}}$
$\qquad\qquad-\sum_{a_{1}>a_{4}\geq a_{2}>a_{3}}\frac{\mathrm{wal}_{2^{a_{1}}}(x)}{2^{a_{1}+a_{2}+a_{3}+4}}-\sum_{a_{1}>a_{2}>a_{4}\geq a_{3}}\frac{\mathrm{wal}_{2^{a_{1}}+2^{a_{2}}}(x)}{2^{a_{1}+a_{2}+a_{3}+4}}\bigg{)}$
$=:\sum_{\beta(x)-1>a_{4}}\left[V_{1}(a_{4})-V_{2}(a_{4})-V_{3}(a_{4})-V_{4}(a_{4})\right].$
Clearly $V_{1}(a_{4})=s_{3}(x)/2$ and $V_{2}$ is easily computed. Now
$V_{3}(a_{4})$ $=\left(\sum_{a_{4}\geq a_{2}>a_{3}}\frac{1}{2^{a_{2}+a_{3}+3}}\right)\left(\sum_{a_{1}>a_{4}}\frac{\mathrm{wal}_{2^{a_{1}}}(x)}{2^{a_{1}+1}}\right)$
$=\left(\sum_{a_{4}\geq a_{2}>a_{3}}\frac{1}{2^{a_{2}+a_{3}+3}}\right)\left(s_{1}(x)-\sum_{a_{4}\geq a_{1}}\frac{1}{2^{a_{1}+1}}\right)$
and
$V_{4}(a_{4})$ $=\left(\sum_{a_{4}\geq a_{3}}\frac{1}{2^{a_{3}+2}}\right)\left(\sum_{a_{1}>a_{2}>a_{4}}\frac{\mathrm{wal}_{2^{a_{1}}+2^{a_{2}}}(x)}{2^{a_{1}+a_{2}+2}}\right)$
$=\left(\sum_{a_{4}\geq a_{3}}\frac{1}{2^{a_{3}+2}}\right)\left(\sum_{a_{1}>a_{2}}\frac{\mathrm{wal}_{2^{a_{1}}+2^{a_{2}}}(x)}{2^{a_{1}+a_{2}+2}}-\sum_{a_{4}\geq a_{1}>a_{2}}\frac{1}{2^{a_{1}+a_{2}+2}}-\sum_{a_{1}>a_{4}\geq a_{2}}\frac{\mathrm{wal}_{2^{a_{1}}}(x)}{2^{a_{1}+a_{2}+2}}\right)$
$=\left(\sum_{a_{4}\geq a_{3}}\frac{1}{2^{a_{3}+2}}\right)\left(s_{2}(x)-\sum_{a_{4}\geq a_{1}>a_{2}}\frac{1}{2^{a_{1}+a_{2}+2}}-\left(s_{1}(x)-\sum_{a_{4}\geq a_{1}}\frac{1}{2^{a_{1}+1}}\right)\left(\sum_{a_{4}\geq a_{2}}\frac{1}{2^{a_{2}+1}}\right)\right).$
The second term is
$T_{2}$ $=\sum_{a_{1}>a_{2}>a_{3}}\frac{\mathrm{wal}_{2^{a_{1}}+2^{a_{2}}+2^{a_{3}}}(x)}{2^{a_{1}+a_{2}+a_{3}+4}}-\sum_{\beta(x)-1>a_{1}>a_{2}>a_{3}}\frac{1}{2^{a_{1}+a_{2}+a_{3}+4}}+\sum_{\beta(x)-1>a_{2}>a_{3}}\frac{1}{2^{\beta(x)+a_{2}+a_{3}+3}}$
$-\sum_{a_{1}>\beta(x)-1>a_{2}>a_{3}}\frac{\mathrm{wal}_{2^{a_{1}}}(x)}{2^{a_{1}+a_{2}+a_{3}+4}}+\sum_{a_{1}>\beta(x)-1>a_{3}}\frac{\mathrm{wal}_{2^{a_{1}}}(x)}{2^{a_{1}+\beta(x)+a_{3}+3}}$
$-\sum_{a_{1}>a_{2}>\beta(x)-1>a_{3}}\frac{\mathrm{wal}_{2^{a_{1}}+2^{a_{2}}}(x)}{2^{a_{1}+a_{2}+a_{3}+4}}+\sum_{a_{1}>a_{2}>\beta(x)-1}\frac{\mathrm{wal}_{2^{a_{1}}+2^{a_{2}}}(x)}{2^{a_{1}+a_{2}+\beta(x)+3}}$
$=:W_{1}-W_{2}+W_{3}-W_{4}+W_{5}-W_{6}+W_{7}.$
Clearly $W_{1}=s_{3}(x)/2$ and both $W_{2}$ and $W_{3}$ are easily computed. Similarity in the next two sums gives
$W_{5}-W_{4}$ $=\left(\sum_{\beta(x)-1>a_{3}}\frac{1}{2^{\beta(x)+a_{3}+2}}-\sum_{\beta(x)-1>a_{2}>a_{3}}\frac{1}{2^{a_{2}+a_{3}+3}}\right)\left(\sum_{a_{1}>\beta(x)-1}\frac{\mathrm{wal}_{2^{a_{1}}}(x)}{2^{a_{1}+1}}\right)$
$=\left(\sum_{\beta(x)-1>a_{3}}\frac{1}{2^{\beta(x)+a_{3}+2}}-\sum_{\beta(x)-1>a_{2}>a_{3}}\frac{1}{2^{a_{2}+a_{3}+3}}\right)\left(s_{1}(x)-\sum_{\beta(x)-1>a_{1}}\frac{1}{2^{a_{1}+1}}+\frac{1}{2^{\beta(x)}}\right)$
Similarity in the final two sums gives
$W_{7}-W_{6}$ $=\left(\frac{1}{2^{\beta(x)+1}}-\sum_{\beta(x)-1>a_{3}}\frac{1}{2^{a_{3}+2}}\right)\left(\sum_{a_{1}>a_{2}>\beta(x)-1}\frac{\mathrm{wal}_{2^{a_{1}}+2^{a_{2}}}(x)}{2^{a_{1}+a_{2}+2}}\right).$
where
$\sum_{a_{1}>a_{2}>\beta(x)-1}\frac{\mathrm{wal}_{2^{a_{1}}+2^{a_{2}}}(x)}{2^{a_{1}+a_{2}+2}}$ $=\sum_{a_{1}>a_{2}}\frac{\mathrm{wal}_{2^{a_{1}}+2^{a_{2}}}(x)}{2^{a_{1}+a_{2}+2}}-\sum_{\beta(x)-1>a_{1}>a_{2}}\frac{1}{2^{a_{1}+a_{2}+2}}+\sum_{\beta(x)-1>a_{2}}\frac{1}{2^{\beta(x)+a_{2}+1}}$
$-\sum_{a_{1}>\beta(x)-1>a_{2}}\frac{\mathrm{wal}_{2^{a_{1}}}(x)}{2^{a_{1}+a_{2}+2}}+\sum_{a_{1}>\beta(x)-1}\frac{\mathrm{wal}_{2^{a_{1}}}(x)}{2^{a_{1}+\beta(x)+1}}$
$=s_{2}(x)-\sum_{\beta(x)-1>a_{1}>a_{2}}\frac{1}{2^{a_{1}+a_{2}+2}}+\sum_{\beta(x)-1>a_{2}}\frac{1}{2^{\beta(x)+a_{2}+1}}$
$+\left(\frac{1}{2^{\beta(x)}}-\sum_{\beta(x)-1>a_{2}}\frac{1}{2^{a_{2}+1}}\right)\left(s_{1}(x)-\sum_{\beta(x)-1>a_{1}}\frac{1}{2^{a_{1}+1}}+\frac{1}{2^{\beta(x)}}\right).$
This implies
${\widetilde{s}}_{4}(x)$ $=\frac{2}{3}\left(2-\beta(x)\right)x^{3}+\left(1-5\ t_{1}(x)\right)x^{2}-\frac{1}{9}\left(1-43t_{2}(x)\right)x$
$-\frac{1}{48}\left(2-\beta(x)\right)\sum_{a_{1}}\frac{\mathrm{wal}_{2^{a_{1}}}(x)}{2^{3a_{1}}}-\frac{1}{294}\left(7\beta(x)+701t_{3}(x)\right)+\frac{5}{98}$
from which the result follows. ∎

### v2

@@
-A Unified Implementation of Quasi-Monte Carlo Generators, Randomization Routines, and Fast Kernel Methods
-Abstract.
-Quasi-random sequences, also called low-discrepancy sequences, have been extensively used as efficient experimental designs across many scientific disciplines.
-This article provides a unified description and software API for methods pertaining to low-discrepancy point sets.
-These methods include low discrepancy point set generators, randomization techniques, and fast kernel methods.
-Specifically, we provide generators for lattices, digital nets, and Halton point sets.
-Supported randomization techniques include random permutations / shifts, linear matrix scrambling, and nested uniform scrambling.
-Routines for working with higher-order digital nets and scramblings are also detailed.
-For kernel methods, we provide implementations of special shift-invariant and digitally-shift invariant kernels along with fast Gram matrix operations facilitated by the bit-reversed FFT, the bit-reversed IFFT, and the FWHT.
-A new digitally-shift-invariant kernel of higher-order smoothness is also derived.
-We also describe methods to quickly update the matrix-vector product or linear system solution after doubling the number of points in a lattice or digital net in natural order.
-Low-discrepancy sequences, low-discrepancy point set randomizations, higher-order digital net scrambling, randomized Quasi-Monte Carlo, fast kernel methods, shift-invariant kernels, digitally-shift-invariant kernels
-copyright: acmlicensedjournalyear: 2025doi: 0.0journal: TOMSjournalvolume: 0journalnumber: 0article: 0publicationmonth: 0submissionid: 0-0-0ccs: Mathematics of computing Statistical softwareccs: Mathematics of computing Mathematical software performanceccs: Mathematics of computing Solversccs: Mathematics of computing Probabilistic algorithms
-1. Introduction
-This article describes and implements a number of algorithms related to low-discrepancy point sets. These include implementations of
-  • low-discrepancy point set generators and randomizations,
-  • higher-order digital net generators and randomizations,
-  • shift-invariant and digitally-shift-invariant kernels of higher order smoothness,
-  • and fast transforms which enable fast Gram matrix operations when special kernels are paired with certain low-discrepancy point sets.
-While these methods have been developed and often extensively studied throughout the literature, we found implementations to be scattered or even missing in some popular programming languages.
-Our main contribution is to provide a unified and accessible software for computations using low-discrepancy sequences in the QMCPy [^1] https://qmcpy.readthedocs.io Python package (Choi et al., 2022b, a).
-We also integrate with the new LDData [^2] https://github.com/QMCSoftware/LDData repository which provides generating vectors for lattices and generating matrices for digital nets in standardized formats.
-This collection is, so far, mainly an aggregation of results from the Magic Point Shop [^3] https://people.cs.kuleuven.be/~dirk.nuyens/qmc-generators/ (Kuo and Nuyens, 2016a) and the websites of Frances Kuo on lattices [^4] https://web.maths.unsw.edu.au/~fkuo/lattice/index.html (Cools et al., 2006; Nuyens and Cools, 2006) and Sobol’ points [^5] https://web.maths.unsw.edu.au/~fkuo/sobol/index.html (Joe and Kuo, 2003, 2008).
-Below we summarize the novel components of our QMCPy software which we detail in this paper.
-Comparisons to existing implementations are also discussed.
-A more comprehensive review of Quasi-Monte Carlo software is given in (Choi et al., 2022b), along with more details on the QMCPy framework.
-Lattice point sets:
-We support generating points in linear, natural, or Gray code order and randomizing points with shifts modulo one.
-Other implementations of shifted lattices include those in MATLAB’s GAIL [^6] http://gailgithub.github.io/GAIL_Dev/ (Guaranteed Automatic Integration Library) (Tong et al., 2022; Hickernell et al., 2018), the C++ LatNet Builder [^7] https://github.com/umontreal-simul/latnetbuilder software (L’Ecuyer et al., 2021), and the multi-language MPS [^8] https://people.cs.kuleuven.be/~dirk.nuyens/qmc-generators/ (Magic Point Shop) (Kuo and Nuyens, 2016b).
-LatNet Builder and MPS also provide search routines for finding good generating vectors which define lattice point sets.
-Digital nets in base $2$:
-We support generating digital nets in natural or Gray code order and randomizing with linear matrix scrambling (LMS) (Owen, 2003), digital shifts, nested uniform scrambling (NUS) (Owen, 1995), and permutation scrambling.
-To the best of our knowledge, we are also the only Python implementation to support digital interlacing for the creating of higher order digital nets and higher order scrambling with either LMS or NUS.
-Early implementations of unrandomized digital sequences, including the Faure, Sobol’ and Niederreiter-Xing constructions, can be found in (Fox, 1986; Bratley et al., 1992; Bratley and Fox, 2003; Pirsic, 2002).
-Considerations for implementing scrambles for digital sequences was discussed in (Hong and Hickernell, 2003).
-Recent digital net implementations supporting LMS and digital shifts include those in MPS, PyTorch (Paszke et al., 2019), SciPy (Virtanen et al., 2020), and MATLAB (Inc., 2022).
-LatNet Builder provides robust support for all randomization routines and digital interlacing for higher order nets and scramblings.
-LatNet Builder and MPS also provide search routines for finding good generating matrices which define digital nets.
-Halton point sets:
-As with digital nets, we support randomizing Halton point sets with LMS, digital shifts, NUS, and permutation scrambles.
-The implementation of Halton point sets and randomizations have been treated in (Owen, 2017; Wang and Hickernell, 2000).
-The QRNG [^9] https://cran.r-project.org/web/packages/qrng/qrng.pdf (Quasi-Random Number Generators) R package (Hofert and Lemieux, 2023) implements generalized digital nets which use optimized digital permutation scrambles; these are also supported in QMCPy.
-LatNet Builder also provides support for Halton points and their randomizations.
-Kernel Methods and Fast Transforms:
-We provide routines to evaluate special shift-invariant and digitally-shift-invariant kernels.
-This includes a new digitally-shift-invariant kernel of higher-order smoothness derived in this paper.
-These are not readily available elsewhere.
-We also provide routines for fast matrix-vector multiplication and inversion of the nicely-structured Gram matrices produced when pairing these special kernels with certain low-discrepancy sequences.
-We also describe novel routines to efficiently update the matrix-vector products and linear system solves in the Gram matrix after doubling the number of lattice or digital net points.
-These fast operations are backed by implementations of the Fast Walsh Hadamard Transform (FWHT), the bit-reversed fast Fourier Transform (FFT), and the bit-reversed inverse FFT (IFFT).
-We are among the first to provide efficient and accessible Python interfaces to such fast kernel methods and the underlying fast transforms.
-GAIL has implemented fast Bayesian cubature routines utilizing this machinery.
-These above routines are most commonly used in Quasi-Monte Carlo methods and for fast kernel methods, which we describe in the following subsections.
-1.1. Quasi-Monte Carlo methods
-Monte Carlo (MC) methods approximate a high dimensional integral over the unit cube by the sample average of function evaluations at certain sampling locations:
-(1) $\mu:=\int_{[0,1]^{d}}f({\boldsymbol{x}})\mathrm{d}{\boldsymbol{x}}\approx\frac{1}{n}\sum_{i=0}^{n-1}f({\boldsymbol{x}}_{i})=:{\widehat{\mu}}.$
+QMCPy: A Python Software for Randomized Low-Discrepancy Sequences, Quasi-Monte Carlo, and Fast Kernel Methods
+Abstract
+Low-discrepancy (LD) sequences have been extensively used as efficient experimental designs across many scientific disciplines.
+QMCPy (https://qmcsoftware.github.io/QMCSoftware/) is an accessible Python library which provides a unified implementation of randomized LD sequences, automatic variable transformations, adaptive Quasi-Monte Carlo error estimation algorithms, and fast kernel methods.
+This article focuses on recent updates to QMCPy which broaden support for randomized LD sequences and add new tools to enable fast kernel methods using LD sequences.
+Specifically, we give a unified description of the supported LD lattices, digital nets, and Halton point sets, along with randomization options including random permutations / shifts, linear matrix scrambling (LMS), and nested uniform scrambling (NUS).
+We also support higher-order digital nets, higher-order scrambling with LMS or NUS, and Halton scrambling with LMS or NUS.
+For fast kernel methods, we provide shift-invariant (SI) and digitally-shift-invariant (DSI) kernels, including a new set of higher-order smoothness DSI kernels.
+When SI and DSI kernels are respectively paired with $n$ LD lattice and digital net points, the resulting Gram matrices permit multiplication and inversion at only $\mathcal{O}(n\log n)$ cost.
+These fast operations utilize QMCPy’s implementation of the fast Fourier transform in bit-reversed order (FFTBR), inverse FFTBR (IFFTBR), and fast Walsh–Hadamard transform (FWHT).
+1 Introduction
+Low-discrepancy (LD) sequences, also called quasi-random sequences, judiciously explore the unit cube in high dimensions.
+They were first developed to replace the independent points used in Monte Carlo simulation for high dimensional numerical integration.
+The resulting Quasi-Monte Carlo (QMC) methods have proven theoretically superior to independent Monte Carlo methods for a large class of nicely behaved functions [Niederreiter, 1992, Dick and Pillichshammer, 2010, Kroese et al., 2013, Dick et al., 2022, Lemieux, 2009, Sloan and Joe, 1994, Dick et al., 2013].
+This advantage has been shown across a variety of scientific disciplines including financial modeling [Joy et al., 1996, Lai and Spanier, 1998, L’Ecuyer, 2004, 2009, Xu and Ökten, 2015, Giles and Waterhouse, 2009], solving PDEs with random coefficients [Graham et al., 2011, Kuo et al., 2012, 2015, Graham et al., 2015, Kuo and Nuyens, 2016, Robbe et al., 2017], and graphics rendering using ray tracing [Jensen et al., 2003, Raab et al., 2006, Waechter and Keller, 2011] among others.
+Randomizing LD sequences provide additional advantages, such as enabling practical QMC error estimation, avoiding boundary observations, and negating the probability of pairing to adverse functions [Owen, 1995, L’Ecuyer, 2016, L’Ecuyer et al., 2023, Owen, 2003, Matoušek, 1998, Tezuka, 2002].
+LD sequences, especially randomized versions, have also been extensively used as efficient experimental designs.
+In reproducing kernel Hilbert spaces (RKHSs), pairing certain LD points with certain kernels yields structured Gram matrices which are diagonalizable by fast transforms [Zeng et al., 2009, 2006].
+For example, pairing an LD rank-1 lattice in linear order with a shift-invariant kernel gives a circulant Gram matrix which is diagonalizable by the fast Fourier transform (FFT) and inverse FFT (IFFT).
+Such structures have been exploited to accelerate kernel interpolation methods for solving PDEs with random coefficients [Kaarnioja et al., 2022, 2023, Sorokin et al., 2024], Bayesian cubature [Rathinavel and Hickernell, 2019, 2022, Rathinavel, 2019], and discrepancy measure computations [Hickernell, 1998a, b].
+LD points have also found recent applications to training scientific machine learning models [Longo et al., 2021, Chen et al., 2021, Keller et al., 2025], although this is an avenue we do not discuss in this paper.
+While randomized LD sequences have been extensively studied for QMC and fast kernel methods, implementations of such methods are often scattered or even missing in some popular programming languages.
+Our prior works [Choi et al., 2022b, a, Hickernell et al., 2025, Sorokin and Rathinavel, 2022] described the current QMC software landscape and introduced QMCPy: a community driven implementation of QMC routines into a unified and accessible Python framework.
+1.1 QMCPy features
+Below we summarize the supported features in QMCPy [^1] https://qmcsoftware.github.io/QMCSoftware/ and provide pointers to related codes:
+Randomized LD Sequences
+We support the point sets and randomization routines described below.
+These features are also supported in the comprehensive C++ LatNet Builder [^2] https://github.com/umontreal-simul/latnetbuilder software [L’Ecuyer et al., 2021].
+LatNet Builder and the multi-language Magic Point Shop (MPS) [^3] https://people.cs.kuleuven.be/~dirk.nuyens/qmc-generators/ [Kuo and Nuyens, 2016] both provide search routines for finding good lattice generating vectors and digital net generating matrices.
+QMCPy integrates with the new LDData [^4] https://github.com/QMCSoftware/LDData repository which contains a variety of these pregenerated vectors and matrices in standardized formats.
+LDData additionally includes popular choices from the websites of Frances Kuo on lattices [^5] https://web.maths.unsw.edu.au/~fkuo/lattice/index.html [Cools et al., 2006, Nuyens and Cools, 2006] and Sobol’ points [^6] https://web.maths.unsw.edu.au/~fkuo/sobol/index.html (a special case of digital nets) [Joe and Kuo, 2003, 2008].
+Lattice Points
+Rank-1 lattices may be generated in radical inverse or linear order, and may be randomized using shifts modulo one.
+Randomly shifted rank-1 lattices are also available in MPS and GAIL [^7] http://gailgithub.github.io/GAIL_Dev/ (MATLAB’s Guaranteed Automatic Integration Library) [Tong et al., 2022, Hickernell et al., 2018].
+Digital Nets
+Base $2$ digital nets, including higher-order digital nets, may be generated in either radical inverse or Gray code order, and may be randomized with linear matrix scrambling (LMS) [Owen, 2003], digital shifts, nested uniform scrambling (NUS) [Owen, 1995], and/or permutation scrambling.
+Early implementations of unrandomized digital sequences, including the Faure, Sobol’, and Niederreiter–Xing constructions, can be found in [Fox, 1986, Bratley et al., 1992, Bratley and Fox, 2003, Pirsic, 2002].
+Considerations for implementing scrambles were discussed in [Hong and Hickernell, 2003].
+Support for combining LMS with digital shifts is also provided in MATLAB, MPS, and both the PyTorch [^8] https://pytorch.org/ [Paszke et al., 2019] and SciPy [^9] https://scipy.org/ [Virtanen et al., 2020] Python packages.
+Halton Points
+As with digital nets, Halton point sets may be randomized with LMS, digital shifts, NUS, and/or permutation scrambles.
+The implementation of Halton point sets and randomizations have been treated in [Owen, 2017, Wang and Hickernell, 2000].
+The QRNG [^10] https://cran.r-project.org/web/packages/qrng/qrng.pdf (Quasi-Random Number Generators) R package [Hofert and Lemieux, 2023] implements generalized Halton point sets [Faure and Lemieux, 2009] which use optimized digital permutation scrambles; these are also supported in QMCPy.
+Variable Transformations
+These define the distribution of stochasticity in the underlying problem and automatically rewrite user-defined functions into QMC-compatible forms.
+The available transforms are mainly wrappers around distributions provided by the SciPy Python package.
+Error Estimators for (Q)MC and Multilevel (Q)MC
+We provide numerous adaptive error estimation algorithms which automatically select the number of points required for a (Q)MC approximation to be within user-specified error tolerances.
+QMC error estimation is treated more broadly in [Owen, 2025], while [Clancy et al., 2014, Hickernell et al., 2017] detail additional considerations for adaptive QMC algorithms.
+QMCPy’s adaptive error estimation procedures are described below.
+The single level algorithms were also implemented in GAIL.
+Monte Carlo with Independent Points
+Confidence intervals are derived from either a central limit theorem (CLT) heuristic or a guaranteed version of CLT for functions with bounded kurtosis [Hickernell et al., 2013].
+QMC with Multiple Randomizations
+Student’s-$t$ confidence intervals are evaluated based on independent mean estimates from independent randomizations of an LD point set, see [L’Ecuyer et al., 2023] or [Owen, 2023, Chapter 17].
+QMC via Decay Tracking using a Single Randomized LD Sequence
+Guaranteed error bounds are available for cones of functions whose Fourier or Walsh coefficients decay predictably [Hickernell and Jiménez Rugama, 2014, Hickernell et al., 2017, Jiménez Rugama and Hickernell, 2014, Ding et al., 2018].
+QMC via Bayesian Cubature using a Single Randomized LD Sequence
+Posterior credible intervals on the integral of a Gaussian process may be computed quickly by exploiting fast kernel computations [Rathinavel and Hickernell, 2019, 2022, Rathinavel, 2019].
+Multilevel Monte Carlo with Independent Points
+We have implemented standard multilevel Monte Carlo [Giles, 2008] and continuation multilevel Monte Carlo algorithms [Collier et al., 2015, Robbe et al., 2016, 2019].
+Multilevel QMC with Multiple Randomizations
+We have implemented standard multilevel QMC [Giles and Waterhouse, 2009] and continuation multilevel QMC algorithms [Robbe et al., 2017].
+Fast Kernel Computations
+Matching a rank-1 lattice in radical inverse order to a shift-invariant (SI) RKHS kernel yields a Gram matrix diagonalizable by the FFT in bit-reversed order (FFTBR) and its inverse (IFFTBR).
+Similarly, matching a base $2$ digital net in radical inverse order to a digitally-shift-invariant (DSI) kernel yields a Gram matrix diagonalizable by the fast Walsh–Hadamard transform (FWHT) [Fino and Algazi, 1976].
+The currently supported kernels and fast transforms are described below.
+RKHS Kernels
+We provide SI kernels and DSI kernels, including those of higher-order smoothness.
+We also maintain an interface to a number of standard kernels, such as squared exponential and Matérn kernels, whose implementations may also be found in the GPyTorch [^11] https://gpytorch.ai/ [Gardner et al., 2018] and the scikit-learn [^12] https://scikit-learn.org/ [Pedregosa et al., 2011] Python packages among others.
+SI kernels of arbitrary smoothness are well known and can be computed based on the Bernoulli polynomials [Kaarnioja et al., 2022, 2023, Cools et al., 2021, 2020, Sloan and Woźniakowski, 2001, Kuo et al., 2004].
+DSI kernels of order $1$ smoothness were derived in [Dick and Pillichshammer, 2005].
+In this paper we derive new higher-order DSI kernels whose RKHSs contain smooth functions.
+Fast Transforms
+We provide interfaces to the FFTBR, IFFTBR, and FWHT algorithms which all have $\mathcal{O}(n\log n)$ complexity in the number of points $n$.
+Our FFTBR and IFFT implementations use the FFT routines in the SciPy.
+An implementation of the FWHT is also available in the SymPy [^13] https://www.sympy.org [Meurer et al., 2017] Python package.
+1.2 Novel contributions and scope
+The novel contributions of this article are to:
+  1. 1.
+    Update QMCPy to be the first Python interface to support higher-order digital nets, higher-order digital net scrambling with either LMS or NUS, or Halton scrambling with either LMS and NUS.
+  2. 2.
+    Introduce the LDData repository, which contains collections of generating vectors and generating matrices in standardized formats, and provide a supported interface in QMCPy.
+  3. 3.
+    Add new tools to QMCPy which enable fast kernel computations. These tools include SI and DSI kernels of varying smoothness as well as the FFTBR, IFFTBR, and FWHT algorithms.
+  4. 4.
+    Derive novel DSI kernels of higher-order smoothness.
+Low order kernel forms appeared in [Baldeaux et al., 2012] as worst-case error bounds on QMC rules using higher-order polynomial lattices, but there they were not interpreted as DSI kernels.
+We also derive a new order $4$ smoothness DSI kernel whose form has not appeared elsewhere in the literature.
+  5. 5.
+    Provide efficient procedures to update the FFTBR, IFFTBR, or FWHT of a sequence when the number of points is doubled and the original points are not discarded.
+This article will focus on our novel contributions to equip QMCPy with broad support for randomized LD sequences and new tools for fast kernel methods.
+1.3 Outline
+The remainder of this article is organized as follows.
+Section 2 gives common notations.
+Section 3 describes two motivating problems using low discrepancy sequences: QMC methods (Section 3.1) and fast kernel computations (Section 3.2).
+Section 4 gives a few notes on our QMCPy implementation and the setup required to reproduce codes in this article.
+Section 5 details randomized rank-1 lattices (Section 5.1), digital sequences (Section 5.2), and Halton sequences (Section 5.3).
+Section 6 describes pairing SI kernels to rank-1 lattices (Section 6.1), DSI kernels to digital nets (Section 6.2), and the resulting fast kernel computations (Section 6.3).
+Section 7 provides numerical experiments showcasing the speed, accuracy, and versatility of QMCPy.
+Section 8 gives a brief conclusion.
+2 Notation
+Bold symbols will denote vectors which are assumed to be column vectors, e.g., ${\boldsymbol{x}}\in[0,1]^{d}$.
+Capital letters in serif font denote matrices, e.g., ${\mathsf{K}}\in{\mathbb{R}}^{n\times n}$.
+Lower case letters in serif font denote digits in a base $b$ expansion, e.g., $i=\sum_{t=0}^{m-1}{\mathsf{i}}_{t}b^{t}$ for $0\leq i<2^{m}$.
+This may be combined with bold notation when denoting the vector of base $b$ digits, e.g., ${\boldsymbol{\mathsf{i}}}=({\mathsf{i}}_{0},{\mathsf{i}}_{1},\dots,{\mathsf{i}}_{m-1})^{\intercal}$.
+Modulo is always taken component-wise, e.g., rank-1 lattices will use the notation ${\boldsymbol{x}}\mod 1=(x_{1}\mod 1,\dots,x_{d}\mod 1)^{\intercal}$ and digital nets will take all matrix operations to be carried modulo $b$.
+Permutations may be specified by vectors, e.g., $\pi:\{0,1,2\}\to\{0,1,2\}$ with $\pi(0)=2$, $\pi(1)=0$, and $\pi(2)=1$ may be denoted by $\pi=(2,0,1)$.
+3 Motivating problems
+3.1 QMC methods
+Monte Carlo (MC) and Quasi-Monte Carlo (QMC) methods approximate a high dimensional integral over the unit cube by the sample average of function evaluations at certain sampling locations:
+$\mu:=\int_{[0,1]^{d}}f({\boldsymbol{x}})\mathrm{d}{\boldsymbol{x}}\approx\frac{1}{n}\sum_{i=0}^{n-1}f({\boldsymbol{x}}_{i})=:{\widehat{\mu}}$. (1)
 Here $f:[0,1]^{d}\to{\mathbb{R}}$ is a given integrand and $\{{\boldsymbol{x}}_{i}\}_{i=0}^{n-1}\in[0,1]^{n\times d}$ is a point set.
-The integral on the left hand side may be viewed as taking the expectation ${\mathbb{E}}[f({\boldsymbol{X}})]$ where ${\boldsymbol{X}}$ is a standard uniform ${\boldsymbol{X}}\sim{\mathcal{U}}[0,1]^{d}$.
+The integral on the left-hand side may be viewed as taking the expectation ${\mathbb{E}}[f({\boldsymbol{X}})]$ where $\boldsymbol{X}$} is a standard uniform ${\boldsymbol{X}}\sim{\mathcal{U}}[0,1]^{d}$.
 Classic MC methods choose the sampling locations to be independent and identically distributed (IID) $d$-dimensional standard uniforms ${\boldsymbol{x}}_{0},\dots,{\boldsymbol{x}}_{n-1}\overset{\mathrm{IID}}{\sim}{\mathcal{U}}[0,1]^{d}$.
-IID-MC methods for (1) have a root-mean-squared-error (RMSE) of ${\mathcal{O}}(n^{-1/2})$.
-Figure 1.
-An independent identically distributed (IID) point set and low-discrepancy (quasi-random) point sets of size $n=2^{13}=8192$.
-Notice the quasi-random points more evenly fill the space than IID points, which leads to faster convergence of Quasi-Monte Carlo methods compared to IID Monte Carlo methods.
-Randomized lattices, digital nets in base $b=2$ (DN2), and Halton points are shown.
-Randomizations include shifts, linear matrix scramblings (LMS), digital shifts (DS), and nested uniform scramblings (NUS).
-Digital interlacing of order $\alpha$ is used to create higher-order randomized digital nets (HOα).
-[…]
-Quasi-Monte Carlo (QMC) methods (Niederreiter, 1992; Dick and Pillichshammer, 2010; Kroese et al., 2013; Dick et al., 2022; Lemieux, 2009; Sloan and Joe, 1994) replace IID point sets with low-discrepancy (LD) point sets which more evenly cover the unit cube $[0,1]^{d}$.
-For integrands with bounded variation, plugging LD point sets into ${\widehat{\mu}}$ (1) yields a worst-case error rate of ${\mathcal{O}}(n^{-1+\delta})$ with $\delta>0$ arbitrarily small.
-Some popular LD point sets are plotted in fig.
-1 including lattices, digital nets (including higher-order versions), and Halton point sets.
+IID MC methods for (1) have a root-mean-squared-error (RMSE) of ${\mathcal{O}}(n^{-1/2})$.
+Figure 1: An independent identically distributed (IID) point set and low-discrepancy (LD) point sets of size $n=2^{13}=8192$.
+Notice the LD points more evenly fill the space than IID points, which leads to faster convergence of Quasi-Monte Carlo methods compared to IID Monte Carlo methods.
+Randomized rank-1 lattices, digital nets in base $2$, and Halton points are shown.
+Randomizations include shifts, linear matrix scramblings (LMS), digital shifts (DS), digital permutations (DP), and nested uniform scramblings (NUS).
+Digital interlacing of order $\alpha$ is used to create higher-order randomized digital nets in base 22 (DNα).
+QMC methods [Niederreiter, 1992, Dick and Pillichshammer, 2010, Kroese et al., 2013, Dick et al., 2022, Lemieux, 2009, Sloan and Joe, 1994, Dick et al., 2013] replace IID point sets with LD point sets which more evenly cover the unit cube $0,1]^{d}$.
+For integrands with bounded variation, plugging LD point sets into (1) yields a worst-case error rate of ${\mathcal{O}}(n^{-1+\delta})$ with $\delta>0$ arbitrarily small.
+Some popular LD point sets are plotted in Figure 1 including randomized rank-1 lattices, base $2$ digital nets (including higher-order versions), and Halton point sets.
 Randomized Quasi-Monte Carlo (RQMC) uses randomized LD point sets to give improved convergence rates and enable practical error estimation.
 Specifically, if we again assume the integrand has bounded variation, then certain RQMC methods can achieve a RMSE of ${\mathcal{O}}(n^{-3/2+\delta})$.
-Moreover, let $\{{\boldsymbol{x}}_{i}^{1}\}_{i=0}^{n-1},\dots,\{{\boldsymbol{x}}_{i}^{R}\}_{i=0}^{n-1}$ denote $R$ IID randomizations of a low-discrepancy point set where typically $R$ is small e.g.
-$R=15$.
-Then the RQMC estimate
-(2) ${\widehat{\mu}}_{R}=\frac{1}{R}\sum_{r=1}^{R}\hat{\mu}^{r}\qquad\text{where}\quad\hat{\mu}^{r}=\frac{1}{n}\sum_{i=0}^{n-1}f({\boldsymbol{x}}_{i}^{r})$
-may be used to construct a practical (approximate) $100(1-\tau)\%$ confidence interval ${\widehat{\mu}}_{R}\pm t_{R-1,\tau/2}{\widehat{\sigma}}_{R}/\sqrt{R}$ for $\mu$ where ${\widehat{\sigma}}_{R}^{2}=R^{-1}(R-1)^{-1}\sum_{r=1}^{R}({\widehat{\mu}}^{r}-{\widehat{\mu}}_{R})^{2}$ and $t_{R-1,\tau/2}$ is the $\tau/2$ quantile of a Student distribution with $R-1$ degrees of freedom.
-RQMC estimates are detailed in (Owen, 2018, Chapter 17), and a recent comparison of RQMC confidence interval methods can be found in (L’Ecuyer et al., 2023).
-QMC error estimation is treated more broadly in (Owen, 2024), while (Clancy et al., 2014; Hickernell et al., 2017; Clancy et al., 2014) detail additional considerations for adaptive QMC algorithms.
-For lattices, randomization is typically done using random shifts modulo $1$.
-For digital nets, randomization is typically done using nested uniform scrambling (NUS) or the cheaper combination of linear matrix scrambling (LMS) with digital shifts / permutations (Owen, 2003, 2017; Owen and Pan, 2024).
+Moreover, let $\{{\boldsymbol{x}}_{1i}\}_{i=0}^{n-1},\dots,\{{\boldsymbol{x}}_{Ri}\}_{i=0}^{n-1}$ denote $R$ IID randomizations of an LD point set where typically $R$ is small, e.g., $R=15$.
+Then, following [L’Ecuyer et al., 2023] and [Owen, 2023, Chapter 17], the RQMC estimate
+${\widehat{\mu}}=\frac{1}{R}\sum_{r=1}^{R}{\widehat{\mu}}_{r}\qquad\text{where}\quad{\widehat{\mu}}_{r}=\frac{1}{n}\sum_{i=0}^{n-1}f({\boldsymbol{x}}_{ri})$ (2)
+gives a practical $100(1-\tau)\%$ confidence interval ${\widehat{\mu}}\pm t_{R-1,\tau/2}{\widehat{\sigma}}/\sqrt{R}$ for $\mu$ where ${\widehat{\sigma}}^{2}=(R-1)^{-1}\sum_{r=1}^{R}({\widehat{\mu}}_{r}-{\widehat{\mu}})^{2}$ and $t_{R-1,\tau/2}$ is the $\tau/2$ quantile of a Student’s-$t$ distribution with R1R-1 degrees of freedom.
+LABEL:code:genz_ex_1 will show how to code up such an estimate in QMCPy.
+For rank-1 lattices, randomization is typically done using random shifts modulo $1$.
+For digital nets, randomization is typically done using NUS or the cheaper combination of LMS with digital shifts / permutations [Matoušek, 1998, Owen, 2003, 2017, Owen and Pan, 2024].
 Higher-order LD point sets were designed to yield faster convergence for integrands with additional smoothness.
-For integrands with square integrable mixed partial derivatives up to order $\alpha>1$, plugging higher-order digital nets into ${\widehat{\mu}}$ (1) yields a worst-case error rate of ${\mathcal{O}}(n^{-\alpha+\delta})$ (Dick, 2008, 2009b, 2009a).
-RQMC using scrambled digital nets (via either NUS or LMS with digital permutations) combined with digital interlacing has been shown to achieve an RMSE of order ${\mathcal{O}}(n^{-\alpha-1/2+\delta})$ (Dick, 2011).
-1.2. Kernel computations
-Low-discrepancy (LD) point sets can also been used to accelerated many kernel computations.
-Let $\{{\boldsymbol{x}}_{i}\}_{i=0}^{n-1}$ be a point set, $K:[0,1]^{d}\times[0,1]^{d}\to{\mathbb{R}}$ be a symmetric positive definite (SPD) kernel, and ${\mathsf{K}}=\left(K({\boldsymbol{x}}_{i},{\boldsymbol{x}}_{k})\right)_{i,k=0}^{n-1}$ be the $n\times n$ SPD Gram matrix.
+For integrands with square integrable mixed partial derivatives up to order $\alpha>1$, plugging higher-order digital nets into $^{\widehat{\mu}}$ (1) yields a worst-case error rate of ${\mathcal{O}}(n^{-\alpha+\delta})$ [Dick, 2008, 2009b, 2009a].
+RQMC using higher-order digital nets with higher-order NUS or LMS scrambling has been shown to achieve an RMSE of order ${\mathcal{O}}(n^{-\alpha-1/2+\delta})$ [Dick, 2011].
+3.2 Fast Kernel computations
+Randomized LD point sets have also been used to accelerated kernel computations.
+Let $\{{\boldsymbol{x}}_{i}\}_{i=0}^{n-1}$ be a point set, $K:[0,1]^{d}\times[0,1]^{d}\to{\mathbb{R}}$ be a symmetric positive definite (SPD) kernel, and ${\mathsf{K}}=\left(K({\boldsymbol{x}}_{i},{\boldsymbol{x}}_{i^{\prime}})\right)_{i,i^{\prime}=0}^{n-1}$ be the $n\times n$ SPD Gram matrix of kernel evaluations at all pairs of points.
 Two motivating kernel computations are described below
-Discrepancy Computation:
-The error of the Monte Carlo approximation (1) can be bounded by the product two terms: a measure of variation of the function $f$ and a measure of the discrepancy of the point set $\{{\boldsymbol{x}}_{i}\}_{i=1}^{n-1}$.
-The most well known bound of this type is the Koksma-Hlawka inequality (Hickernell, 1998; Dick et al., 2013; Hickernell, 1999; Niederreiter, 1992).
-The discrepancy is frequently computed when designing and evaluating low-discrepancy point sets, see (Rusch et al., 2024) for a newer idea in this area where good point sets are generated using a neural networks trained with a discrepancy-based loss function.
+Discrepancy Computation
+The error of the approximation (1) can be bounded by the product of two terms: a measure of variation of the function $f$ and a measure of the discrepancy of the point set $\{{\boldsymbol{x}}_{i}\}_{i=0}^{n-1}$.
+The most well known bound of this type is the Koksma–Hlawka inequality [Hickernell, 1998a, Dick et al., 2013, Hickernell, 1999, Niederreiter, 1992].
+The discrepancy is frequently computed when designing and evaluating low-discrepancy point sets, see [Rusch et al., 2024] for a newer idea in this area where good point sets are generated using neural networks trained with a discrepancy-based loss function.
 Let us consider the more general setting of approximating the mean $\mu$ in (1) by a weighted cubature rule $\sum_{i=0}^{n-1}\omega_{i}f({\boldsymbol{x}}_{i})$.
-If $f$ is assumed to lie in the RKHS $H$, the discrepancy in the error bound of such a cubature rule takes the form
-$\int\int K({\boldsymbol{u}},{\boldsymbol{v}})\mathrm{d}{\boldsymbol{u}}\mathrm{d}{\boldsymbol{v}}-2\sum_{i=0}^{n-1}\omega_{i}\int K({\boldsymbol{u}},{\boldsymbol{x}}_{i})\mathrm{d}{\boldsymbol{u}}+\sum_{i,k=0}^{n-1}\omega_{i}\omega_{k}K({\boldsymbol{x}}_{i},{\boldsymbol{x}}_{k})$
-following (Hickernell, 1998).
-Here the integrals are understood to be over the unit cube $[0,1]^{d}$.
-Evaluating this discrepancy above requires computing ${\mathsf{K}}{\boldsymbol{\omega}}$ where ${\boldsymbol{\omega}}=\{\omega_{i}\}_{i=0}^{n-1}$.
-Moreover, the discrepancy is minimized for a given point set by setting the weights to ${\boldsymbol{\omega}}^{\star}={\mathsf{K}}^{-1}{\boldsymbol{\kappa}}$ where $\kappa_{i}=\int K({\boldsymbol{u}},{\boldsymbol{x}}_{i})\mathrm{d}{\boldsymbol{u}}$ for $0\leq i<n$.
-Kernel Interpolation:
-Suppose we would like to approximate $f:[0,1)^{d}\to{\mathbb{R}}$ given observations ${\boldsymbol{y}}=\{y_{i}\}_{i=0}^{n-1}$ of $f$ at $\{{\boldsymbol{x}}_{i}\}_{i=0}^{n-1}$ satisfying $y_{i}=f(x_{i})$.
+If $f$ is assumed to lie in the RKHS $H$ with kernel $K$, the discrepancy in the error bound of such a cubature rule takes the form
+$\int_{[0,1]^{d}}\int_{[0,1]^{d}}K({\boldsymbol{u}},{\boldsymbol{v}})\mathrm{d}{\boldsymbol{u}}\mathrm{d}{\boldsymbol{v}}-2\sum_{i=0}^{n-1}\omega_{i}\int_{[0,1]^{d}}K({\boldsymbol{u}},{\boldsymbol{x}}_{i})\mathrm{d}{\boldsymbol{u}}+\sum_{i,i^{\prime}=0}^{n-1}\omega_{i}\omega_{i^{\prime}}K({\boldsymbol{x}}_{i},{\boldsymbol{x}}_{i^{\prime}})$
+following [Hickernell, 1998a].
+Evaluating the discrepancy above requires computing ${\mathsf{K}}{\boldsymbol{\omega}}$ where ${\boldsymbol{\omega}}=\{\omega_{i}\}_{i=0}^{n-1}$.
+Moreover, this discrepancy is minimized for a given point set by setting the weights to ${\boldsymbol{\omega}}^{\star}={\mathsf{K}}^{-1}{\boldsymbol{\kappa}}$ where $\kappa_{i}=\int_{[0,1]^{d}}K({\boldsymbol{u}},{\boldsymbol{x}}_{i})\mathrm{d}{\boldsymbol{u}}$ for $0\leq i<n$.
+Kernel Interpolation
+Suppose we would like to approximate $f:[0,1)^{d}\to{\mathbb{R}}$ given observations ${\boldsymbol{y}}=\{y_{i}\}_{i=0}^{n-1}$ of $f$ at $\{{\boldsymbol{x}}_{i}\}_{i=0}^{n-1}$ satisfying $y_{i}=f(x_{i})$ for $0\leq i<n$.
 Then a kernel interpolant approximates $f$ by ${\widehat{f}}\in H$ where
@@
 and ${\boldsymbol{\omega}}={\mathsf{K}}^{-1}{\boldsymbol{y}}$.
-The above kernel interpolant may be reinterpreted as the posterior mean in Gaussian process regression, see (Williams and Rasmussen, 2006).
-Fitting a Gaussian process regression model often includes optimizing a kernel’s hyperparameters, which also may be done by computing ${\mathsf{K}}{\boldsymbol{\omega}}$ and ${\mathsf{K}}^{-1}{\boldsymbol{y}}$.
-Underlying these problems, and many others, is the requirement to compute the matrix-vector product ${\mathsf{K}}{\boldsymbol{y}}$ or solve the linear system ${\mathsf{K}}\boldsymbol{\omega}={\boldsymbol{y}}$ for some length $n$ vector ${\boldsymbol{y}}$.
+The above kernel interpolant may be reinterpreted as the posterior mean of a Gaussian process regression model, see [Williams and Rasmussen, 2006].
+Fitting such a Gaussian process often includes optimizing a kernel’s hyperparameters, which may also be done by computing ${\mathsf{K}}{\boldsymbol{\omega}}$ and ${\mathsf{K}}^{-1}{\boldsymbol{y}}$.
+Underlying these problems, and many others, is the requirement to compute the matrix-vector product ${\mathsf{K}}{\boldsymbol{y}}$ or solve the linear system ${\mathsf{K}}^{-1}{\boldsymbol{y}}$.
 The standard cost of these computations are ${\mathcal{O}}(n^{2})$ and ${\mathcal{O}}(n^{3})$ respectively.
 One method to reduce these high costs is to induce structure into $\mathsf{K}$}.
-Structure-inducing methods pairing lattices and digital nets with special kernels were proposed in (Zeng et al., 2009) and shown to reduce the cost of both computations to ${\mathcal{O}}(n\log n)$.
-Specifically, when a lattice point set $\{{\boldsymbol{x}}_{i}\}_{i=0}^{n-1}$ in linear order is paired with a shift-invariant kernel, ${\mathsf{K}}$ becomes circulant and computations can be performed performed using the fast Fourier transforms (FFT) and inverse FFT (IFFT).
-Similarly, when a base $2$ digital net $\{{\boldsymbol{x}}_{i}\}_{i=0}^{n-1}$ is paired with a digitally-shift-invariant kernel, ${\mathsf{K}}$ becomes nested-block-Toeplitz and these computations can be performed using the Fast Walsh-Hadamard Transforms (FWHT) (Fino and Algazi, 1976).
-Such methods have recently been used to accelerate Bayesian cubature (Jagadeeswaran and Hickernell, 2019, 2022; Rathinavel, 2019) and solving PDEs with random coefficients (Kaarnioja et al., 2022, 2023).
-We provide novel methods which exploit past solutions to ${\mathsf{K}}{\boldsymbol{y}}$ or ${\mathsf{K}}^{-1}{\boldsymbol{y}}$ when doubling the sample size from the LD sequence.
-This requires lattices are generated in natural order so we do not need to reorder outputs after doubling the sample size.
-We show that the same speedups can be achieved for lattices in natural order using bit-reversed FFT and bit-reversed IFFT algorithms.
-These bit-reversed algorithms simply skip the initial and final bit-reversal steps in the decimation-in-time FFT and IFFT respectively.
-We provide implementations of the bit-reversed FFT, bit-reversed IFFT, and FWHT.
-Shift-invariant kernels of arbitrary smoothness are well known and can be computed based on the Bernoulli polynomials (Kaarnioja et al., 2022, 2023; Cools et al., 2021, 2020; Sloan and Woźniakowski, 2001; Kuo et al., 2004).
-Until now, digitally-shift-invariant kernels were only known for order $1$ smoothness (Dick and Pillichshammer, 2005).
-We present higher-order digitally-shift-invariant kernels whose corresponding RKHSs contain functions of arbitrary smoothness.
-We derive closed forms of low order kernels (up to order $4$ smoothness) in base $b=2$, with the order $4$ kernel not appearing elsewhere in the literature.
-These kernels are also of interest for computing the worst-case error of higher-order polynomial-lattice rules (Baldeaux et al., 2012), although we do not explore this connection here.
-1.3. Outline
-The remainder of this article is organized as follows.
-First, we describe our implementation in section 2.
-Section 3 discusses notation.
-Section 4 details rank $1$ lattices, digital sequences, and Halton sequences along with their randomization routines.
-Section 5 describes accelerated kernel computation with LD sequences, matching kernels, and fast transform algorithms.
-Section 6 provides numerical experiments.
-Finally, section 7 gives a brief conclusion and outlines directions for future work.
-2. Implementation
-We have implemented this work into QMCPy [^10] https://qmcpy.readthedocs.io version 1.6, installable with pip install -U qmcpy.
-The architecture and philosophy underlying QMCPy are described in (Choi et al., 2022b, a).
-To follow along with code snippets presented in this article, please import QMCPy and NumPy (Oliphant, 2006) via
+Zeng et al.
+[2009, 2006] proposed structure-inducing methods which use certain kernels $K$ and points $\{{\boldsymbol{x}}_{i}\}_{i=0}^{n-1}$ which enable both computations to be done at only ${\mathcal{O}}(n\log n)$ cost.
+Two such pairings exist:
+  1. 1.
+    When a rank-1 lattice point set $\{{\boldsymbol{x}}_{i}\}_{i=0}^{n-1}$ in linear order is paired with a SI kernel, $\mathsf{K}$} is circulant and thus diagonalizable by the FFT and IFFT.
+Using lattices in the extensible radical inverse order requires using the bit-reversed fast transforms FFTBR and IFFTBR.
+  2. 2.
+    When a base $2$ digital net $\{{\boldsymbol{x}}_{i}\}_{i=0}^{n-1}$ in radical inverse order is paired with a DSI kernel, $\mathsf{K}$} becomes RSBT (recursive symmetric block Toeplitz) and thus diagonalizable by the FWHT.
+4 Implementation
+The features discussed in this article are available in QMCPy version 2.0 or later, which may be readily installed using the command pip install -U qmcpy.
+To follow along with the code snippets in this paper, first
+Listing 1: Import the QMCPy and NumPy Python packages to be used in examples throughout this paper.
 import qmcpy as qp
 import numpy as np
-These code snippets and code producing the plots in this article are available in notebook form [^11] https://github.com/QMCSoftware/QMCSoftware/blob/master/demos/ld_randomizations_and_higher_order_nets.ipynb.
-Most of the QMCPy objects and methods presented here are wrappers around the lower level QMCToolsCL [^12] https://qmcsoftware.github.io/QMCToolsCL/ package.
-The Python functions in QMCToolsCL often call lower level C implementations for improved performance.
-QMCToolsCL interfaces for other languages which support C extensions is a direction of future work.
-3. Notation
-Bold symbols denote vectors which are assumed to be column vectors e.g.
-${\boldsymbol{x}}\in[0,1]^{d}$.
-Capital letters in serif font denote matrices e.g.
-${\mathsf{K}}\in{\mathbb{R}}^{n\times n}$.
-Lower case letters in serif font denote digits in a base $b$ expansion e.g.
-$i=\sum_{t=0}^{m-1}{\mathsf{i}}_{t}b^{t}$ for $0\leq i<2^{m}$.
-This may be combined with bold notation when denoting the vector of base $b$ digits e.g.
-${\boldsymbol{\mathsf{i}}}=({\mathsf{i}}_{0},{\mathsf{i}}_{1},\dots,{\mathsf{i}}_{m-1})^{\intercal}$.
-Modulo is always taken component-wise e.g.
-${\boldsymbol{x}}\mod 1=(x_{1}\mod 1,\dots,x_{d}\mod 1)^{\intercal}$.
-Permutations may be specified by vectors e.g.
-$\pi:\{0,1,2\}\to\{0,1,2\}$ with $\pi(0)=2$, $\pi(1)=0$, and $\pi(2)=1$ may be denoted by $\pi=(2,0,1)$.
-4. Low-discrepancy sequences and randomization routines
+All codes and plots are reproducible in a public notebook [^14] https://qmcsoftware.github.io/QMCSoftware/demos/talk_paper_demos/ACMTOMS_Sorokin_2025/acm_toms_sorokin_2025/.
+Most of the QMCPy functions presented in this paper are wrappers around efficient C implementations in our QMCToolsCL [^15] https://qmcsoftware.github.io/QMCToolsCL/ package.
+These low level routines could be used to rapidly build similar interfaces in other programming languages which support C extensions.
+5 Randomized low-discrepancy sequences
 For a fixed base $b\in{\mathbb{N}}$, write $i\in{\mathbb{N}}_{0}$ as $i=\sum_{t=0}^{\infty}{\mathsf{i}}_{t}b^{t}$ so ${\mathsf{i}}_{t}$ is the $t^{\mathrm{th}}$ digit in the base $b$ expansion of $i$.
@@
 ${\boldsymbol{D}}_{m}(i)=({\mathsf{i}}_{0},{\mathsf{i}}_{1},\dots,{\mathsf{i}}_{m-1})^{\intercal}.$
-For ${\boldsymbol{\mathsf{i}}}={\boldsymbol{D}}_{m}(i)$ we use the notation
-$F_{m}({\boldsymbol{\mathsf{i}}})=\sum_{t=1}^{m}{\mathsf{i}}_{t-1}b^{-t}$
-to flip the digits of $i$ about the decimal and convert back to a floating point number in $[0,1)$. Finally, let
-(3) $v(i)=\sum_{t=1}^{\infty}{\mathsf{i}}_{t-1}b^{-t}$
-so that $v(i)=F_{m}({\boldsymbol{D}}_{m}(i))$ when $i<b^{m}$, which is always the case in this article. The Van der Corput sequence in base $b$ is $(v(i))_{i\geq 0}$.
+For ${\boldsymbol{\mathsf{i}}}={\boldsymbol{D}}_{m}(i)$, we denote the radical inverse of $i$ by
+$F_{m}({\boldsymbol{\mathsf{i}}})=\sum_{t=1}^{m}{\mathsf{i}}_{t-1}b^{-t}\in[0,1)$
+Finally, let
+$v(i)=\sum_{t=1}^{\infty}{\mathsf{i}}_{t-1}b^{-t}$ (3)
+so that $v(i)=F_{m}({\boldsymbol{D}}_{m}(i))$ when $i<b^{m}$, which is always the case in this article. The van der Corput sequence in base $b$ is $\{v(i)\}_{i\geq 0}$.
 […]
-$x$$z$shifted digital net
-$x$$z$permuted digital net
-$x$$z$NUS digital net
-Figure 2.
-Low-discrepancy randomization routines in dimension $d=1$.
-Each vertical line is a unit interval with 00 at the bottom and $1$ at the top.
+xxzzdigitally shifted DN
+xxzzpermuted DN
+xxzzNUS DN
+Figure 2: Low-discrepancy randomization routines in dimension $d=1$.
+Each vertical line is a unit interval with 0 at the bottom and 11 at the top.
 A given interval is partitioned at the horizontal ticks extending to the right, and then rearranged following the arrows to create the partition of the right interval as shown by the ticks extending to the left.
-For the shifted lattice and digitally shifted digital net, when the arrow hits a horizontal gray bar it is wrapped around to the next gray bar below.
-See for example the dotted line in the shifted digital net.
+For the shifted rank-1 lattice and digitally shifted digital net, when the arrow hits a horizontal gray bar it is wrapped around to the next gray bar below.
+See for example the dotted line in the digitally shifted digital net.
 The lattice shift is $\Delta=0.227$.
-All digital nets use base $b=3$ and $t_{\mathrm{max}}=2$ digits of precision.
+All digital nets (DNs) use base $b=3$ and $t_{\mathrm{max}}=2$ digits of precision.
 The digital shift is ${\boldsymbol{\Delta}}=(2,1)^{\intercal}$.
 Dropping the $j$ subscript for dimension, the digital permutations in the third panel are $\pi_{1}=(2,0,1)$ and $\pi_{2}=(2,1,0)$.
-Notice $\pi_{1}$ is equivalent to a digital shift by $2$, but $\pi_{2}$ cannot be written as a digital shift.
+Notice $\pi_{1}$ is equivalent to a digital shift by 22, but $\pi_{2}$ cannot be written as a digital shift.
 The nested uniform scramble (NUS) has digital permutations $\pi=(2,0,1)$, $\pi_{0}=(2,1,0)$, $\pi_{1}=(0,1,2)$, and $\pi_{2}=(1,2,0)$.
-Notice the permuted digital net in third panel has permutations depending only the current digit in the base $b$ expansion.
+Notice the permuted digital net in the third panel has permutations depending only on the current digit in the base $b$ expansion.
 In contrast, the full NUS scrambling in the fourth panel has permutations which depend on all previous digits in the base $b$ expansion.
-4.1. Rank 1 lattices
+5.1 Rank-1 lattices
 Consider a fixed generating vector ${\boldsymbol{g}}\in{\mathbb{N}}^{d}$ and fixed prime base $b$. Then we define the lattice sequence
-${\boldsymbol{z}}_{i}=v(i){\boldsymbol{g}}\mod 1,\qquad i\geq 0.$
-If $n=b^{m}$ for some $m\in{\mathbb{N}}_{0}$, then the lattice point set $\{{\boldsymbol{z}}_{i}\}_{i=0}^{n-1}\in[0,1)^{n\times d}$ is equivalent to $\{i/n{\boldsymbol{g}}\mod 1\}_{i=0}^{n-1}$ where we say the former is in natural order while the latter is in linear order.
-Shifted lattice:
+${\boldsymbol{z}}_{i}=v(i){\boldsymbol{g}}\mod 1,\qquad i\geq 0.$ (4)
+If $n=b^{m}$ for some $m\in{\mathbb{N}}_{0}$, then the lattice point set $\{{\boldsymbol{z}}_{i}\}_{i=0}^{n-1}\in[0,1)^{n\times d}$ is equivalent to $\{{\boldsymbol{g}}i/n\mod 1\}_{i=0}^{n-1}$ where we say the former is in radical inverse order while the latter is in linear order.
+When $n$ is not of the form $b^{m}$, then linear and radical inverse order will not generate the same lattice point set.
+Shifted lattice
 For a shift ${\boldsymbol{\Delta}}\in[0,1)^{d}$, we define the shifted point set
 ${\boldsymbol{x}}_{i}=({\boldsymbol{z}}_{i}+{\boldsymbol{\Delta}})\mod 1$
-where ${\boldsymbol{z}}_{i}$ are lattice point set.
-The shift operation mod 1 is visualized in fig.
-2.
+where ${\boldsymbol{z}}_{i}$ denotes the unrandomized lattice point in (4).
+The shift operation mod 1 is visualized in Figure 2.
 Randomized lattices use ${\boldsymbol{\Delta}}\sim{\mathcal{U}}[0,1]^{d}$.
 In the following code snippet we generate $R$ independently shifted lattices with shifts ${\boldsymbol{\Delta}}_{1},\dots,{\boldsymbol{\Delta}}_{R}\overset{\mathrm{IID}}{\sim}{\mathcal{U}}[0,1]^{d}$.
+Listing 2: Generate randomized lattices.
 lattice = qp.Lattice(
 dimension = 52,
-randomize = "shift", # for unrandomized lattice set randomize = False
+randomize = "shift", # for unrandomized set randomize = None
 replications = 16, # R
-order = "natural", # also supports "linear" and "Gray" for Gray code
+order = "radical inverse", # also supports "linear"
 seed = None, # pass integer seed for reproducibility
-generating_vector = "kuo.lattice-33002-1024-1048576.9125.txt")
-x = lattice(2**16) # a numpy.ndarray with shape 16 x 65536 x 52
-Here we have used a generating vector from (Cools et al., 2006) which is stored in a standardized format in the LDData repository.
-Other generating vectors from LDData may be used by passing in a file name from https://github.com/QMCSoftware/LDData/tree/main/lattice/ or by passing an explicit array.
-4.2. Digital nets
-Consider a fixed prime base $b$ and generating matrices ${\mathsf{C}}_{1},\dots\mathsf{C}_{d}\in\{0,...,b-1\}^{t_{\mathrm{max}}\times m}$ where $m\in{\mathbb{N}}$ is fixed and $b$ is a given prime base.
+generating_vector = "mps.exod2_base2_m20_CKN.txt")
+x = lattice(2**16) # a numpy.ndarray with shape (16, 65536, 52)
+Here we have used a generating vector from [Cools et al., 2006] which is stored in a standardized format in the LDData repository.
+Other generating vectors from LDData may be used by passing in a file name from https://github.com/QMCSoftware/LDData/tree/main/lattice/ or by passing in an explicit array.
+5.2 Digital nets
+Consider a fixed prime base $b$ and generating matrices ${\mathsf{C}}_{1},\dots\mathsf{C}_{d}\in\{0,...,b-1\}^{t_{\mathrm{max}}\times m}$ where $t_{\mathrm{max}},m\in{\mathbb{N}}$ are fixed and $b$ is a given prime base.
 One may relax the assumption that $b$ is prime, but we do not consider that here or in the implementation.
-The first $n=b^{m}$ points of a digital sequence form a digital net $\{{\boldsymbol{z}}_{i}\}_{i=1}^{b^{m}-1}\in[0,1)^{b^{m}\times d}$ in natural order.
-For $1\leq j\leq d$ and $0\leq i<b^{m}$ with digits ${\boldsymbol{\mathsf{i}}}={\boldsymbol{D}}_{m}(i)$,
-${\boldsymbol{\mathsf{z}}}_{ij}={\mathsf{C}}_{j}{\boldsymbol{\mathsf{i}}}\mod b\qquad\text{and}\qquad z_{ij}=F_{t_{\mathrm{max}}}({\boldsymbol{\mathsf{z}}}_{ij})$
-where ${\boldsymbol{\mathsf{z}}}_{ij}$ is the base $b$ expansion vector of $z_{ij}$.
-Digitally shifted digital net:
+The first $n=b^{m}$ points of a digital sequence form a digital net $\{{\boldsymbol{z}}_{i}\}_{i=0}^{b^{m}-1}\in[0,1)^{b^{m}\times d}$ where, in radical inverse order, for $1\leq j\leq d$ and $0\leq i<b^{m}$ with base $b$ digit vector ${\boldsymbol{\mathsf{i}}}={\boldsymbol{D}}_{m}(i)$,
+$z_{ij}=F_{t_{\mathrm{max}}}({\boldsymbol{\mathsf{z}}}_{ij})\qquad\text{with}\qquad{\boldsymbol{\mathsf{z}}}_{ij}={\mathsf{C}}_{j}{\boldsymbol{\mathsf{i}}}\mod b.$
+Digitally-shifted digital net
 Similar to lattices, one may apply a shift ${\mathsf{\Delta}}\in[0,1)^{t_{\mathrm{max}}\times d}$ to the digital net to get a digitally-shifted digital net $\{{\boldsymbol{x}}_{i}\}_{i=0}^{b^{m}-1}$ where
@@
 and ${\boldsymbol{\Delta}}_{j}$ is the $j^{\mathrm{th}}$ column of $\mathsf{\Delta}$}.
-Randomly shifted digital nets use ${\mathsf{\Delta}}\overset{\mathrm{IID}}{\sim}{\mathcal{U}}\{0,\dots,b-1\}^{t_{\mathrm{max}}\times d}$ i.e.
-is each digit is chosen uniformly from $\{0,\dots,b-1\}$.
-Digitally permuted digital net:
-In what follows will denote permutations of $\{0,\dots,b-1\}$ by $\pi$. Suppose we are given a set of permutations
+Randomly shifted digital nets use ${\mathsf{\Delta}}\overset{\mathrm{IID}}{\sim}{\mathcal{U}}\{0,\dots,b-1\}^{t_{\mathrm{max}}\times d}$, i.e., each digit is chosen uniformly from $\{0,\dots,b-1\}$.
+Digitally permuted digital net
+In what follows we will denote permutations of $\{0,\dots,b-1\}$ by $\pi$. Suppose we are given a set of permutations
 ${\mathsf{\Pi}}=\{\pi_{j,t}:\quad 1\leq j\leq d,\quad 0\leq t<t_{\mathrm{max}}\}.$
-Then we may construct the digitally permuted digital net $\{{\boldsymbol{x}}_{i}\}_{i=1}^{b^{m}-1}$ where
+Then we may construct the digitally permuted digital net $\{{\boldsymbol{x}}_{i}\}_{i=0}^{b^{m}-1}$ where
 $x_{ij}=F_{t_{\mathrm{max}}}((\pi_{j,0}({\mathsf{z}}_{ij0}),\pi_{j,1}({\mathsf{z}}_{ij1}),\dots,\pi_{j,t_{\mathrm{max}}-1}({\mathsf{z}}_{ij(t_{\mathrm{max}}-1)}))^{\intercal}).$
 Randomly permuted digital nets use independent permutations chosen uniformly over all permutations of $\{0,\dots,b-1\}$.
-Nested Uniform Scrambling (NUS):
-NUS is often called Owen scrambling for its conception in (Owen, 1995). As before, $\pi$ values denote permutations of $\{0,\dots,b-1\}$. Now suppose we are given a set of permutations
+Nested Uniform Scrambling (NUS)
+NUS is often called Owen scrambling for its conception in [Owen, 1995]. As before, $\pi$ denotes permutations of $\{0,\dots,b-1\}$. Now suppose we are given a set of permutations
 ${\mathcal{P}}=\{\pi_{j,{\mathsf{v}}_{1}\cdots{\mathsf{v}}_{t}}:\quad 1\leq j\leq d,\quad 0\leq t<t_{\mathrm{max}},\quad{\mathsf{v}}_{k}\in\{0,\dots,b-1\}\text{ for }0\leq k\leq t\}.$
-Then a nested uniform scrambling of a digital net is $\{{\boldsymbol{x}}_{i}\}_{i=1}^{b^{m}-1}$ where
+Then a nested uniform scrambling of a digital net is $\{{\boldsymbol{x}}_{i}\}_{i=0}^{b^{m}-1}$ where
 $x_{ij}=F_{t_{\mathrm{max}}}((\pi_{j,}({\mathsf{z}}_{ij0}),\pi_{j,{\mathsf{z}}_{ij0}}({\mathsf{z}}_{ij1}),\pi_{j,{\mathsf{z}}_{ij0}{\mathsf{z}}_{ij1}}({\mathsf{z}}_{ij2}),\dots,\pi_{j,{\mathsf{z}}_{ij0}{\mathsf{z}}_{ij1}\cdots{\mathsf{z}}_{ij(t_{\mathrm{max}}-2)}}({\mathsf{z}}_{ij(t_{\mathrm{max}}-1)}))^{\intercal}).$
@@
 $\lvert{\mathcal{P}}\rvert=d(1+b+b^{2}+\dots+b^{t_{\mathrm{max}}-1})=d\frac{b^{t_{\mathrm{max}}}-1}{b-1},$
-our implementation cannot afford to generated all permutations apriori.
+our implementation cannot afford to generate all permutations a priori.
 Instead, permutations are generated and stored only as needed.
 As with digitally-permuted digital nets, NUS uses independent uniform random permutations.
-The randomization routines described above are visualized in fig. 2.
+The randomization routines described above are visualized in Figure 2.
 Linear matrix scrambling (LMS) is a computationally cheaper but less complete version of NUS which has proven sufficient for many practical problems.
@@
 ${\widetilde{\mathsf{C}}}_{j}={\mathsf{S}}_{j}{\mathsf{C}}_{j}\mod b$
-for $j=1,\dots,d$. Following (Owen, 2003), let us denote elements in $\{1,\dots,b\}$ by $h$ and elements in $\{0,\dots,b\}$ by $g$. Then common structures for ${\mathsf{S}}_{j}$ include
+for $j=1,\dots,d$. Following [Owen, 2003], let us denote elements in $\{1,\dots,b-1\}$ by $h$ and elements in $\{0,\dots,b-1\}$ by $g$. Then common structures for ${\mathsf{S}}_{j}$ include
 $\begin{pmatrix}h_{11}&0&0&0&\dots\\ g_{21}&h_{22}&0&0&\dots\\ g_{31}&g_{32}&h_{33}&0&\dots\\ g_{41}&g_{42}&g_{43}&h_{44}&\dots\\ \vdots&\vdots&\vdots&\vdots&\ddots\end{pmatrix},\begin{pmatrix}h_{1}&0&0&0&\dots\\ g_{2}&h_{1}&0&0&\dots\\ g_{3}&g_{2}&h_{1}&0&\dots\\ g_{4}&g_{3}&g_{2}&h_{1}&\dots\\ \vdots&\vdots&\vdots&\vdots&\ddots\end{pmatrix},\quad\text{and}\quad\begin{pmatrix}h_{1}&0&0&0&\dots\\ h_{1}&h_{2}&0&0&\dots\\ h_{1}&h_{2}&h_{3}&0&\dots\\ h_{1}&h_{2}&h_{3}&h_{4}&\dots\\ \vdots&\vdots&\vdots&\vdots&\ddots\end{pmatrix}$
-which corresponds to Matousek’s linear scrambling (Matoušek, 1998), Tezuka’s $i$-binomial scrambling (Tezuka, 2002), and Owen’s striped LMS (Owen, 2003) (not to be confused with NUS which is often called Owen scrambling).
-Random LMS chooses $g$ and $h$ values all independently and uniformly from $\{1,\dots,b\}$ and $\{0,\dots,b-1\}$ respectively.
-These scrambling are described in more detail and connected to NUS in (Owen, 2003).
+which corresponds to Matoušek’s linear scrambling [Matoušek, 1998], Tezuka’s $i$-binomial scrambling [Tezuka, 2002], and Owen’s striped LMS [Owen, 2003] (not to be confused with NUS which is often called Owen scrambling).
+Random LMS chooses $g$ and $h$ values all independently and uniformly from $\{1,\dots,b-1\}$ and $\{0,\dots,b-1\}$ respectively.
+Owen [2003] analyzes these scramblings and their connection to NUS.
 Digital interlacing enables the construction of higher-order digital nets.
-Denote the digitally interlaced version of ${\mathsf{C}}_{1},{\mathsf{C}}_{2},\dots\in\{0,\dots,b\}^{t_{\mathrm{max}}\times m}$ with interlacing factor $\alpha\in{\mathbb{N}}$ by ${\widetilde{\mathsf{C}}}_{1},{\widetilde{\mathsf{C}}}_{2},\dots\in\{0,\dots,b\}^{\alpha t_{\mathrm{max}}\times m}$.
-Then we have ${\widetilde{\mathsf{C}}}_{jtk}=C_{{\widetilde{j}},\tt,k}$ where ${\widetilde{j}}=\alpha(j-1)+(t\mod\alpha)+1$ and $\tt=\lfloor t/\alpha\rfloor$ for $j\geq 1$ and $0\leq t<\alpha t_{\mathrm{max}}$ and $1\leq k\leq m$.
-For example, with $m=2$, $t_{\mathrm{max}}=2$, and $\alpha=2$ we have the following
-${\mathsf{C}}_{1}=\begin{pmatrix}c_{101}&c_{102}\\ c_{111}&c_{112}\end{pmatrix}\qquad{\mathsf{C}}_{2}=\begin{pmatrix}c_{201}&c_{202}\\ c_{211}&c_{212}\end{pmatrix}\qquad{\mathsf{C}}_{3}=\begin{pmatrix}c_{301}&c_{302}\\ c_{311}&c_{312}\end{pmatrix}\qquad{\mathsf{C}}_{4}=\begin{pmatrix}c_{401}&c_{402}\\ c_{411}&c_{412}\end{pmatrix}\qquad\dots$
-${\widetilde{\mathsf{C}}}_{1}=\begin{pmatrix}c_{101}&c_{102}\\ c_{201}&c_{202}\\ c_{111}&c_{112}\\ c_{211}&c_{212}\end{pmatrix}\qquad{\widetilde{\mathsf{C}}}_{2}=\begin{pmatrix}c_{301}&c_{302}\\ c_{401}&c_{402}\\ c_{311}&c_{312}\\ c_{411}&c_{412}\end{pmatrix}\qquad\dots$
-Higher-order NUS requires NUS be applied to the digital net point sets generated by ${\mathsf{C}}_{1},\dots,{\mathsf{C}}_{\alpha d}$ and then interlacing be performed to the digits $\{{\boldsymbol{\mathsf{z}}}_{i1}\}_{i=0}^{b^{m}-1},\dots,\{{\boldsymbol{\mathsf{z}}}_{i,\alpha d}\}_{i=0}^{b^{m}-1}$ (Dick, 2011).
-For higher-order LMS, we apply LMS to the generating matrices ${\mathsf{C}}_{1},\dots,{\mathsf{C}}_{\alpha d}$, then interlace them, then generate the digital net.
-As we shown in the numerical experiments in section 6, LMS is significantly faster than NUS (especially for higher-order nets) while still achieving higher-order rates of RMSE convergence.
+For integer order $\alpha\geq 1$, interlacing ${\mathsf{A}}_{1},{\mathsf{A}}_{2},\dots\in\{0,\dots,b\}^{t_{\mathrm{max}}\times m}$ gives $\widehat{{\mathsf{A}}}_{1},\widehat{{\mathsf{A}}}_{2},\dots\in\{0,\dots,b\}^{\alpha t_{\mathrm{max}}\times m}$ satisfying $\widehat{{\mathsf{A}}}_{jtk}=A_{\widehat{j},\widehat{t},k}$ where $\widehat{j}=\alpha(j-1)+(t\mod\alpha)+1$ and $\widehat{t}=\lfloor t/\alpha\rfloor$ for $j\geq 1$ and $0\leq t<\alpha t_{\mathrm{max}}$ and $1\leq k\leq m$.
+For example, with $m=2$, $t_{\mathrm{max}}=2$, and $\alpha=2$ we have
+${\mathsf{A}}_{1}=\begin{pmatrix}a_{101}&a_{102}\\ a_{111}&a_{112}\end{pmatrix}\qquad{\mathsf{A}}_{2}=\begin{pmatrix}a_{201}&a_{202}\\ a_{211}&a_{212}\end{pmatrix}\qquad{\mathsf{A}}_{3}=\begin{pmatrix}a_{301}&a_{302}\\ a_{311}&a_{312}\end{pmatrix}\qquad{\mathsf{A}}_{4}=\begin{pmatrix}a_{401}&a_{402}\\ a_{411}&a_{412}\end{pmatrix}\qquad\dots$
+$\widehat{{\mathsf{A}}}_{1}=\begin{pmatrix}a_{101}&a_{102}\\ a_{201}&a_{202}\\ a_{111}&a_{112}\\ a_{211}&a_{212}\end{pmatrix}\qquad\widehat{{\mathsf{A}}}_{2}=\begin{pmatrix}a_{301}&a_{302}\\ a_{401}&a_{402}\\ a_{311}&a_{312}\\ a_{411}&a_{412}\end{pmatrix}\qquad\dots$
+Without scrambling, higher-order digital nets may be directly generated from $\widehat{{\mathsf{C}}}_{1},\dots,\widehat{{\mathsf{C}}}_{d}$, the interlaced generating matrices resulting from interlacing ${\mathsf{C}}_{1},\dots,{\mathsf{C}}_{\alpha d}$.
+Higher-order NUS [Dick, 2011] requires generating a digital net from ${\mathsf{C}}_{1},\dots,{\mathsf{C}}_{\alpha d}$, applying NUS to the resulting $\alpha d$ dimensional digital net, and then interlacing the resulting digit matrices $\{{\boldsymbol{\mathsf{z}}}_{i,1}^{\intercal}\}_{i=0}^{b^{m}-1},\dots,\{{\boldsymbol{\mathsf{z}}}_{i,\alpha d}^{\intercal}\}_{i=0}^{b^{m}-1}\in\{0,\dots,b\}^{t_{\mathrm{max}}\times m}$.
+For higher-order LMS, one applies LMS to the generating matrices ${\mathsf{C}}_{1},\dots,{\mathsf{C}}_{\alpha d}$ to get ${\widetilde{\mathsf{C}}}_{1},\dots,{\widetilde{\mathsf{C}}}_{\alpha d}$, then interlaces ${\widetilde{\mathsf{C}}}_{1},\dots,{\widetilde{\mathsf{C}}}_{\alpha d}$ to get $\widehat{{\widetilde{\mathsf{C}}}}_{1},\dots,\widehat{{\widetilde{\mathsf{C}}}}_{d}$, then generates the digital net from $\widehat{{\widetilde{\mathsf{C}}}}_{1},\dots,\widehat{{\widetilde{\mathsf{C}}}}_{d}$.
+As we show in the numerical experiments in Section 7, LMS is significantly faster than NUS (especially for higher-order nets) while still achieving higher-order rates of RMSE convergence.
 A subtle difference between the above presentation and practical implementation is that $t_{\mathrm{max}}$ may change with randomization in practice.
@@
 Then we should generate ${\boldsymbol{\Delta}}\in\{0,\dots,b-1\}^{t_{\mathrm{max}}\times d}$ with $t_{\mathrm{max}}=64$ and treat ${\mathsf{C}}_{j}$ as $t_{\mathrm{max}}\times t_{\mathrm{max}}$ matrices with appropriate rows and columns zeroed out.
-Gray code ordering of digital nets enables computing the next point ${\boldsymbol{x}}_{i+1}$ from ${\boldsymbol{x}}_{i}$ by only adding a single column from each generating matrix rather.
-Specifically, the $q^{\mathrm{th}}$ column of each generating matrix gets digitally added to the previous point where $q-1$ is the index of the only digit to be changed in Gray code ordering.
+Gray code ordering of digital nets enables computing the next point ${\boldsymbol{x}}_{i+1}$ from ${\boldsymbol{x}}_{i}$ by only adding a single column from each generating matrix.
+Specifically, the $q^{\mathrm{th}}$ column of each generating matrix gets digitally added to the previous point where q1q-1 is the index of the only digit to be changed in Gray code ordering.
 Gray code orderings for $b=2$ and $b=3$ are shown in Table 1.
 $i$ $i_{2}$ Gray code $i_{2}$ $i_{3}$ Gray code $i_{3}$
-00 $0000_{2}$ $0000_{2}$ $00_{3}$ $00_{3}$
-$1$ $0001_{2}$ $0001_{2}$ $01_{3}$ $01_{3}$
-$2$ $0010_{2}$ $0011_{2}$ $02_{3}$ $02_{3}$
-$3$ $0011_{2}$ $0010_{2}$ $10_{3}$ $12_{3}$
-$4$ $0100_{2}$ $0110_{2}$ $11_{3}$ $11_{3}$
-$5$ $0101_{2}$ $0111_{2}$ $12_{3}$ $10_{3}$
-$6$ $0110_{2}$ $0101_{2}$ $20_{3}$ $20_{3}$
-$7$ $0111_{2}$ $0100_{2}$ $21_{3}$ $21_{3}$
-$8$ $1000_{2}$ $1000_{2}$ $22_{3}$ $22_{3}$
-Table 1. Gray code order for bases $b=2$ and $b=3$. In Gray code order only one digit is incremented or decremented by $1$ (modulo $b$) when $i$ is incremented by $1$.
-The following code generates a base $2$ digital net with $R$ independent LMS, $R$ independent digital shifts, and $\alpha=2$ interlacing.
+0 $00_{2} 0000$$0_{2} 00$$_{3} 00$$0_{3}$
+11 $001_{2} 00$$001_{2}$ $01_{3}$ $01_{3}$
+22 $010_{2} 00$$011_{2}$ $02_{3}$ $02_{3}$
+33 $011_{2} 00$$010_{2}$ $10_{3}$ $12_{3}$
+44 $00_{2} 0110$$0_{2} 11$$_{3} 11$$1_{3}$
+55 $0101_{2}$ $0111_{2}$ $12_{3}$ $10_{3}$
+66 $10_{2} 0101$$01_{2}$ $20_{3}$ $20_{3}$
+77 $0111_{2}$ $0100_{2}$ $21_{3}$ $21_{3}$
+88 $1000_{2}$ $1000_{2}$ $2_{3} 22$$2_{3}$
+Table 1: Gray code order for bases $b=2$ and $b=3$. In Gray code order only one digit is incremented or decremented by 11 (modulo $b$) when $i$ is incremented by $1$.
+The following code generates $\alpha=2$ higher-order digital nets in base $2$ with $R$ independent LMS and digital shift combinations.
+Listing 3: Generate randomized higher-order digital nets.
 dnb2 = qp.DigitalNetB2(
 dimension = 52,
-randomize = "LMS_DS", # Matousek’s LMS with a digital shift
-# options are [LMS_DS, NUS, DS, LMS, False]
-t_lms = 64, # number of LMS bits i.e. number of rows in S_j
+randomize = "LMS DS", # Matousek’s LMS then a digital shift
+# other options ["NUS", "DS", "LMS", None]
+t = 64, # number of LMS bits, i.e., number of rows in S_j
 alpha = 2, # interlacing factor for higher order digital nets
-replications = 15, # R
-graycode = False, # use natural order instead of Gray code order
+replications = 16, # R
+order = "radical inverse", # also supports "Gray code"
 seed = None, # pass integer seed for reproducibility
-generating_matrices = "LDData:mps.sobol_Cs.txt")
-x = dnb2(2**16) # a numpy.ndarray with shape 15 x 65536 x 52
-Here we have used a set of Sobol’ generating matrices from Joe and Kuo131313the “new-joe-kuo-6.21201” direction numbers from https://web.maths.unsw.edu.au/~fkuo/sobol/index.html (Joe and Kuo, 2008) which are stored in a standardized format in the LDData repository.
+generating_matrices = "joe_kuo.6.21201.txt")
+x = dnb2(2**16) # a numpy.ndarray with shape (16, 65536, 52)
+Here we have used a set of Sobol’ generating matrices from Joe and Kuo161616“new-joe-kuo-6.21201” direction numbers from https://web.maths.unsw.edu.au/~fkuo/sobol/index.html [Joe and Kuo, 2008] which are stored in a standardized format in the LDData repository.
 Other generating matrices from LDData may be used by passing in a file name from https://github.com/QMCSoftware/LDData/blob/main/dnet/ or by passing an explicit array.
-4.3. Halton sequences
-The digital sequences described in the previous section used a fix prime base $b$.
+5.3 Halton sequences
+The digital sequences described in the previous section used a fixed prime base $b$.
 One may allow each dimension $j\in\{1,\dots,d\}$ to have its own prime base $b_{j}$.
@@
 However, digital interlacing is not generally applicable when the bases differ.
-Halton with random starting points has also been explored in (Wang and Hickernell, 2000), although we do not consider this here.
-The following code generates a Halton point set with $R$ independent LMS and $R$ independent digital permutations.
+Halton with random starting points has also been explored in [Wang and Hickernell, 2000], although we do not consider this here.
+The following code generates a Halton point set with $R$ independent LMS and digital permutation combinations.
+Specifically, we generate $R$ independent LMS-Halton point sets and then apply an independent permutation scramble to each.
+Listing 4: Generate randomized Halton points.
 halton = qp.Halton(
 dimension = 52,
-randomize = "LMS_PERM", # Matousek’s LMS with a digital permutation
-# options are [LMS_PERM, LMS_DS, LMS, PERM, DS, NUS, QRNG, FALSE]
-t_lms = 64, # number of LMS digits i.e. number of rows in S_j
-replications = 15, # R
+randomize = "LMS PERM", # Matousek’s LMS then a digital perm
+# or ["LMS DS", "LMS", "PERM", "DS", "NUS", "QRNG", None]
+t = 64, # number of LMS digits, i.e., number of rows in S_j
+replications = 16, # R
 seed = None) # pass integer seed for reproducibility
-x = halton(2**10) # a numpy.ndarray with shape 15 x 1024 x 52
-The QRNG randomization follows the QRNG software package (Hofert and Lemieux, 2023) in generating a generalized Halton point set (Faure and Lemieux, 2009) using a deterministic set of permutation scrambles and random digital shifts.
-5. Fast transforms and kernel computations
-Recall from the introduction that we would like to compute the matrix-vector product ${\mathsf{K}}{\boldsymbol{y}}$ and solve the linear system ${\mathsf{K}}^{-1}{\boldsymbol{y}}$ where ${\boldsymbol{y}}$ is some known length $n$ vector and ${\mathsf{K}}=\{K({\boldsymbol{x}}_{i},{\boldsymbol{x}}_{k})\}_{i,k=0}^{n-1}$ is a $n\times n$ symmetric positive definite (SPD) Gram matrix based on an SPD kernel $K$ and point set $\{{\boldsymbol{x}}_{i}\}_{i=0}^{n-1}$.
-We may reduce the standard ${\mathcal{O}}(n^{2})$ and ${\mathcal{O}}(n^{3})$ costs of solving these respective problems to both be ${\mathcal{O}}(n\log n)$ by inducing structure into ${\mathsf{K}}$ using special point set-kernel pairings.
-Specifically, pairing lattice points with a shift invariance kernel yields a circulant Gram matrix ${\mathsf{K}}$ for which computations can be done quickly using the bit-reversed fast Fourier transform (FFT) and bit-reversed inverse FFT (IFFT).
-Similarly, pairing a digital net with digitally-shift-invariant kernel yields a nested block Toeplitz matrix for which computations can be done quickly using the fast Walsh Hadamard Transform (FWHT).
-Shift-invariant and digitally-shift-invariant kernels are shown in fig.
-3.
-The reduced costs achieved by inducing these structures are summarized in table 2.
-Figure 3.
-Shift-invariant kernels (top row) and digitally-shift-invariant kernels (bottom row).
-Here $\alpha$ denotes the smoothness of the kernel.
-The $\alpha=3$ and $\alpha=4$ digitally-shift-invariant kernels are nearly indistinguishable.
-Shift-invariant kernels are described in section 5.1 while digitally-shift-invariant kernels are described in section 5.2.
-[…]
+x = halton(2**10) # a numpy.ndarray with shape (16, 1024, 52)
+The "QRNG" randomization follows the QRNG software package [Hofert and Lemieux, 2023] in generating a generalized Halton point set [Faure and Lemieux, 2009] using a deterministic set of permutation scrambles and random digital shifts.
+6 Fast transforms and kernel computations
+Recall from the motivation in Section 3.2 that we would like to compute the matrix-vector product ${\mathsf{K}}{\boldsymbol{y}}$ and solve the linear system ${\mathsf{K}}^{-1}{\boldsymbol{y}}$ where $\boldsymbol{y}$} is some known length $n$ vector and ${\mathsf{K}}=\{K({\boldsymbol{x}}_{i},{\boldsymbol{x}}_{i^{\prime}})\}_{i,i^{\prime}=0}^{n-1}$ is an $n\times n$ symmetric positive definite (SPD) Gram matrix based on a SPD kernel $K$ and point set $\{{\boldsymbol{x}}_{i}\}_{i=0}^{n-1}$.
+Table 2 compares our fast kernel methods against standard techniques in terms of both storage and computational costs.
+Figure 3 gives schematics of the FFTBR and FWHT algorithms which enable these fast kernel methods.
+The following subsections detail these special kernel-point set pairings and how the structure induce in $\mathsf{K}$} can be exploited for fast kernel computations.
 $\{{\boldsymbol{x}}_{i}\}_{i=0}^{n-1}$ $K$ structure $\mathsf{K}$} storage $\mathsf{K}$} decomposition ${\mathsf{K}}{\boldsymbol{y}}$ cost ${\mathsf{K}}^{-1}{\boldsymbol{y}}$ cost methods
-general SPD ${\mathcal{O}}(n^{2})$ ${\mathcal{O}}(n^{3})$ ${\mathcal{O}}(n^{2})$ ${\mathcal{O}}(n^{2})$ standard
-lattice SPD SI ${\mathcal{O}}(n)$ ${\mathcal{O}}(n\log n)$ ${\mathcal{O}}(n\log n)$ ${\mathcal{O}}(n\log n)$ FFT-based
-digital net SPD DSI ${\mathcal{O}}(n\log n)$ ${\mathcal{O}}(n)$ ${\mathcal{O}}(n\log n)$ ${\mathcal{O}}(n\log n)$ FWHT-based
-Table 2.
-Comparison of storage and cost requirements for kernel methods.
+any general SPD ${\mathcal{O}}(n^{2})$ ${\mathcal{O}}(n^{3})$ ${\mathcal{O}}(n^{2})$ ${\mathcal{O}}(n^{2})$ Cholesky
+rank-1 lattice SPD SI ${\mathcal{O}}(n)$ ${\mathcal{O}}(n\log n)$ ${\mathcal{O}}(n\log n)$ ${\mathcal{O}}(n\log n)$ FFTBR/IFFTBR
+digital net SPD DSI ${\mathcal{O}}(n)$ ${\mathcal{O}}(n\log n)$ ${\mathcal{O}}(n\log n)$ ${\mathcal{O}}(n\log n)$ FWHT
+Table 2: Comparison of storage and cost requirements for kernel methods.
 Decomposition of the symmetric positive definite (SPD) Gram matrix $\mathsf{K}$} is the cost of computing the eigendecomposition or Cholesky factorization.
@@
 Both storage and kernel computation costs are greatly reduced by pairing certain low-discrepancy point sets with special shift-invariant (SI) or digitally-shift-invariant (DSI) kernels.
-The following code can be used to construct a Gram matrix with a lattice point set and shift-invariant kernel.
-The Gram matrix is then used perform matrix-vector multiplication and solve a linear system.
-Analogous code using qp.DigitalNetB2, qp.KernelDigShiftInvar, and qp.FastGramMatrixDigitalNetB2 is also supported.
-d = 5 # dimension
-n = 2**20 # 2^20 points
-dd_obj = qp.Lattice(dimension=d, seed=7)
-# qp.Lattice defaults to natural ordering with a single random shift
-kernel_obj = qp.KernelShiftInvar( # matching kernel
-dimension = d,
-lengthscales = 0.1*np.ones(d,dtype=float), # product weights
-alpha = 4*np.ones(d,dtype=int)) # smoothness parameters
-K = qp.FastGramMatrixLattice( # n x n Gram matrix
-kernel_obj = kernel_obj,
-dd_obj = dd_obj,
-n1 = n,
-n2 = n)
-y = np.random.rand(n) # get a random length n vector
-a = K@y # compute the matrix-vector product a=Ky at O(n log n) cost
-b = K.solve(y) # solve linear system Kb=y for b at O(n log n) cost
-Schematics of the bit-reversed FFT and FWHT underlying these fast computations are shown in fig.
-4.
-The one-dimensional bit-reversed FFT, one-dimensional bit-reversed IFFT, and one-dimensional FWHT are implemented into the QMCPy functions qp.fftbr, qp.ifftbr, and qp.fwht respectively.
-These fast transforms act along the last dimension which is required to to have length $2^{m}$ for some $m\geq 0$.
-The following subsections detail the matching point set-kernel pairings and provide methods to quickly update the Gram matrix product or linear system solution after doubling the point size.
+These fast algorithms rely on the fast Fourier transform in bit-reversed order (FFTBR), its inverse (IFFTBR), and the fast Walsh–Hadamard transform (FWHT).
 […]
 […]
-Figure 4.
-Bit-reversed FFT and FWHT schemes.
+Figure 3: Bit-reversed FFT and FWHT schemes.
 The bit-reversed FFT is performed via a decimation in time algorithm without the initial reordering of inputs.
 The bit-reversed IFFT may be performed by propagating $\{{\widetilde{y}}_{i}\}_{i=0}^{2^{m}-1}$ to $\{y_{i}\}_{i=0}^{2^{m}-1}$ in the other direction through the bit-reversed FFT algorithm.
-[…]
-5.1. Shift invariant kernels, lattices, and the bit-reversed FFT/IFFT
-A kernel is said to be shift-invariant when $K({\boldsymbol{u}},{\boldsymbol{v}})={\widetilde{K}}(({\boldsymbol{u}}-{\boldsymbol{v}})\mod 1)$ for some ${\widetilde{K}}$ i.e.
-the kernel is only a function of the component-wise difference between inputs modulo $1$.
-One set of shift-invariant kernels take the form
-(4) $K({\boldsymbol{u}},{\boldsymbol{v}})=\sum_{{\mathfrak{u}}\subseteq{\mathcal{D}}}\gamma_{\mathfrak{u}}\prod_{j\in{\mathfrak{u}}}\eta_{\alpha_{j}}((u_{j}-v_{j})\mod 1)$
+6.1 SI kernels, rank-1 lattices, and the FFTBR/IFFTBR
+A kernel is said to be shift-invariant (SI) when
+$K({\boldsymbol{x}},{\boldsymbol{x}}^{\prime})={\widetilde{K}}(({\boldsymbol{x}}-{\boldsymbol{x}}^{\prime})\mod 1)$
+for some $\widetilde{K}$}, i.e., the kernel is only a function of the component-wise difference between inputs modulo $1$. One set of SI kernels take the form
+$K({\boldsymbol{x}},{\boldsymbol{x}}^{\prime})=\gamma\sum_{{\mathfrak{u}}\subseteq\{1,\dots,d\}}\eta_{\mathfrak{u}}\prod_{j\in{\mathfrak{u}}}K_{\alpha_{j}}(x_{j},x_{j}^{\prime})$ (5)
 where
-$\eta_{\alpha}(\delta)=\frac{(2\pi)^{2\alpha}}{(-1)^{\alpha+1}(2\alpha)!}B_{2\alpha}(\delta)$
-with ${\boldsymbol{\alpha}}\in{\mathbb{N}}^{d}$, ${\mathcal{D}}=\{1,\dots,d\}$, weights $\{\gamma_{\mathfrak{u}}\}_{{\mathfrak{u}}\subseteq{\mathcal{D}}}$, and $B_{\ell}$ denoting the Bernoulli polynomial of degree $\ell$.
-The corresponding reproducing kernel Hilbert space (RKHS) $H$ is a weighted periodic unanchored Sobolev space of smoothness ${\boldsymbol{\alpha}}\in{\mathbb{N}}^{d}$ with norm
-$\lVert f\rVert_{H}^{2}:=\sum_{{\mathfrak{u}}\subseteq{\mathcal{D}}}\frac{1}{(2\pi)^{2\lvert{\mathfrak{u}}\rvert}\gamma_{\mathfrak{u}}}\int_{[0,1]^{\lvert{\mathfrak{u}}\rvert}}\left\lvert\int_{[0,1]^{s-\lvert{\mathfrak{u}}\rvert}}\left(\prod_{j\in{\mathfrak{u}}}\frac{\partial^{\alpha_{j}}}{\partial y_{j}^{\alpha_{j}}}\right)f({\boldsymbol{y}})\mathrm{d}{\boldsymbol{y}}_{-{\mathfrak{u}}}\right\rvert^{2}\mathrm{d}{\boldsymbol{y}}_{\mathfrak{u}},$
-where $f:[0,1]^{d}\to{\mathbb{R}}$, ${\boldsymbol{y}}_{\mathfrak{u}}=({\boldsymbol{y}}_{j})_{j\in{\mathfrak{u}}}$, ${\boldsymbol{y}}_{-{\mathfrak{u}}}:=(y_{j})_{j\in{\mathcal{D}}\setminus{\mathfrak{u}}}$, and $\lvert{\mathfrak{u}}\rvert$ is the cardinality of ${\mathfrak{u}}$.
-The space $H$ is a special case of the weighted Korobov space which has real smoothness parameter $\alpha$ characterizing the rate of decay of Fourier coefficients, see e.g.
-(Kaarnioja et al., 2022, 2023; Cools et al., 2021, 2020; Sloan and Woźniakowski, 2001; Kuo et al., 2004).
+$K_{\alpha}(x,x^{\prime})=\frac{(2\pi)^{2\alpha}}{(-1)^{\alpha+1}(2\alpha)!}B_{2\alpha}((x-x^{\prime})\mod 1)$
+with smoothness parameters ${\boldsymbol{\alpha}}\in{\mathbb{N}}^{d}$, global scale $\gamma>0$, weights (lengthscales) $\{\eta_{\mathfrak{u}}\}_{{\mathfrak{u}}\subseteq\{1,\dots,d\}}$, and $B_{\ell}$ denoting the Bernoulli polynomial of degree $\ell$.
+The corresponding RKHS $H$ is a weighted periodic unanchored Sobolev space of smoothness ${\boldsymbol{\alpha}}\in{\mathbb{N}}^{d}$ with norm
+$\lVert f\rVert_{H}^{2}:=\sum_{{\mathfrak{u}}\subseteq\{1,\dots,d\}}\frac{1}{(2\pi)^{2\lvert{\mathfrak{u}}\rvert}\eta_{\mathfrak{u}}}\int_{[0,1]^{\lvert{\mathfrak{u}}\rvert}}\left\lvert\int_{[0,1]^{s-\lvert{\mathfrak{u}}\rvert}}\left(\prod_{j\in{\mathfrak{u}}}\frac{\partial^{\alpha_{j}}}{\partial y_{j}^{\alpha_{j}}}\right)f({\boldsymbol{y}})\mathrm{d}{\boldsymbol{y}}_{-{\mathfrak{u}}}\right\rvert^{2}\mathrm{d}{\boldsymbol{y}}_{\mathfrak{u}},$
+where $f:[0,1]^{d}\to{\mathbb{R}}$, ${\boldsymbol{y}}_{\mathfrak{u}}=({\boldsymbol{y}}_{j})_{j\in{\mathfrak{u}}}$, ${\boldsymbol{y}}_{-{\mathfrak{u}}}:=(y_{j})_{j\in\{1,\dots,d\}\setminus{\mathfrak{u}}}$, and $\lvert{\mathfrak{u}}\rvert$ is the cardinality of $\mathfrak{u}$}.
+The space $H$ is a special case of the weighted Korobov space which has real smoothness parameter $\alpha$ characterizing the rate of decay of Fourier coefficients [Kaarnioja et al., 2022, 2023, Cools et al., 2021, 2020, Sloan and Woźniakowski, 2001, Kuo et al., 2004].
 The kernel (5) is the sum over $2^{d}$ terms and thus becomes impractical to compute for large $d$.
-A simplified space assumes $\gamma_{\mathfrak{u}}$ are product weights which take the form $\gamma_{\mathfrak{u}}=\prod_{j\in{\mathfrak{u}}}\gamma_{j}$ for some $\{\gamma_{j}\}_{j=1}^{d}$.
+A simplified space assumes $\eta_{\mathfrak{u}}$ are product weights which take the form $\eta_{\mathfrak{u}}=\prod_{j\in{\mathfrak{u}}}\eta_{j}$ for some $\{\eta_{j}\}_{j=1}^{d}$.
 The kernel with product weights takes the simplified form
-(5) $K({\boldsymbol{u}},{\boldsymbol{v}})=\prod_{j=1}^{d}\left(1+\gamma_{j}\eta_{\alpha_{j}}((u_{j}-v_{j})\mod 1)\right).$
-Figure 3 shows one dimensional shift-invariant kernels with unit weight and varying smoothness.
-See (Kaarnioja et al., 2023) for a review of lattice-based kernel interpolation in such weighted spaces along with the development of serendipitous weights, which are a special form of high-performance product weights.
+$K({\boldsymbol{x}},{\boldsymbol{x}}^{\prime})=\gamma\prod_{j=1}^{d}\left(1+\eta_{j}K_{\alpha_{j}}(x_{j},x_{j}^{\prime})\right).$ (6)
+See [Kaarnioja et al., 2023] for a review of lattice-based kernel interpolation in such weighted spaces along with the development of serendipitous weights, which are a special form of high-performance product weights.
 Suppose $n=2^{m}$ for some $m\in{\mathbb{N}}_{0}$ and let $\{{\boldsymbol{x}}_{i}\}_{i=0}^{b^{m}-1}$ be a shifted lattice with shift ${\boldsymbol{\Delta}}\in[0,1)^{d}$ generated in linear order.
@@
 Let $R_{m}(i)$ flip the first $m$ bits of $0\leq i<2^{m}$ in base $b=2$ so that if $i=\sum_{t=0}^{m-1}{\mathsf{i}}_{t}2^{t}$ then $R(i)=\sum_{t=0}^{m-1}{\mathsf{i}}_{m-t-1}2^{t}$.
-The first step in the FFT with decimation-in-time is to reorder the inputs $\{y_{i}\}_{i=0}^{2^{m}-1}$ into bit reversed order $\{y_{R(i)}\}_{i=0}^{2^{m}-1}$.
-The last step in computing the IFFT is to reorder the outputs $\{y_{R(i)}\}_{i=0}^{2^{m}-1}$ into bit reversed order $\{y_{i}\}_{i=0}^{2^{m}-1}$.
-Therefore, one should generate lattice points in natural order and skip the first step of the FFT and last step of the inverse FFT.
-The Fourier matrix is $\overline{{\mathsf{F}}^{(m)}}=\{W_{m}^{ij}\}_{i,j=0}^{2^{m}-1}$ where $W_{m}=\exp(-2\pi\sqrt{-1}/2^{m})$.
-Let $\overline{{\mathsf{T}}^{(m)}}=\{W_{m}^{iR_{m}(j)}\}_{i,j=0}^{2^{m}-1}$ so that
-$\overline{{\mathsf{F}}^{(m)}\{y_{i}\}_{j=0}^{2^{m}-1}}=\overline{{\mathsf{T}}^{(m)}\{y_{R(j)}\}_{j=0}^{2^{m}-1}}.$
-For $\{{\boldsymbol{x}}_{i}\}_{i=0}^{2^{m}-1}$ a lattice in natural order, we have
-${\mathsf{K}}=\frac{1}{n}{\mathsf{T}}^{(m)}{\mathsf{\Lambda}}^{(m)}\overline{{\mathsf{T}}^{(m)}}.$
-Notice that $\overline{{\mathsf{T}}^{(m+1)}}=\left\{W_{m+1}^{iR_{m+1}(j)}\right\}_{i,j=0}^{2^{m+1}-1}$.
+The first step in the FFT with decimation-in-time is to reorder the inputs $\{y_{i}\}_{i=0}^{2^{m}-1}$ into bit-reversed order $\{y_{R(i)}\}_{i=0}^{2^{m}-1}$.
+The last step in computing the IFFT is to reorder the outputs $\{y_{R(i)}\}_{i=0}^{2^{m}-1}$ into bit-reversed order $\{y_{i}\}_{i=0}^{2^{m}-1}$.
+Therefore, one should generate lattice points in radical inverse order and skip the first step of the FFT and last step of the inverse FFT.
+We call such algorithms the FFTBR and IFFTBR respectively.
+The Fourier matrix is $\overline{{\mathsf{F}}_{m}}=\{W_{m}^{ij}\}_{i,j=0}^{2^{m}-1}$ where $W_{m}=\exp(-2\pi\sqrt{-1}/2^{m})$.
+Let $\overline{{\mathsf{T}}_{m}}=\{W_{m}^{iR_{m}(j)}\}_{i,j=0}^{2^{m}-1}$ so that
+$\overline{{\mathsf{F}}_{m}\{y_{i}\}_{j=0}^{2^{m}-1}}=\overline{{\mathsf{T}}_{m}\{y_{R(j)}\}_{j=0}^{2^{m}-1}}.$
+For $\{{\boldsymbol{x}}_{i}\}_{i=0}^{2^{m}-1}$ a lattice in radical inverse order, we have
+${\mathsf{K}}=\frac{1}{n}{\mathsf{T}}_{m}{\mathsf{\Lambda}}_{m}\overline{{\mathsf{T}}_{m}}.$
+Notice that $\overline{{\mathsf{T}}_{m+1}}=\left\{W_{m+1}^{iR_{m+1}(j)}\right\}_{i,j=0}^{2^{m+1}-1}$.
 For $0\leq j<2^{m}$ we have $R_{m+1}(j)=2R_{m}(j)$ so $W^{iR_{m+1}(j)}_{m+1}=W^{iR_{m}(j)}_{m}$.
@@
 Moreover, for $0\leq i<2^{m}$ we have $W_{m+1}^{2^{m}+i}=-W_{m+1}^{i}$.
-Define ${\widetilde{\boldsymbol{w}}}^{(m)}=\{W_{m+1}^{i}\}_{i=0}^{2^{m}-1}$.
+Define ${\widetilde{\boldsymbol{w}}}_{m}=\{W_{m+1}^{i}\}_{i=0}^{2^{m}-1}$.
 Then
-$\overline{{\mathsf{T}}^{(m+1)}}=\begin{pmatrix}\overline{{\mathsf{T}}^{(m)}}&\mathrm{diag}({\widetilde{\boldsymbol{w}}}^{(m)})\overline{{\mathsf{T}}^{(m)}}\\ \overline{{\mathsf{T}}^{(m)}}&-\mathrm{diag}({\widetilde{\boldsymbol{w}}}^{(m)})\overline{{\mathsf{T}}^{(m)}}\end{pmatrix}$
-and ${\widetilde{\boldsymbol{w}}}^{(m+1)}$ is the $\alpha=2$ interlacing of ${\widetilde{\boldsymbol{w}}}^{(m)}$ and $W_{m+1}{\widetilde{\boldsymbol{w}}}^{(m)}$.
-5.2. Digital nets, digitally-shift-invariant kernels, and the FWHT
+$\overline{{\mathsf{T}}_{m+1}}=\begin{pmatrix}\overline{{\mathsf{T}}_{m}}&\mathrm{diag}({\widetilde{\boldsymbol{w}}}_{m})\overline{{\mathsf{T}}_{m}}\\ \overline{{\mathsf{T}}_{m}}&-\mathrm{diag}({\widetilde{\boldsymbol{w}}}_{m})\overline{{\mathsf{T}}_{m}}\end{pmatrix}$
+and ${\widetilde{\boldsymbol{w}}}_{m+1}$ is the $\alpha=2$ interlacing of ${\widetilde{\boldsymbol{w}}}_{m}$ and $W_{m+1}{\widetilde{\boldsymbol{w}}}_{m}$.
+6.2 DSI kernels, digital nets, and the FWHT
 For $x,y\in[0,1)$ with $x=\sum_{t=1}^{m}{\mathsf{x}}_{t-1}b^{-t}$ and $y=\sum_{t=1}^{m}{\mathsf{y}}_{t-1}b^{-t}$ let
@@
 For vectors ${\boldsymbol{x}},{\boldsymbol{y}}\in[0,1)^{d}$, digital operations ${\boldsymbol{x}}\oplus{\boldsymbol{y}}$ and ${\boldsymbol{x}}\ominus{\boldsymbol{y}}$ act component-wise.
-A kernel is said to be digitally-shift-invariant when
-$K({\boldsymbol{u}},{\boldsymbol{v}})={\widetilde{K}}({\boldsymbol{u}}\ominus{\boldsymbol{v}})$
+A kernel is said to be digitally-shift-invariant (DSI) when
+$K({\boldsymbol{x}},{\boldsymbol{x}}^{\prime})={\widetilde{K}}({\boldsymbol{x}}\ominus{\boldsymbol{x}}^{\prime})$
 for some $\widetilde{K}$}.
-Suppose $n=b^{m}$ for some $m\in{\mathbb{N}}_{0}$ and let $\{{\boldsymbol{x}}_{i}\}_{i=0}^{b^{m}-1}$ be a digitally shifted digital net in natural order (possibly of higher order and/or with LMS).
-Following (Jagadeeswaran and Hickernell, 2022, Theorem 5.3.1,Theorem 5.3.2), for $b=2$, the Gram matrix ${\mathsf{K}}$ is nested block Toeplitz which implies the eigendecomposition
-${\mathsf{K}}=\frac{1}{n}{\mathsf{H}}^{(m)}{\mathsf{\Lambda}}^{(m)}{\mathsf{H}}^{(m)}$
-where ${\mathsf{H}}^{(m)}$ is the $2^{m}\times 2^{m}$ Hadamard matrix defined by ${\mathsf{H}}^{(0)}=(1)$ and the relationship
-${\mathsf{H}}^{(m+1)}=\begin{pmatrix}{\mathsf{H}}^{(m)}&{\mathsf{H}}^{(m)}\\ {\mathsf{H}}^{(m)}&-{\mathsf{H}}^{(m)}\end{pmatrix}.$
-Multiplying by ${\mathsf{H}}^{(m)}$ can be done at at ${\mathcal{O}}(n\log n)$ cost using the FWHT. Note that ${\mathsf{K}}$ is not necessarily nested block Toeplitz under NUS.
-We now describe one dimensional digitally-shift-invariant kernels $K(u,v)={\widetilde{K}}(u\ominus v)$ of higher-order smoothness $\alpha$.
-Weighted sums over products these one dimensional kernels may be used to construct higher-dimensional kernels e.g.
-those with product weights
-$K({\boldsymbol{u}},{\boldsymbol{v}})=\prod_{j=1}^{d}\left(1+\gamma_{j}\left(K(u_{j},v_{j})-1\right)\right).$
-To begin, let us write $k\in{\mathbb{N}}_{0}$ as $k=\sum_{\ell=1}^{\#k}{\mathsf{k}}_{a_{\ell}}b^{a_{\ell}}$ where $a_{1}>\cdots>a_{\#k}\geq 0$ are the #$\#k$ indices of non-zero digits ${\mathsf{k}}_{a_{\ell}}\in\{1,\dots,b-1\}$ in the base $b$ expansion of $k$.
+Suppose $n=b^{m}$ for some $m\in{\mathbb{N}}_{0}$ and let $\{{\boldsymbol{x}}_{i}\}_{i=0}^{b^{m}-1}$ be a digitally-shifted digital net in radical inverse order (possibly of higher-order and/or with LMS).
+Following [Rathinavel and Hickernell, 2022, Theorem 5.3.1,Theorem 5.3.2], for $b=2$, the Gram matrix $\mathsf{K}$} is RSBT (recursive symmetric block Toeplitz) which implies the eigendecomposition
+${\mathsf{K}}=\frac{1}{n}{\mathsf{H}}_{m}{\mathsf{\Lambda}}_{m}{\mathsf{H}}_{m}$
+where ${\mathsf{H}}_{m}$ is the $2^{m}\times 2^{m}$ Hadamard matrix defined by ${\mathsf{H}}^{(0)}=(1)$ and the relationship
+${\mathsf{H}}_{m+1}=\begin{pmatrix}{\mathsf{H}}_{m}&{\mathsf{H}}_{m}\\ {\mathsf{H}}_{m}&-{\mathsf{H}}_{m}\end{pmatrix}$.
+Multiplying by ${\mathsf{H}}_{m}$ can be done at ${\mathcal{O}}(n\log n)$ cost using the FWHT. Note that $\mathsf{K}$} is not generally RSBT when $\{{\boldsymbol{x}}_{i}\}_{i=0}^{b^{m}-1}$ has a NUS.
+We now describe one dimensional DSI kernels $K_{\alpha}(x,x^{\prime})$ of higher-order smoothness $\alpha\geq 2$.
+Weighted sums over products these one dimensional kernels may be used to construct higher-dimensional kernels, e.g., those with product weights
+$K({\boldsymbol{x}},{\boldsymbol{x}}^{\prime})=\gamma\prod_{j=1}^{d}\left(1+\eta_{j}K_{\alpha_{j}}(x_{j},x_{j}^{\prime})\right).$
+To begin, let us write $k\in{\mathbb{N}}_{0}$ as $k=\sum_{\ell=1}^{\#k}{\mathsf{k}}_{a_{\ell}}b^{a_{\ell}}$ where $a_{1}>\cdots>a_{\#k}\geq 0$ are the #k\#k indices of non-zero digits ${\mathsf{k}}_{a_{\ell}}\in\{1,\dots,b-1\}$ in the base $b$ expansion of $k$.
 Then the $k^{\mathrm{th}}$ Walsh coefficient is
@@
 $\mu_{\alpha}(k)=\sum_{\ell=1}^{\min\{\alpha,\#k\}}(a_{\ell}+1).$
-For $\alpha>1$, let the digitally-shift-invariant kernel
-(6) $K(x,z)=\sum_{k\in{\mathbb{N}}_{0}}\frac{\mathrm{wal}_{k}(x\ominus z)}{b^{\mu_{\alpha}(k)}}={\widetilde{K}}_{\alpha}(x\ominus z)$
-have corresponding RKHS ${\widetilde{H}}_{\alpha}$.
-The following theorem shows ${\widetilde{H}}_{\alpha}$ contains smooth functions.
-The case when $\alpha=1$ is treated separately in (Dick and Pillichshammer, 2005).
+For $\alpha\geq 2$, let the DSI kernel
+$K_{\alpha}(x,x^{\prime})=\sum_{k\in{\mathbb{N}}}\frac{\mathrm{wal}_{k}(x\ominus x^{\prime})}{b^{\mu_{\alpha}(k)}}$ (7)
+have corresponding RKHS $H_{\alpha}$. The following theorem shows $H_{\alpha}$ contains smooth functions. The case when $\alpha=1$ is treated separately in [Dick and Pillichshammer, 2005].
 Theorem 6.1.
-${\widetilde{H}}_{\alpha}\supset H_{\alpha}$ where $H_{\alpha}$ is an RKHS with inner product
-$\langle f,g\rangle_{H_{\alpha}}=\sum_{l=1}^{\alpha-1}\int_{0}^{1}f^{(l)}(x)\mathrm{d}x\int_{0}^{1}g^{(l)}(x)\mathrm{d}x+\int_{0}^{1}f^{(\alpha)}(x)g^{(\alpha)}(x)\mathrm{d}x.$
+Let $H_{\alpha}^{\prime}$ denote the RKHS with inner product
+$\langle f,g\rangle_{H_{\alpha}^{\prime}}=\sum_{l=1}^{\alpha-1}\int_{0}^{1}f^{(l)}(x)\mathrm{d}x\int_{0}^{1}g^{(l)}(x)\mathrm{d}x+\int_{0}^{1}f^{(\alpha)}(x)g^{(\alpha)}(x)\mathrm{d}x$
+and $f\in H_{\alpha}^{\prime}$ satisfying $\int_{0}^{1}f(x)\mathrm{d}x=0$. Then $H_{\alpha}^{\prime}\subset H_{\alpha}$ where $H_{\alpha}$ is the RKHS with kernel defined in Equation 7.
 Proof.
-See appendix A
-The following theorem gives explicit forms of a few higher-order digitally-shift-invariant kernels in base $b=2$ which are plotted in fig.
-3.
-These kernels are also useful in the computation of the worst-case error of QMC with higher order polynomial-lattice rules as in (Baldeaux et al., 2012).
-In fact, their paper details how to compute ${\widetilde{K}}_{\alpha}$ at ${\mathcal{O}}(\alpha\#x)$ cost where #$\#x$ is the number of non-zero digits in the base $b$ expansion of $x$
+See Appendix B
+The following theorem gives explicit forms of a few higher-order DSI kernels in base $b=2$.
+These kernels are also useful in the computation of the worst-case error of QMC with higher-order polynomial-lattice rules as in [Baldeaux et al., 2012].
+In fact, their paper details how to compute $K_{\alpha}$ in (7) at ${\mathcal{O}}(\alpha\#x)$ cost where #x\#x is the number of non-zero digits in the base $b$ expansion of $x$.
 Theorem 6.2.
-Fix the base $b=2$. Let $\beta(x)=-\lfloor\log_{2}(x)\rfloor$ and $t_{\nu}(x)=2^{-\nu\beta(x)}$ where for $x=0$ we set $\beta(x)=t_{\nu}(x)=0$. Then
-${\widetilde{K}}_{2}(x)$ $=-\beta(x)x+\frac{5}{2}\left[1-t_{1}(x)\right],$
-${\widetilde{K}}_{3}(x)$ $=\beta(x)x^{2}-5\left[1-t_{1}(x)\right]x+\frac{43}{18}\left[1-t_{2}(x)\right],$
-${\widetilde{K}}_{4}(x)$ $=-\frac{2}{3}\beta(x)x^{3}+5\left[1-t_{1}(x)\right]x^{2}-\frac{43}{9}\left[1-t_{2}(x)\right]x$
+Fix the base $b=2$. Let $\beta(x)=-\lfloor\log_{2}(x)\rfloor$ and for $\nu\in{\mathbb{N}}$ define $t_{\nu}(x)=2^{-\nu\beta(x)}$ where for $x=0$ we set $\beta(x)=t_{\nu}(x)=0$. Then (7) satisfies
+$K_{2}(x)$ $=-1+-\beta(x)x+\frac{5}{2}\left[1-t_{1}(x)\right],$
+$K_{3}(x)$ $=-1+\beta(x)x^{2}-5\left[1-t_{1}(x)\right]x+\frac{43}{18}\left[1-t_{2}(x)\right],$
+$K_{4}(x)$ $=-1+-\frac{2}{3}\beta(x)x^{3}+5\left[1-t_{1}(x)\right]x^{2}-\frac{43}{9}\left[1-t_{2}(x)\right]x$
 $+\frac{701}{294}\left[1-t_{3}(x)\right]+\beta(x)\left[\frac{1}{48}\sum_{a=0}^{\infty}\frac{\mathrm{wal}_{2^{a}}(x)}{2^{3a}}-\frac{1}{42}\right].$
 Proof.
-The $\alpha\in\{2,3\}$ forms are due to (Baldeaux et al., 2012). The $\alpha=4$ form is derived in appendix A. ∎
-5.3. Fast kernel computations and efficient updates
-Let us make the general assumptions satisfied by the formulations in the previous two subsections.
-  1. (1)
-    For $n=2^{m}$, the Gram matrix ${\mathsf{K}}^{(m)}\in{\mathbb{R}}^{n\times n}$ can be written as
-    ${\mathsf{K}}^{m}={\mathsf{V}}^{(m)}{\mathsf{\Lambda}}^{(m)}\overline{{\mathsf{V}}^{(m)}}$
-    where ${\mathsf{V}}^{(m)}\overline{{\mathsf{V}}^{(m)}}={\mathsf{I}}$ and the first column of ${\mathsf{V}}^{(m)}$, denoted ${\boldsymbol{v}}_{1}^{(m)}$, is the constant $1/\sqrt{n}$
-  2. (2)
-    $V^{(m)}\in{\mathbb{C}}^{n\times n}$ satisfies ${\mathsf{V}}^{(0)}=(1)$ and
-    ${\mathsf{V}}^{(m+1)}=\begin{pmatrix}{\mathsf{V}}^{(m)}&{\mathsf{V}}^{(m)}\\ {\mathsf{V}}^{(m)}\overline{\mathrm{diag}({\widetilde{\boldsymbol{w}}}^{(m)})}&-{\mathsf{V}}^{(m)}\overline{\mathrm{diag}({\widetilde{\boldsymbol{w}}}^{(m)})}\end{pmatrix}/\sqrt{2}.$
-  3. (3)
-    ${\mathsf{V}}^{(m)}{\boldsymbol{y}}$ and $\overline{{\mathsf{V}}^{(m)}}{\boldsymbol{y}}$ are each computable at cost ${\mathcal{O}}(2^{m}m)$.
-For the case of lattice points with shift-invariant kernels and resulting circulant Gram matrices, ${\mathsf{V}}^{(m)}={\mathsf{T}}^{(m)}/\sqrt{2^{m}}$ is the scaled and permuted Fourier matrix and ${\widetilde{\boldsymbol{w}}}^{(m)}=\{\exp(-\pi\sqrt{-1}/2^{m}i)\}_{i=0}^{2^{m}-1}$, so ${\mathsf{V}}^{(m)}{\boldsymbol{y}}$ and $\overline{{\mathsf{V}}^{(m)}}{\boldsymbol{y}}$ can be computed using a bit-reversed FFT and bit-reversed IFFT respectively.
-For the case of digital nets with digitally-shift-invariant kernels and resulting nested block Toeplitz Gram matrices, ${\mathsf{V}}^{(m)}={\mathsf{H}}^{(m)}/\sqrt{n}$ is the scaled Hadamard matrix and ${\widetilde{\boldsymbol{w}}}^{(m)}=\{1\}_{i=0}^{2^{m}-1}$, so ${\mathsf{V}}^{(m)}{\boldsymbol{y}}$ and $\overline{{\mathsf{V}}^{(m)}}{\boldsymbol{y}}$ can be computed using a FWHT and inverse FWHT respectively.
-For ${\mathsf{\Lambda}}=\mathrm{diag}({\boldsymbol{\lambda}})$, we have
-${\boldsymbol{\lambda}}=\sqrt{n}{\mathsf{\Lambda}}^{(m)}\overline{{\boldsymbol{v}}_{1}^{(m)}}=\sqrt{n}\overline{{\mathsf{V}}^{(m)}}\left({\mathsf{V}}^{(m)}{\mathsf{\Lambda}}^{(m)}\overline{{\boldsymbol{v}}_{1}^{(m)}}\right)=\sqrt{n}\overline{{\mathsf{V}}^{(m)}}{\boldsymbol{k}}_{1}$
-which can be computed at ${\mathcal{O}}(n\log n)$ cost. Moreover,
-${\mathsf{K}}{\boldsymbol{y}}={\mathsf{V}}^{(m)}({\widetilde{\boldsymbol{y}}}\odot{\boldsymbol{\lambda}})\qquad\text{and}\qquad{\mathsf{K}}^{-1}{\boldsymbol{y}}={\mathsf{V}}^{(m)}({\widetilde{\boldsymbol{y}}}./{\boldsymbol{\lambda}})$
-may each be evaluated at cost ${\mathcal{O}}(n\log n)$ where $\odot$ denotes the Hadamard (component-wise) product and ././. / denotes component-wise division.
-To update the solutions of the motivating problems, it is easily shown that
-${\mathsf{K}}^{(m+1)}{\boldsymbol{y}}^{(m+1)}$ $=\sqrt{2}\begin{pmatrix}{\mathsf{V}}^{(m)}\left({\widetilde{\boldsymbol{y}}}^{(1)}\odot{\boldsymbol{\lambda}}^{(1)}+{\widetilde{\boldsymbol{w}}}^{(m)}\odot{\widetilde{\boldsymbol{w}}}^{(m)}\odot{\widetilde{\boldsymbol{y}}}^{(2)}\odot{\boldsymbol{\lambda}}^{(2)}\right)\\ {\mathsf{V}}^{(m)}\left({\widetilde{\boldsymbol{y}}}^{(2)}\odot{\boldsymbol{\lambda}}^{(1)}+{\widetilde{\boldsymbol{y}}}^{(1)}\odot{\boldsymbol{\lambda}}^{(2)}\right)\end{pmatrix}$
-${\mathsf{K}}^{-(m+1)}{\boldsymbol{y}}^{(m+1)}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}{\mathsf{V}}^{(m)}\left(\left({\widetilde{\boldsymbol{y}}}^{(1)}\odot{\boldsymbol{\lambda}}^{(1)}-{\widetilde{\boldsymbol{w}}}^{(m)}\odot{\widetilde{\boldsymbol{w}}}^{(m)}\odot{\widetilde{\boldsymbol{y}}}^{(2)}\odot{\boldsymbol{\lambda}}^{(2)}\right)./{\widetilde{\boldsymbol{\gamma}}}\right)\\ {\mathsf{V}}^{(m)}\left(\left({\widetilde{\boldsymbol{y}}}^{(2)}\odot{\boldsymbol{\lambda}}^{(1)}-{\widetilde{\boldsymbol{y}}}^{(1)}\odot{\boldsymbol{\lambda}}^{(2)}\right)./{\widetilde{\boldsymbol{\gamma}}}\right)\end{pmatrix}$
-where ${\widetilde{\boldsymbol{\gamma}}}={\boldsymbol{\lambda}}^{(1)}\odot{\boldsymbol{\lambda}}^{(1)}-{\widetilde{\boldsymbol{w}}}^{(m)}\odot{\widetilde{\boldsymbol{w}}}^{(m)}\odot{\boldsymbol{\lambda}}^{(2)}\odot{\boldsymbol{\lambda}}^{(2)}$.
-6. Numerical experiments
-All numerical experiments were carried out on a 2019 MacBook Pro with an Intel Core i7 processor.
-Figure 5 compares the wall-clock time required to generate point sets and perform fast transforms.
-Since scaling in the number dimensions and number of randomizations is linear, point sets are only generated in $d=1$ with a single randomization.
-IID points, lattices with shifts, and digital nets with LMS and digital shifts (including higher-order versions) are the fastest sequences to generate.
+The $\alpha\in\{2,3\}$ forms are due to [Baldeaux et al., 2012]. The $\alpha=4$ form is derived in Appendix B. ∎
+6.3 Fast kernel computations and efficient eigenvalue updates
+The formulations in the previous two subsections have the following properties.
+  1. 1.
+    For $n=2^{m}$, the Gram matrix ${\mathsf{K}}_{m}\in{\mathbb{R}}^{n\times n}$ can be written as
+    ${\mathsf{K}}_{m}={\mathsf{V}}_{m}{\mathsf{\Lambda}}_{m}\overline{{\mathsf{V}}_{m}}$
+    where ${\mathsf{V}}_{m}\overline{{\mathsf{V}}_{m}}={\mathsf{I}}$ and the first column of ${\mathsf{V}}_{m}$, denoted ${\boldsymbol{v}}_{m1}$, is the constant $1/\sqrt{n}$.
+  2. 2.
+    ${\mathsf{V}}_{m}\in{\mathbb{C}}^{n\times n}$ satisfies ${\mathsf{V}}_{0}=1$ and
+    ${\mathsf{V}}_{m+1}=\begin{pmatrix}{\mathsf{V}}_{m}&{\mathsf{V}}_{m}\\ {\mathsf{V}}_{m}\overline{\mathrm{diag}({\widetilde{\boldsymbol{w}}}_{m})}&-{\mathsf{V}}_{m}\overline{\mathrm{diag}({\widetilde{\boldsymbol{w}}}_{m})}\end{pmatrix}/\sqrt{2}.$ (8)
+  3. 3.
+    ${\mathsf{V}}_{m}{\boldsymbol{y}}$ and $\overline{{\mathsf{V}}_{m}}{\boldsymbol{y}}$ are each computable at cost ${\mathcal{O}}(2^{m}m)={\mathcal{O}}(n\log n)$.
+For the case of lattice points with SI kernels, ${\mathsf{V}}_{m}={\mathsf{T}}_{m}/\sqrt{2^{m}}$ is the scaled and permuted Fourier matrix and ${\widetilde{\boldsymbol{w}}}_{m}=\{\exp(-\pi\sqrt{-1}/2^{m}i)\}_{i=0}^{2^{m}-1}$, so ${\mathsf{V}}_{m}{\boldsymbol{y}}$ and $\overline{{\mathsf{V}}_{m}}{\boldsymbol{y}}$ can be computed using the FFTBR and IFFTBR respectively.
+For the case of digital nets with DSI kernels, ${\mathsf{V}}_{m}={\mathsf{H}}_{m}/\sqrt{n}$ is the real symmetric Hadamard matrix and ${\widetilde{\boldsymbol{w}}}_{m}=\{1\}_{i=0}^{2^{m}-1}$, so both ${\mathsf{V}}_{m}{\boldsymbol{y}}$ and $\overline{{\mathsf{V}}_{m}}{\boldsymbol{y}}$ can be computed using a FWHT.
+With ${\mathsf{\Lambda}}_{m}=\mathrm{diag}({\boldsymbol{\lambda}}_{,})$, we have
+${\boldsymbol{\lambda}}_{m}=\sqrt{n}{\mathsf{\Lambda}}_{m}\overline{{\boldsymbol{v}}_{m1}}=\sqrt{n}\overline{{\mathsf{V}}_{m}}\left({\mathsf{V}}_{m}{\mathsf{\Lambda}}_{m}\overline{{\boldsymbol{v}}_{m1}}\right)=\sqrt{n}\overline{{\mathsf{V}}_{m}}{\boldsymbol{k}}_{m1}$
+which can be computed at ${\mathcal{O}}(n\log n)$ cost and only requires storing the first column of ${\mathsf{K}}^{(m)}$ which we denote by ${\boldsymbol{k}}_{m1}$. Moreover,
+${\mathsf{K}}{\boldsymbol{y}}={\mathsf{V}}_{m}({\widetilde{\boldsymbol{y}}}\odot{\boldsymbol{\lambda}}_{m})\qquad\text{and}\qquad{\mathsf{K}}^{-1}{\boldsymbol{y}}={\mathsf{V}}_{m}({\widetilde{\boldsymbol{y}}}./{\boldsymbol{\lambda}}_{m})$
+may each be evaluated at cost ${\mathcal{O}}(n\log n)$ where $\odot$ denotes the Hadamard (component-wise) product and ././ denotes component-wise division.
+The following code implements these fast Gram matrix operations for the SI-lattice variant.
+It also includes an application of (8) to update transformed values, i.e., to compute ${\widetilde{\boldsymbol{y}}}_{\mathrm{full}}=\overline{{\mathsf{V}}_{m+1}}\begin{pmatrix}{\boldsymbol{y}}\\ {\boldsymbol{y}}_{\mathrm{new}}\end{pmatrix}$ given ${\widetilde{\boldsymbol{y}}}=\overline{{\mathsf{V}}_{m}}{\boldsymbol{y}}$.
+An analogous code for the DSI-digital net variant is given in Listing LABEL:code:dsi_dnet.
+Listing 5: Efficient Gram matrix operations and transform updates for the SI-lattice variant.
+d = 3 # dimension
+m = 10 # will generate 2^m points
+n = 2**m # number of points
+lattice = qp.Lattice(d) # defaults to radical inverse order
+kernel = qp.KernelShiftInvar(
+d, # dimension
+alpha = [1,2,3], # per-dimension smoothness parameters
+lengthscales = [1, 1/2, 1/4]) # per-dimension product weights
+x = lattice(n_min=0,n_max=n) # shape=(n, d) lattice points
+y = np.random.rand(n) # shape=(n,) random uniforms
+# fast matrix multiplication and linear system solve
+k1 = kernel(x,x[0]) # shape=(n,) first column of Gram matrix
+lam = np.sqrt(n)*qp.fftbr(k1) # vector of eigenvalues
+yt = qp.fftbr(y)
+u = qp.ifftbr(yt*lam) # fast matrix multiplication
+v = qp.ifftbr(yt/lam) # fast linear system solve
+# efficient fast transform updates
+ynew = np.random.rand(n) # shape=(n,) new random uniforms
+omega = qp.omega_fftbr(m) # shape=(n,)
+ytnew = qp.fftbr(ynew) # shape=(n,)
+ytfull1 = yt+omega*ytnew # shape=(n,) first half of ytfull
+y2full2 = yt-omega*ytnew # shape=(n,) second half of ytfull
+ytfull = 1/np.sqrt(2)*np.hstack([ytfull1,y2full2]) # shape=(2n,)
+7 Numerical experiments
+All numerical experiments were carried out on a 2023 MacBook Pro with an M3 Max processor.
+Figure 4 compares the wall-clock time required to generate point sets and perform fast transforms for high dimensions and/or number of randomizations.
+Following the theory, our implementation scales linearly in the number of dimensions and randomizations.
+For large numbers of randomizations, our vectorized QMCPy generators significantly outperform the looped implementations in SciPy and PyTorch.
+For fast transforms, our FFTBR and IFFTBR implementations are the same speed as SciPy’s FFT and IFFT algorithms.
+Our FWHT implementation is significantly faster than the SymPy version, especially when applying the FWHT to multiple sequences simultaneously.
+IID points, randomly shifted lattices, and digital nets with LMS and digital shifts (including higher-order versions) are the fastest sequences to generate.
 Digital nets in base $b=2$ exploit Gray code order, integer storage of bit-vectors, and exclusive or (XOR) operations to perform digital addition.
-Generators of Halton point sets are slower to generate as they cannot exploit these advantages.
-NUS, especially higher order versions, are significantly slower to generate then LMS randomizations.
-Even so, higher-order LMS scrambling with digital shifts are sufficient to achieve higher-order RMSE convergence as we shown in the next experiment.
-Figure 5.
-Comparison of time required to generate IID and low-discrepancy point sets and perform fast transforms.
-All point sets are generated in a single dimension and with a single randomizations.
-BR indicates the bit-reversed FFT and IFFT.
+Halton point sets are slower to generate as they cannot exploit these advantages.
+NUS, especially higher-order versions, are significantly slower to generate than LMS randomizations.
+Even so, higher-order LMS scrambling with digital shifts are sufficient to achieve higher-order RMSE convergence as we show in the next experiment.
+Figure 4: Comparison of time required to generate point sets and perform fast transforms.
+For IID and randomized LD generators, we vary the number of randomizations $R$, the number of dimensions $d$, and the sequence size $n$.
+Timings include initialization, and a new point set is generating for each plotted $(R,n,d)$.
+Fast transforms are applied to $r$ sequences of size $n$ in a vectorized fashion.
+Figure 5 plots the RMSE convergence of RQMC methods applied to the integrands described below.
+Simple function, $d=1$,
+has $f(x)=xe^{x}-1$. This was used in [Dick, 2011] where higher-order digital net scrambling was first proposed.
+Simple function, $d=2$,
+has $f({\boldsymbol{x}})=x_{2}e^{x_{1}x_{2}}/(e-2)-1$. This was also considered in [Dick, 2011].
+Oakley & O’Hagan, $d=2$,
+has $f({\boldsymbol{x}})=g(({\boldsymbol{x}}-1/2)/50)$ for $g({\boldsymbol{t}})=5+t_{1}+t_{2}+2\cos(t_{1})+2\cos(t_{2})$, see [Oakley and O’hagan, 2002] or the VLSE [^17] https://www.sfu.ca/~ssurjano/uq.html (Virtual Library of Simulation Experiments).
+G-Function, $d=3$,
+has $f({\boldsymbol{x}})=\prod_{j=1}^{d}\frac{\lvert 4x_{j}-2\rvert-a_{j}}{1+a_{j}}$ with $a_{j}=(j-2)/2$ for $1\leq j\leq d$, see [Crestaux et al., 2007, Marrel et al., 2009] or the VLSE.
+Oscillatory Genz, $d=3$,
+has $f({\boldsymbol{x}})=\cos\left(-\sum_{j=1}^{d}c_{j}x_{j}\right)$ with coefficients of the third kind $c_{j}=4.5{\widetilde{c}}_{j}/\sum_{j=1}^{d}{\widetilde{c}}_{j}$ where ${\widetilde{c}}_{j}=\exp\left(j\log\left(10^{-8}\right)/d\right)$.
+This is a common test function for uncertainty quantification which is available in the Dakota [^18] https://dakota.sandia.gov/ software [Adams et al., 2020] among others.
+Corner-peak Genz, $d=3$,
+has $f({\boldsymbol{x}})=\left(1+\sum_{j=1}^{d}c_{j}x_{j}\right)^{-(d+1)}$ with coefficients of the second kind $c_{j}=0.25{\widetilde{c}}_{j}/\sum_{j=1}^{d}{\widetilde{c}}_{j}$ where ${\widetilde{c}}_{j}=1/j^{2}$.
+This is also available in Dakota.
 […]
-Figure 6 shows higher-order LMS with digital shifts for base $b=2$ digital nets achieve higher-order RMSE convergence for RQMC methods.
-Specifically, we are able to achieve an RMSE of ${\mathcal{O}}(n^{-\min\{\alpha,{\widetilde{\alpha}}\}-1/2-\delta})$ where $\alpha$ is the higher-order digital interlacing of the net, ${\widetilde{\alpha}}$ is the smoothness of the integrand, and $\delta>0$ is arbitrarily small.
-The first two integrands reproduce the experiments of (Dick, 2011).
-The next two problems are test functions from Derek Bingham’s Virtual Library of Simulation Experiments (VLSE) [^14] https://www.sfu.ca/~ssurjano/uq.html.
-The last two Genz functions are used for forward uncertainty quantification in the Dakota software package (Adams et al., 2020) among other places.
-For each problem, the RMSE of the (Q)MC estimator (1) with ${\boldsymbol{X}}\sim{\mathcal{U}}[0,1]^{d}$ is approximated using $300$ independent randomizations of a LD (or IID) point sets.
-For lattices, we periodize the integrand using a Baker-transform i.e.
-we use $\tilde{f}({\boldsymbol{x}})=f(1-2\lvert{\boldsymbol{x}}-1/2\rvert)$ in place of $f({\boldsymbol{x}})$ in (1).
-The Baker-transform does not change the expectation.
-A description of the integrands follows
-Simple function, $d=1$:
-has $f(X)=Xe^{X}-1$. This was used in (Dick, 2011) where higher-order digital net scrambling was first proposed.
-Simple function, $d=2$:
-has $f({\boldsymbol{X}})=X_{2}e^{X_{1}X_{2}}/(e-2)-1$. This was also used in (Dick, 2011).
-Oakley & O’Hagan, $d=2$:
-has $f({\boldsymbol{X}})=g(({\boldsymbol{X}}-1/2)/50)$ for $g({\boldsymbol{T}})=5+T_{1}+T_{2}+2\cos(T_{1})+2\cos(T_{2})$, see (Oakley and O’hagan, 2002) or the VLSE14.
-G-Function, $d=3$:
-has $f({\boldsymbol{X}})=\prod_{j=1}^{d}\frac{\lvert 4X_{j}-2\rvert-a_{j}}{1+a_{j}}$ with $a_{j}=(j-2)/2$ for $1\leq j\leq d$, see (Crestaux et al., 2007; Marrel et al., 2009) or the VLSE14.
-Oscillatory Genz, $d=3$:
-has $f({\boldsymbol{X}})=\cos\left(-\sum_{j=1}^{d}c_{j}X_{j}\right)$ with coefficients of the third kind $c_{j}=\exp\left(k\log\left(10^{-8}\right)/d\right)$.
-This is a common test function for uncertainty quantification which is is available in the Dakota software (Adams et al., 2020) among other.
-Corner-peak Genz, $d=3$:
-has $f({\boldsymbol{x}})=\left(1+\sum_{j=1}^{d}c_{j}X_{j}\right)^{-(d+1)}$ with coefficients of the second kind $c_{j}=1/j^{2}$.
-This is also used in the Dakota software (Adams et al., 2020) among others.
-Figure 6. The RMSE of the RQMC estimate for a few different integrands. Higher-order digital nets with linear matrix scrambling and digital shifts achieve higher-order convergence.
-[…]
-7. Conclusion and future work
-This work has reviewed routines for generating low-discrepancy point sets, randomizing them, and applying them to fast kernel methods.
-We have reviewed lattices, digital nets, and Halton point sets with randomizations spanning shifts, digital permutations, digital shifts, linear matrix scrambling, and nested uniform scrambling.
-Higher-order scramblings for digital nets were also considered.
-For kernel methods, we presented classes of kernels which pair with low-discrepancy sequences to accelerate kernel computations.
-This includes new higher-order digitally shift invariant kernels.
-The accelerated methods utilize modified fast Fourier transforms and the fast Walsh-Hadamard transform.
-The presented material has been implemented into the accessible open-source Python library QMCPy.
-Acknowledgements.
-Thank you to Fred J Hickernell, Sou-Cheng T Choi, and Aadit Jain for helpful comments and prototyping. The author would like to acknowledge the support of NSF Grant DMS-2316011.
-Appendix A Proofs of theorems
+Figure 5: The RMSE of the RQMC estimate for a few different integrands.
+Higher-order digital nets with linear matrix scrambling and digital shifts achieve higher-order convergence and significantly outperform both IID points and randomly shifted lattices.
+In the following code, we use QMCPy to build an RQMC estimate to the mean of the Corner-peak Genz function in $d=50$ dimensions.
+Using $R$ replications and a fixed number of points $n$, we construct a Student’s-$t$ confidence interval as in (2) and the discussion thereafter.
+SciPy is used to compute the necessary quantile.
+Listing LABEL:code:genz_ex_2 gives an example of an adaptive version of this algorithm which automatically selects the number of points $n$ required to meet a user-specified absolute error tolerance.
+Listing 6: Randomized QMC approximation and error estimation for a fixed number of points.
+import scipy.stats
+def gen_corner_peak_2(x):
+d = x.shape[-1] # x.shape=(...,n,d), e.g., (n,d) or (R,n,d)
+c_tilde = 1/np.arange(1,d+1)**2
+c = 0.25*c_tilde/np.sum(c_tilde)
+y = (1+np.sum(c*x,axis=-1))**(-(d+1))
+return y # y.shape=(...,n), e.g., (n,) or (R,n)
+R = 10 # number of randomizations
+n = 2**15 # number of points
+d = 50 # dimension
+dnb2 = qp.DigitalNet(dimension=d, replications=R, seed=7, alpha=3)
+x = dnb2(n) # x.shape=(R,n,d)
+y = gen_corner_peak_2(x) # y.shape=(R,n)
+muhats_r = np.mean(y,axis=1) # muhats_r.shape=(R,)
+muhat = np.mean(muhats_r) # muhat is a scalar
+print(muhat)
+""" 0.014936813948394042 """
+alpha = 0.01 # uncertainty level
+t_star = -scipy.stats.t.ppf(alpha/2,df=R-1) # quantile
+sigmahat = np.std(muhats_r,ddof=1) # unbiased estimate
+std_error = t_star*sigmahat/np.sqrt(R)
+print(std_error)
+""" 5.247445301861484e-07 """
+conf_int = [muhat-std_error,muhat+std_error]
+8 Conclusion
+This work has reviewed our QMCPy implementations of routines for generating randomized LD sequences and applying them to fast kernel methods.
+We presented lattices, digital nets, and Halton point sets with randomizations spanning shifts, digital permutations, digital shifts, LMS, and NUS.
+Higher-order scramblings for digital nets were also considered, including higher-order LMS and NUS.
+We described classes of kernels which pair with LD sequences to enable fast kernel computations.
+These included known SI kernels as well as new DSI kernels of higher-order smoothness.
+The fast kernel methods utilized the fast transform algorithms FFTBR, IFFTBR, and FWHT.
+We showed our implementations achieves state-of-the-art speed, accuracy, and flexibility.
+Acknowledgements
+Thank you to Fred J Hickernell, Sou-Cheng T Choi, and Aadit Jain for helpful comments and prototyping.
+The author would like to acknowledge the support of NSF Grant DMS-2316011 and the DOE Office of Science Graduate Student Research (SCGSR) Program.
+Appendix A Extra Code Snippets
+Listing 7: Efficient Gram matrix operations and transform updates for the DSI-digital net variant.
+d = 3 # dimension
+m = 10 # will generate 2^m points
+n = 2**m # number of points
+dnb2 = qp.DigitalNetB2(d) # defaults to radical inverse order
+kernel = qp.KernelDigShiftInvar(
+d, # dimension
+t = dnb2.t, # number of bits in integer representation of points
+alpha = [1,2,3], # per-dimension smoothness parameters
+lengthscales = [1, 1/2, 1/4]) # per-dimension product weights
+x = dnb2(n_min=0,n_max=n) # shape=(n, d) digital net
+y = np.random.rand(n) # shape=(n,) random uniforms
+# fast matrix multiplication and linear system solve
+k1 = kernel(x,x[0]) # shape=(n,) first column of Gram matrix
+lam = np.sqrt(n)*qp.fwht(k1) # vector of eigenvalues
+yt = qp.fwht(y)
+u = qp.fwht(yt*lam) # fast matrix multiplication
+v = qp.fwht(yt/lam) # fast linear system solve
+# efficient fast transform updates
+ynew = np.random.rand(n) # shape=(n,) new random uniforms
+omega = qp.omega_fwht(m) # shape=(n,)
+ytnew = qp.fwht(ynew) # shape=(n,)
+ytfull1 = yt+omega*ytnew # shape=(n,) first half of ytfull
+y2full2 = yt-omega*ytnew # shape=(n,) second half of ytfull
+ytfull = 1/np.sqrt(2)*np.hstack([ytfull1,y2full2]) # shape=(2n,)
+Listing 8: Adaptive randomized QMC approximation and error estimation.
+def gen_corner_peak_2(x):
+d = x.shape[-1] # x.shape=(...,n,d), e.g., (n,d) or (R,n,d)
+c_tilde = 1/np.arange(1,d+1)**2
+c = 0.25*c_tilde/np.sum(c_tilde)
+y = (1+np.sum(c*x,axis=-1))**(-(d+1))
+return y # y.shape=(...,n), e.g., (n,) or (R,n)
+R = 10
+d = 50
+dnb2 = qp.DigitalNet(dimension=d, replications=R, seed=7, alpha=3)
+true_measure = qp.Uniform(dnb2, lower_bound=0, upper_bound=1)
+integrand = qp.CustomFun(true_measure=true_measure, g=gen_corner_peak_2)
+# equivalent to
+# integrand = qp.Genz(dnb2, kind_func="CORNER PEAK", kind_coeff=2)
+qmc_algo = qp.CubQMCRepStudentT(integrand, abs_tol=1e-4)
+solution,data = qmc_algo.integrate() # run adaptive QMC algorithm
+print(solution)
+""" 0.014950908095474802 """
+conf_int = [data.comb_bound_low,data.comb_bound_high]
+std_error = (conf_int[1]-conf_int[0])/2
+print(std_error)
+""" 2.7968149935497788e-05 """
+print(data)
+Data (Data)
+solution 0.015
+comb_bound_low 0.015
+comb_bound_high 0.015
+comb_bound_diff 5.59e-05
+n 10240
+n_rep 2^(10)
+time_integrate 0.019
+CubQMCRepStudentT (AbstractStoppingCriterion)
+alpha 0.010
+abs_tol 1.00e-04
+rel_tol 0
+n_init 2^(8)
+DigitalNetB2 (AbstractLDDiscreteDistribution)
+d 50
+replications 10
+randomize LMS DS
+gen_mats_source joe_kuo.6.21201.txt
+order RADICAL INVERSE
+t 63
+alpha 3
+entropy 7
+Appendix B Proofs of theorems
 For $k\in{\mathbb{N}}_{0}$ write
@@
 Proof of Theorem 6.1.
-Suppose $f\in H_{\alpha}$. Using the fact that (6) is a Mercer kernel and (Dick, 2009a, Theorem 14, Remark 19), we have
-$\left\lVert f\right\rVert_{{\widetilde{H}}_{\alpha}}^{2}=\sum_{k\in{\mathbb{N}}_{0}}\left\lvert{\widehat{f}}(k)\right\rvert^{2}b^{\mu_{\alpha}(k)}\leq C_{\alpha}\sum_{k\in{\mathbb{N}}}b^{-\mu_{\alpha}(k)}<\infty$
-for some $C_{\alpha}<\infty$, so $f\in{\widetilde{H}}_{K}$. ∎
-Lemma A.1 (Walsh coefficients of low order monomials).
+Suppose $f\in H_{\alpha}^{\prime}$. Using (7) and [Dick, 2009a, Theorem 14, Remark 19], we have
+$\left\lVert f\right\rVert_{H_{\alpha}}^{2}=\sum_{k\in{\mathbb{N}}_{0}}\left\lvert{\widehat{f}}(k)\right\rvert^{2}b^{\mu_{\alpha}(k)}\leq C_{\alpha}\sum_{k\in{\mathbb{N}}}b^{-\mu_{\alpha}(k)}<\infty$
+for some $C_{\alpha}<\infty$, so $f\in H_{K}$. ∎
+lemma B.1 (Walsh coefficients of low order monomials).
 Fix $b=2$. Let $f_{p}(x):=x^{p}$. When $k\in{\mathbb{N}}$ write
@@
 where $a_{1}>a_{2}>\dots>a_{\#k}\geq 0$. Then we have
-${\widehat{f}}_{1}(k)$ =$=\begin{cases}1/2,&k=0\\ -2^{-a_{1}-2},&k=2^{a_{1}}\\ 0,&\mathrm{otherwise}\end{cases},$
-${\widehat{f}}_{2}(k)$ =$=\begin{cases}1/3,&k=0\\ -2^{-a_{1}-2},&k=2^{a_{1}}\\ 2^{-a_{1}-a_{2}-3},&k=2^{a_{1}}+2^{a_{2}}\\ 0,&\mathrm{otherwise}\end{cases},$
-${\widehat{f}}_{3}(k)$ =$=\begin{cases}1/4,&k=0\\ -2^{-a_{1}-2}+2^{-3a_{1}-5},&k=2^{a_{1}}\\ 3\cdot 2^{-a_{1}-a_{2}-4},&k=2^{a_{1}}+2^{a_{2}}\\ -3\cdot 2^{-a_{1}-a_{2}-a_{3}-5},&k=2^{a_{1}}+2^{a_{2}}+2^{a_{3}}\\ 0,&\mathrm{otherwise}\end{cases}.$
+${\widehat{f}}_{1}(k)$ $=\begin{cases}1/2,&k=0\\ -2^{-a_{1}-2},&k=2^{a_{1}}\\ 0,&\mathrm{otherwise}\end{cases},$
+${\widehat{f}}_{2}(k)$ $=\begin{cases}1/3,&k=0\\ -2^{-a_{1}-2},&k=2^{a_{1}}\\ 2^{-a_{1}-a_{2}-3},&k=2^{a_{1}}+2^{a_{2}}\\ 0,&\mathrm{otherwise}\end{cases},$
+${\widehat{f}}_{3}(k)$ $=\begin{cases}1/4,&k=0\\ -2^{-a_{1}-2}+2^{-3a_{1}-5},&k=2^{a_{1}}\\ 3\cdot 2^{-a_{1}-a_{2}-4},&k=2^{a_{1}}+2^{a_{2}}\\ -3\cdot 2^{-a_{1}-a_{2}-a_{3}-5},&k=2^{a_{1}}+2^{a_{2}}+2^{a_{3}}\\ 0,&\mathrm{otherwise}\end{cases}.$
 Proof.
-The forms for ${\widehat{f}}_{1}$ and ${\widehat{f}}_{2}$ follow from (Dick and Pillichshammer, 2010, Example 14.2, Example 14.3). For $k=0$ and any $x\in[0,1)$ we have $\mathrm{wal}_{0}(x)=1$, so
+The forms for ${\widehat{f}}_{1}$ and ${\widehat{f}}_{2}$ follow from [Dick and Pillichshammer, 2010, Example 14.2, Example 14.3]. For $k=0$ and any $x\in[0,1)$ we have $\mathrm{wal}_{0}(x)=1$, so
 ${\widehat{f}}_{3}(x)=\int_{0}^{1}x^{3}\mathrm{d}x=1/4.$
-Assume $k\in{\mathbb{N}}$ going forward. For $k=2^{a_{1}}+k^{\prime}$ with $0\leq k^{\prime}<2^{a_{1}}$, (Fine, 1949, Equation 3.6) implies
+Assume $k\in{\mathbb{N}}$ going forward. For $k=2^{a_{1}}+k^{\prime}$ with $0\leq k^{\prime}<2^{a_{1}}$, [Fine, 1949, Equation 3.6] implies
 $J_{k}(x):=\int_{0}^{x}\mathrm{wal}_{k}(t)\mathrm{d}t=2^{-a_{1}-2}\left[\mathrm{wal}_{k^{\prime}}(x)-\sum_{r=1}^{\infty}2^{-r}\mathrm{wal}_{2^{a_{1}+r}+k}(x)\right].$
-Using integration by parts and the fact that $J_{k}(0)=J_{k}(1)=0$
+Using integration by parts and the fact that $J_{k}(0)=J_{k}(1)=0$,
 ${\widehat{f}}_{3}(k)$ $=\int_{0}^{1}x^{3}\mathrm{wal}_{k}(x)\mathrm{d}x=\left[x^{3}J_{k}(x)\right]_{x=0}^{x=1}-3\int_{0}^{1}x^{2}J_{k}(x)\mathrm{d}x$
 $=-3*2^{-a_{1}-2}\left[{\widehat{f}}_{2}(k^{\prime})-\sum_{r=1}^{\infty}2^{-r}{\widehat{f}}_{2}(2^{a_{1}+r}+k)\right].$
-  • If #$\#k=1$, i.e. $k=2^{a_{1}}$ then
+  • If #k=1\#k=1, i.e., $k=2^{a_{1}}$ then
     ${\widehat{f}}_{3}(k)$ $=-3*2^{-a_{1}-2}\left[{\widehat{f}}_{2}(0)-\sum_{r=1}^{\infty}2^{-r}{\widehat{f}}_{2}(2^{a_{1}+r}+2^{a_{1}})\right]$
@@
     $=2^{-3a_{1}-5}-2^{-a_{1}-2}.$
-  • If #$\#k=2$ then
+  • If #k=2\#k=2 then
     ${\widehat{f}}_{3}(k)=-3*2^{-a_{1}-2}{\widehat{f}}_{2}(2^{a_{2}})=3\cdot 2^{-a_{1}-a_{2}-4}.$
-  • If #$\#k=3$ then
+  • If #k=3\#k=3 then
     ${\widehat{f}}_{3}(k)=-3*2^{-a_{1}-2}{\widehat{f}}_{2}(2^{a_{2}}+2^{a_{3}})=-3\cdot 2^{-a_{1}-a_{2}-a_{3}-5}.$
-  • If #$\#k>3$ then ${\widehat{f}}_{3}(k)=0$.
+  • If #k>3\#k>3 then ${\widehat{f}}_{3}(k)=0$.
 Proof of Theorem 6.2.
 Write
-${\widetilde{K}}_{\alpha}(x)=1+\sum_{1\leq\nu<\alpha}s_{\nu}(x)+{\widetilde{s}}_{\alpha}(x)$
-where $s_{\nu}$ sums over all $k$ with #$\#k=\nu$ and ${\widetilde{s}}_{\alpha}$ sums over all $k$ with #$\#k\geq\alpha$. In (Baldeaux et al., 2012, Corollary 1) it was shown that
+$K_{\alpha}(x)=\sum_{1\leq\nu<\alpha}s_{\nu}(x)+{\widetilde{s}}_{\alpha}(x)$
+where $s_{\nu}$ sums over all $k$ with $k=\nu$ and ${\widetilde{s}}_{\alpha}$ sums over all $k$ with $k\geq\alpha$. In [Baldeaux et al., 2012, Corollary 1] it was shown that
 $s_{1}(x)$ $=-2x+1,$
@@
 ${\widetilde{s}}_{3}(x)$ $=-\left[2-\beta(x)\right]x^{2}-\left[1-5t_{1}(x)\right]x+\frac{1}{18}\left[1-43t_{2}(x)\right]$
-from which ${\widetilde{K}}_{2}$ and ${\widetilde{K}}_{3}$ follow. We now find expressions for $s_{3}$ and ${\widetilde{s}}_{4}$ from which ${\widetilde{K}}_{4}$ follows.
-Assume sums over $a_{i}$ are over ${\mathbb{N}}_{0}$ unless otherwise restricted. Recall from Lemma A.1 that
+from which $K_{2}$ and $K_{3}$ follow. We now find expressions for $s_{3}$ and ${\widetilde{s}}_{4}$ from which $K_{4}$ follows.
+Assume sums over $a_{i}$ are over ${\mathbb{N}}_{0}$ unless otherwise restricted. Lemma B.1 gives
 $x$ $=\frac{1}{2}-\sum_{a_{1}}\frac{\mathrm{wal}_{2^{a_{1}}}(x)}{2^{a_{1}+2}},$
@@
 and
-$s^{[3}(x)=\sum_{a_{1}>a_{2}>a_{3}}\frac{\mathrm{wal}_{2^{a_{1}}+2^{a_{2}}+2^{a_{3}}}(x)}{2^{a_{1}+a_{2}+a_{3}+3}}=-\frac{4}{3}x^{3}+2x^{2}-\frac{2}{3}x+\frac{1}{24}\sum_{a_{1}}\frac{\mathrm{wal}_{2^{a_{1}}}(x)}{2^{3a_{1}}}.$
+$s_{3}(x)=\sum_{a_{1}>a_{2}>a_{3}}\frac{\mathrm{wal}_{2^{a_{1}}+2^{a_{2}}+2^{a_{3}}}(x)}{2^{a_{1}+a_{2}+a_{3}+3}}=-\frac{4}{3}x^{3}+2x^{2}-\frac{2}{3}x+\frac{1}{24}\sum_{a_{1}}\frac{\mathrm{wal}_{2^{a_{1}}}(x)}{2^{3a_{1}}}.$
 Now,
 ${\widetilde{s}}_{4}(x)$ $=\sum_{\begin{subarray}{c}a_{1}>a_{2}>a_{3}>a_{4}\\ 0\leq k<2^{a_{4}}\end{subarray}}\frac{\mathrm{wal}_{2^{a_{1}}+2^{a_{2}}+2^{a_{3}}+2^{a_{4}}+k}(x)}{2^{a_{1}+a_{2}+a_{3}+a_{4}+4}}$
-$=\sum_{a_{1}>a_{2}>a_{3}>a_{4}}\frac{\mathrm{wal}_{2^{a_{1}}+2^{a_{2}}+2^{a_{3}}+2^{a_{4}}}(x)}{2^{a_{1}+a_{2}+a_{3}+a_{4}+4}}\sum_{0\leq k<2^{a_{4}}}\mathrm{wal}_{k}(x)$
-If $x=0$ then
+$=\sum_{a_{1}>a_{2}>a_{3}>a_{4}}\frac{\mathrm{wal}_{2^{a_{1}}+2^{a_{2}}+2^{a_{3}}+2^{a_{4}}}(x)}{2^{a_{1}+a_{2}+a_{3}+a_{4}+4}}\sum_{0\leq k<2^{a_{4}}}\mathrm{wal}_{k}(x).$
+If $x=0$, then
 ${\widetilde{s}}_{4}(0)=\sum_{a_{1}>a_{2}>a_{3}>a_{4}}\frac{1}{2^{a_{1}+a_{2}+a_{3}+4}}=\frac{1}{294}.$
@@
 Similarity in the final two sums gives
-$W_{7}-W_{6}$ $=\left(\frac{1}{2^{\beta(x)+1}}-\sum_{\beta(x)-1>a_{3}}\frac{1}{2^{a_{3}+2}}\right)\left(\sum_{a_{1}>a_{2}>\beta(x)-1}\frac{\mathrm{wal}_{2^{a_{1}}+2^{a_{2}}}(x)}{2^{a_{1}+a_{2}+2}}\right).$
+$W_{7}-W_{6}$ $=\left(\frac{1}{2^{\beta(x)+1}}-\sum_{\beta(x)-1>a_{3}}\frac{1}{2^{a_{3}+2}}\right)\left(\sum_{a_{1}>a_{2}>\beta(x)-1}\frac{\mathrm{wal}_{2^{a_{1}}+2^{a_{2}}}(x)}{2^{a_{1}+a_{2}+2}}\right)$
 where
@@
 ${\widetilde{s}}_{4}(x)$ $=\frac{2}{3}\left(2-\beta(x)\right)x^{3}+\left(1-5\ t_{1}(x)\right)x^{2}-\frac{1}{9}\left(1-43t_{2}(x)\right)x$
-$-\frac{1}{48}\left(2-\beta(x)\right)\sum_{a_{1}}\frac{\mathrm{wal}_{2^{a_{1}}}(x)}{2^{3a_{1}}}-\frac{1}{294}\left(7\beta(x)+701t_{3}(x)\right)+\frac{5}{98}$
+$-\frac{1}{48}\left(2-\beta(x)\right)\sum_{a_{1}}\frac{\mathrm{wal}_{2^{a_{1}}}(x)}{2^{3a_{1}}}-\frac{1}{294}\left(7\beta(x)+701t_{3}(x)\right)+\frac{5}{98},$
 from which the result follows. ∎